### v2

# Rustam Ibragimova,b    Jihyun Kimc    Anton Skrobotovd a Imperial College Business School b New Economic School

Abstract
This paper develops robust inference methods for predictive regressions that address key challenges posed by endogenously persistent or heavy-tailed regressors, as well as persistent volatility in errors. Building on the Cauchy estimation framework, we propose two novel tests: one based on $t$-statistic group inference and the other employing a hybrid approach that combines Cauchy and OLS estimation. These methods effectively mitigate size distortions that commonly arise in standard inference procedures under endogeneity, near nonstationarity, heavy tails, and persistent volatility. The proposed tests are simple to implement and applicable to both continuous- and discrete-time models. Extensive simulation experiments demonstrate favorable finite-sample performance across a range of realistic settings. An empirical application examines the predictability of excess stock returns using the dividend–price and earnings–price ratios as predictors. The results suggest that the dividend–price ratio possesses predictive power, whereas the earnings–price ratio does not significantly forecast returns.
Keywords: predictive regressions, robust inference, near nonstationarity, heterogeneity, heavy tails, persistent volatility, endogeneity.
JEL Codes: C12, C22
1 Introduction
Predictive regressions play a central role in empirical finance, providing a framework for assessing whether financial or macroeconomic variables can forecast future returns. Prominent applications include the forecasting of equity and aggregate returns (see, among others, CampbellYogo2006, CampbellYogo2006; GoyalWelch, GoyalWelch; campbell, campbell, Hirshleifer, Hirshleifer; KJ, KJ; Rapach, Rapach; MR, MR; Goyal2024, Goyal2024, and references therein) and tests of market efficiency (e.g., Fama1, Fama1, Fama, Fama2, the review in Martin, Martin, and references therein). Numerous studies have examined the econometric properties of predictive regressions for stock returns (see Phillips2015, Phillips2015, for a review), highlighting several statistical challenges that complicate inference on return predictability. Widely used predictors, including valuation ratios such as the dividend–price and earnings–price ratios, typically exhibit near–unit–root dynamics, and their innovations are correlated with stock returns over long horizons. This combination of persistence and endogeneity induces substantial biases in conventional hypothesis tests (see, e.g., stambaugh1999predictive, stambaugh1999predictive; KimPark1, KimPark1). Moreover, stock return volatility is stochastic and highly persistent (jacquier2004bayesian, jacquier2004bayesian; hansen2014estimating, hansen2014estimating), and cavaliere2004testing shows that such volatility can lead to severe size distortions in tests assuming stationarity. Predictive regression data also often exhibit heavy tails, jumps, structural breaks, and regime shifts, further undermining standard inference (op. cit.).
A large literature has addressed persistent endogeneity in predictive regressions. Notably, CampbellYogo2006, ChenDeo2009, PhillipsMagdalinos2009, and KMS2015, among others, proposed inference procedures designed to handle persistence and endogeneity. While these methods perform well in such environments, they do not adequately address other empirical features such as persistent volatility, structural breaks, or regime switching. The simulations in ibragimov2023new show that standard tests can suffer from severe size distortions under persistent volatility.
CJP2016 proposed an inference method (the Cauchy RT) based on the Cauchy estimator and a time-change transformation in a continuous-time framework to address endogenous persistence and volatility in the errors. [^2] See also BuKimWang for an alternative method robust to endogenously persistent or heavy-tailed regressors and persistent volatility in continuous time. ibragimov2023new introduced another approach (the Cauchy VC), also based on the Cauchy estimator but with a nonparametric volatility correction. Unlike CJP2016, it applies to both continuous- and discrete-time models.
This paper proposes two practical tests that serve as robust alternatives to these methods. The proposed tests are robust to heterogeneous and persistent volatility, as well as to endogenous, persistent, and/or heavy-tailed regressors. Both employ Cauchy estimation procedures, as in CJP2016 and ibragimov2023new, to address endogeneity, persistence, and heavy tails. The two methods differ in their treatment of heterogeneous volatility: the first extends the $t$-statistic-based group inference of IbragimovMuller2010 to asymptotically normal Cauchy estimators, while the second is a hybrid test combining Cauchy and OLS estimation that uses the Cauchy estimator for the coefficient and OLS residuals for the standard error.
The proposed methods are easy to implement and avoid the technical complexities of the time-change transformation in CJP2016 and the nonparametric volatility correction in ibragimov2023new. Although they rely on an asymptotically exogenous volatility condition, they perform well in finite samples and compare favorably to alternative procedures across empirically relevant scenarios, including mild violations of this condition. Moreover, the proposed methods apply to both continuous- and discrete-time models, as in ibragimov2023new. Overall, the two tests developed here and the inference procedures of CJP2016 and ibragimov2023new should be viewed as complementary, particularly in environments with heteroskedastic and persistent volatility.
The remainder of the paper is organized as follows. Section 2 discusses the predictive regression model and the Cauchy estimator; Section 3 develops the inference procedures and theoretical results; Section 4 extends them to multiple predictors and models with intercepts; Sections 5 and 6 present simulations and an empirical illustration; and Section 7 concludes. All proofs are in the Appendix.
2 Predictive Regressions and the Cauchy Estimator
2.1 Model and Issues
Throughout the paper, we consider $(\mathcal{F}_{t})$-adapted processes defined on a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t}),P)$ equipped with an increasing filtration $(\mathcal{F}_{t})$ of sub-$\sigma$-fields of $\mathcal{F}$. Our objective is to test the (un)predictability of the process $(y_{t})$ (e.g., the time series of excess stock returns) based on a covariate process $(x_{t})$ (e.g., the time series of price–dividend ratios). As usual, we consider the linear predictive regression model
$y_{t}$ $=\alpha+\beta x_{t-1}+u_{t},\quad t=1,\dots,T.$ (1)
Following the standard specification for a volatility model, we assume that
$u_{t}=v_{t}\varepsilon_{t},$
where $(v_{t})$ is a volatility process and $(\varepsilon_{t})$ is a martingale difference sequence (MDS) with respect to $(\mathcal{F}_{t})$. We impose the following regularity conditions on $(v_{t},\varepsilon_{t})$.
Assumption 2.1.
(a) $E(\varepsilon_{t}^{2}|\mathcal{F}_{t-1})=1$; (b) $(v_{t})$ is $(\mathcal{F}_{t-1})$-adapted and nonnegative; and (c) $T^{-1}\sum_{t=1}^{T}E\!\left[\varepsilon_{t}^{2}1\{|\varepsilon_{t}|\geq\delta\sqrt{T}\}\middle|\mathcal{F}_{t-1}\right]\!\to_{p}0$ for any $\delta>0$.
Conditions (a) and (b) are standard and ensure that the conditional variance of $u_{t}$ is identified: $E(u_{t}^{2}|\mathcal{F}_{t-1})=v_{t}^{2}$. Condition (c) is a conditional Lindeberg condition, which holds, for example, if $\sup_{1\leq t\leq T}E(|\varepsilon_{t}|^{2+\delta}|\mathcal{F}_{t-1})$ is bounded for some $\delta>0$ with probability one. See ibragimov2023new and references therein for further discussion and examples of processes satisfying Assumption 2.1.
The hypothesis of unpredictability of $(y_{t})$ corresponds to $H_{0}:\beta=0$ in regression (1). It is well-known that standard OLS $t$-statistic inference is not robust to many empirically relevant features of financial data. For instance, the OLS estimator of $\beta$ is not asymptotically Gaussian under $H_{0}$ if $(x_{t})$ is endogenous and (nearly) nonstationary (see elliott1994inference, elliott1994inference; PhillipsNear, PhillipsNear; GP, GP; PM, PM; KMS2015, KMS2015), or even if $(x_{t})$ is stationary but has infinite variance (e.g., GrangerOrr, GrangerOrr; EKM, EKM; ibragimov2015heavy, ibragimov2015heavy). This non-Gaussianity persists even when the errors are homoskedastic with $v_{t}^{2}=\sigma^{2}$ for all tt. [^3] As usual, the endogeneity of $(x_{t-1})$ refers to the existence of nonzero long-run covariance between the innovations of $(y_{t})$ and $(x_{t-1})$. Furthermore, stock return data exhibit time-varying and stochastically persistent volatility, which causes the distribution of the OLS $t$-statistic to deviate from standard normality, leading to size distortions in conventional tests (see CJP2016, CJP2016; ibragimov2023new, ibragimov2023new).
2.2 The Cauchy Estimator
Both inference methods proposed in this paper build upon the following Cauchy estimator of $\beta$ (assuming no intercept, i.e., $\alpha=0$):
$\check{\beta}=\left(\sum_{t=1}^{T}|x_{t-1}|\right)^{-1}\sum_{t=1}^{T}\text{sign}(x_{t-1})\,y_{t},$
where $\text{sign}(\cdot)$ denotes the sign function, $\text{sign}(x)=1$ for $x\geq 0$ and $\text{sign}(x)=-1$ for $x<0$. The estimator $\check{\beta}$ can be interpreted as an instrumental variable (IV) estimator using $\text{sign}(x_{t-1})$ as an instrument for $x_{t-1}$ (see, e.g., SoShin, SoShin; BREITUNG2015, BREITUNG2015; kim-meddahi-2019, kim-meddahi-2019; shephard2020, shephard2020).
Under Assumption 2.1, $\text{sign}(x_{t-1})\varepsilon_{t}$ (denoted by $\xi_{t}$) is a unit-variance MDS with respect to $(\mathcal{F}_{t})$. Define the continuous-time partial sum process $(W^{T}(r),0\leq r\leq 1)$ by
$W^{T}(r)=T^{-1/2}\sum_{t=1}^{[Tr]}\xi_{t},$
which takes values in $\mathbf{D}_{\mathbb{R}}[0,1]$, the space of càdlàg functions on $[0,1]$ with values in $\mathbb{R}^{d}$. By the functional central limit theorem for martingales (Theorem 18.2 of billingsley1986convergence, billingsley1986convergence), we have $W^{T}\Rightarrow W$ in $\mathbf{D}_{\mathbb{R}}[0,1]$, where $W$ is a standard Brownian motion.
For the volatility process $(v_{t})$, define $\sigma^{T}(r)=v_{[Tr]}$ on $\mathbf{D}_{\mathbb{R}^{+}}[0,1]$. Then the Cauchy estimator can be expressed in terms of $\sigma^{T}$ and $W^{T}$ as
$\left(\sum_{t=1}^{T}|x_{t-1}|/\sqrt{T}\right)\!\bigl(\check{\beta}-\beta\bigr)$ $=T^{-1/2}\sum_{t=1}^{T}\text{sign}(x_{t-1})v_{t}\varepsilon_{t}=\int_{0}^{1}\sigma^{T}(r)\,dW^{T}(r).$
Following ibragimov2023new, we assume that the volatility process $\sigma^{T}$ is persistent in the sense that it admits a limiting process $\sigma$ defined on $[0,1]$ such that $(W^{T},\sigma^{T})\Rightarrow(W,\sigma)$ jointly.
Assumption 2.2.
There exists a nonnegative process $\sigma$ on $\mathbf{D}_{\mathbb{R}^{+}}[0,1]$ such that
$0<\int_{0}^{1}\sigma^{2}(r)\,dr<\infty,\quad\text{and}\quad(W^{T},\sigma^{T})\Rightarrow(W,\sigma)\quad\text{in}\quad\mathbf{D}_{\mathbb{R}\times\mathbb{R}^{+}}[0,1],$
where $W$ is a standard Brownian motion adapted to the same filtration as $\sigma$.
Assumption 2.2 encompasses a wide class of models, including those with nonstationary volatility, regime switching, or structural breaks. [^4] Assumptions 2.1 and 2.2 exclude some globally homoskedastic processes, such as stationary GARCH models. However, the hybrid testing procedure proposed later remains valid under $T^{-1}\sum_{t=1}^{T}v_{t}^{2}\to_{p}\omega^{2}>0$, which includes conditionally heteroskedastic but globally homoskedastic processes, such as stationary GARCH models (see also Section 4 of ibragimov2023new, ibragimov2023new). It also covers cases with deterministic volatility $v_{t}=\sigma(t/T)$, as in cavaliere-taylor-2007, cavaliere-taylor-2008, xu-phillips-2008, and harvey-leybourne-zu, among others. [^5] Assumption 2.2 is a simplified version of the condition $v_{[Tr]}/a_{T}\Rightarrow\sigma_{r}$ in Assumption 2 of cavaliere-taylor-2009. We focus on stochastically bounded volatilities with $a_{T}=1$, excluding explosive volatility settings ($a_{T}\to\infty$) for simplicity. It further includes nonstationary volatility processes such as those in hansen1995regression and chung2007nonstationary, where $v_{t}^{2}$ is a smooth positive transformation of a (near) unit root process. Overall, Assumptions 2.1 and 2.2 are general enough to allow for stochastic and discontinuous volatility—features commonly observed in financial returns.
Under Assumptions 2.1 and 2.2, the properly normalized Cauchy estimator satisfies
$\left(\sum_{t=1}^{T}|x_{t-1}|/\sqrt{T}\right)\!(\check{\beta}-\beta)\Rightarrow\int_{0}^{1}\sigma(r)\,dW(r),$
by standard results on the convergence of stochastic integrals (see hansen-1992, hansen-1992; kurtz1991weak, kurtz1991weak; ibragimov2023new, ibragimov2023new). The limit $\int_{0}^{1}\sigma(r)\,dW(r)$ is in general a non-Gaussian martingale, becoming Gaussian only if $W$ and $\sigma$ are independent. In that case, $\int_{0}^{1}\sigma(r)\,dW(r)$ is a scale mixture of normals with variance $\int_{0}^{1}\sigma^{2}(r)\,dr$, denoted
$\int_{0}^{1}\sigma(r)\,dW(r)\;=_{d}\;\mathbb{MN}\!\left(0,\int_{0}^{1}\sigma^{2}(r)\,dr\right).$
We formalize the independence assumption as follows.
Assumption 2.3.
The processes $\sigma$ and $W$ in Assumption 2.2 are independent.
Assumption 2.3 requires the volatility process $\sigma^{T}$ to be asymptotically independent of the martingale $W^{T}$, but does not preclude finite-sample dependence. For example, consider
$\sigma^{T}\!\left(t/T\right)=T^{-\delta}f(x_{t-1},\varepsilon_{t})+\sigma_{0}^{T}\!\left(t/T\right),\quad\delta>0,$
where $f:\mathbb{R}^{2}\to\mathbb{R}^{+}$ is bounded and $\sigma_{0}^{T}$ is independent of $W^{T}$ with $(W^{T},\sigma_{0}^{T})\Rightarrow(W,\sigma)$, where $W$ and $\sigma$ are independent. For any $\delta>0$, the volatility process $\sigma^{T}$ in this example satisfies Assumption 2.3, even though $\sigma^{T}$ and $W^{T}$ may be dependent for any fixed $T>0$.
In the following sections, we develop inference methods based on the Cauchy estimator. Section 3 focuses on predictive regressions with a single predictor and no intercept, while Section 4 extends the analysis to models with multiple predictors and an intercept.
3 Robust Inference for Predictive Regressions
3.1 Robust $t$-Statistic Approach
The first approach relies on $t$-statistic-based inference using group estimates of $\beta$, as proposed by IbragimovMuller2010 (see also IM1, IM1; Section 3.3 of ibragimov2015heavy, ibragimov2015heavy). The method is based on normalized Cauchy estimators—specifically, the numerator of the Cauchy estimator divided by $\sqrt{T}$ in the full-sample case:
$\check{\gamma}=\frac{\sum_{t=1}^{T}|x_{t-1}|}{\sqrt{T}}\check{\beta}=\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\text{sign}(x_{t-1})y_{t}.$ (2)
Following the $t$-statistic approach, we partition the sample into a fixed number $q\geq 2$ of approximately equal groups of consecutive observations. The observation $(y_{t},x_{t-1})$ at time $t$ belongs to the $j$th group $\mathcal{G}_{j}$ if
$t\in\mathcal{G}_{j}=\{s:(j-1)[T/q]<s\leq j[T/q]\},\quad j=1,\dots,q.$
We compute the normalized Cauchy statistic in (2) within each group:
$\check{\gamma}_{j}=\sqrt{\frac{q}{T}}\sum_{t=(j-1)[T/q]+1}^{j[T/q]}\text{sign}(x_{t-1})y_{t}.$ (3)
The $t$-statistic based on the $q$ group statistics $\{\check{\gamma}_{j}\}_{j=1}^{q}$ is given by
$t_{q}(\check{\gamma})=\sqrt{q}\,\frac{\bar{\gamma}}{s_{\gamma}},$ (4)
where
$\bar{\gamma}=q^{-1}\sum_{j=1}^{q}\check{\gamma}_{j},\qquad s_{\gamma}^{2}=(q-1)^{-1}\sum_{j=1}^{q}(\check{\gamma}_{j}-\bar{\gamma})^{2}.$
Under the null hypothesis $H_{0}:\beta=0$, the test rejects $H_{0}$ in favor of $H_{A}:\beta\neq 0$ if $|t_{q}(\check{\gamma})|>cv_{q}(\alpha)$, where $cv_{q}(\alpha)$ denotes the two-sided $t$-critical value at level $\alpha$, i.e. $P(|T_{q-1}|>cv_{q}(\alpha))=\alpha$ for $T_{q-1}\sim t_{q-1}$ (one-sided tests are analogous).
To study the asymptotic behavior of $\{\check{\gamma}_{j}\}_{j=1}^{q}$, we decompose
$\check{\gamma}_{j}=\zeta_{j}+\psi_{j},$
where
$\zeta_{j}=\beta\sqrt{\frac{q}{T}}\sum_{t\in\mathcal{G}_{j}}|x_{t-1}|,\qquad\psi_{j}=\sqrt{\frac{q}{T}}\sum_{t\in\mathcal{G}_{j}}\text{sign}(x_{t-1})u_{t}.$
Under Assumption 2.1, $\{\psi_{j}\}_{j=1}^{q}$ forms a sequence of martingale differences uncorrelated across groups, yielding the following asymptotic characterization.
Lemma 3.1.
Let Assumptions 2.1, 2.2, and 2.3 hold. For any fixed $q\geq 2$ and $\beta\in\mathbb{R}$,
$(\check{\gamma}_{1}-\zeta_{1},\dots,\check{\gamma}_{q}-\zeta_{q})^{\prime}\to_{d}\mathbb{MN}\bigl(0,\mathrm{diag}(q\omega_{1}^{2},\dots,q\omega_{q}^{2})\bigr),$
where $\omega_{j}^{2}=\int_{(j-1)/q}^{j/q}\sigma^{2}(r)\,dr$ for $j=1,\dots,q$.
The statistics $\{\check{\gamma}_{j}\}_{j=1}^{q}$ do not satisfy the standard condition in IbragimovMuller2010, which requires estimators $\{\tilde{\beta}_{j}\}_{j=1}^{q}$ such that
$\{m_{T}(\tilde{\beta}_{j}-\beta)\}_{j=1}^{q}\to_{d}\{V_{j}Z_{j}\}_{j=1}^{q},$
for some $m_{T}\to\infty$, $Z_{j}\stackrel{{\scriptstyle iid}}{{\sim}}\mathbb{N}(0,1)$, and $\{V_{j}\}$ independent of $\{Z_{j}\}$. By contrast, Lemma 3.1 shows that $\{\check{\gamma}_{j}\}_{j=1}^{q}$ lack such a diverging normalization. Consequently, as shown in Proposition 3.2, the $t$-statistic approach yields correct asymptotic size but is consistent only for a restricted class of covariates, excluding (near) unit-root processes. This inconsistency arises precisely because the asymptotics of $\{\check{\gamma}_{j}\}_{j=1}^{q}$ do not involve a diverging sequence (see proofs of Proposition 3.2 and Corollary 3.3).
Nevertheless, with additional regularity conditions, if $\{c_{T}^{-1}\sum_{t\in\mathcal{G}_{j}}|x_{t-1}|\}_{j=1}^{q}\to_{d}\{D_{j}\}_{j=1}^{q}$ for positive random variables $\{D_{j}\}$ and a sequence $c_{T}/\sqrt{T}\to\infty$, then the Cauchy estimator $\check{\beta}_{j}$ computed within each group satisfies
$\{m_{T}(\check{\beta}_{j}-\beta)\}_{j=1}^{q}\to_{d}\{P_{j}\}_{j=1}^{q},$
for $m_{T}=c_{T}\sqrt{q/T}$. In general, however, $\{P_{j}\}_{j=1}^{q}$ are non-Gaussian, especially when $(x_{t})$ is (near) unit root and endogenous. Applying the $t$-statistic approach to $\{\check{\beta}_{j}\}_{j=1}^{q}$ thus yields consistency for broader classes of covariates but may incur size distortions due to non-Gaussianity.
Proposition 3.2.
Let Assumptions 2.1, 2.2, and 2.3 hold, with fixed $q\geq 2$ and $\alpha\leq 0.83$.
(a) Under $H_{0}:\beta=0$,
$\lim_{T\to\infty}\mathbb{P}(|t_{q}(\check{\gamma})|>cv_{q}(\alpha)\mid H_{0})\leq\alpha.$
(b) Under $H_{A}:\beta\neq 0$, suppose $(x_{t})$ is stationary with $E|x_{t}|<\infty$ and satisfies
$\sup_{1\leq s\leq T-T/q}\Biggl|E|x_{t}|-\frac{1}{T/q}\sum_{t=s}^{s+T/q}|x_{t}|\Biggr|\to_{p}0.$
Then
$\lim_{T\to\infty}\mathbb{P}(|t_{q}(\check{\gamma})|>cv_{q}(\alpha)\mid H_{A})=1.$
Proposition 3.2 shows that the $t$-statistic approach is conservative under $H_{0}$ and consistent under $H_{A}$ when $(x_{t})$ is stationary with a finite first moment. It is thus valid and robust to persistent heteroskedasticity and endogenously heavy-tailed covariates. However, it becomes inconsistent for highly persistent covariates, such as (near) unit-root processes. To illustrate, consider the generalized local-to-unity framework of dou2021generalized, where $X^{T}(r)=x_{[Tr]}$ for $r\in[0,1]$ and
$T^{-1/2}\bigl(X^{T}(\cdot)-X^{T}(0)\bigr)\to_{d}X(\cdot)-X(0),$ (5)
with $X$ a stationary continuous-time Gaussian ARMA process. [^6] See dou2021generalized for a detailed discussion.
Corollary 3.3.
Let Assumptions 2.12.3 hold and suppose $(x_{t})$ satisfies (5). Under $\beta\neq 0$, $t_{q}(\check{\gamma})\to_{d}\text{sign}(\beta)\mathcal{D}_{q}$ for $q\geq 2$, where
$\mathcal{D}_{q}=\int_{0}^{1}|X(r)|dr\left(\frac{q(q-1)}{\sum_{j=1}^{q}\bigl(\int_{0}^{1}|X(r)|dr-q\int_{(j-1)/q}^{j/q}|X(r)|dr\bigr)^{2}}\right)^{1/2},$
and $\mathcal{D}_{q}>(q-1)^{-1/2}$ with probability one.
When $(x_{t})$ is highly persistent, $t_{q}(\check{\gamma})$ converges to $\mathcal{D}_{q}$ rather than diverging, with lower bound $(q-1)^{-1/2}$. Simulations in Section 4 confirm that rejection probabilities remain high even when $t_{q}(\check{\gamma})$ is asymptotically bounded. For $q=2$,
$\mathcal{D}_{2}=\frac{\int_{0}^{1}|X(r)|dr}{\bigl|\int_{0}^{1/2}|X(r)|dr-\int_{1/2}^{1}|X(r)|dr\bigr|}>1.$ (6)
The ratio form in (6) implies large realizations of $\mathcal{D}_{2}$ in finite samples, producing high rejection rates even under inconsistency. Figure 1 plots the simulated density of $\mathcal{D}_{2}$ when $X$ is Brownian motion. [^7] Based on 100,000 simulated draws. The minimum simulated value is 1.15, and $\mathbb{P}(|\mathcal{D}_{2}|>cv_{2}(0.05))=0.15$ with $cv_{2}(0.05)=4.303$.
Figure 1: Simulated density of $\mathcal{D}_{2}$ in (6).
3.2 A Hybrid Test
We now propose a simple hybrid test that remains consistent for a broad class of covariates. Under Assumptions 2.12.3,
$\frac{\sum_{t=1}^{T}|x_{t-1}|}{\sqrt{T}}(\check{\beta}-\beta)=\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\text{sign}(x_{t-1})u_{t}\to_{d}\int_{0}^{1}\sigma(r)\,dW(r)=\omega Z,$
where $Z\sim\mathbb{N}(0,1)$ and $\omega^{2}=\int_{0}^{1}\sigma^{2}(r)\,dr$.
A key feature of the Cauchy estimator $\check{\beta}$ is that its properly normalized limit distribution is invariant to the data-generating process of $(x_{t})$. By contrast, the OLS estimator’s variance depends on both $(u_{t})$ and $(x_{t})$, complicating variance estimation even under homoskedasticity. For $\check{\beta}$, the asymptotic variance depends solely on $u_{t}$, requiring only heteroskedasticity-robust adjustments. [^8] See shephard2020, Section 4.3, for related discussion.
We define the hybrid test statistic as
$\tau(\check{\beta})=\frac{\check{\gamma}}{\hat{\omega}}$,
where $\check{\gamma}=(\sum_{t=1}^{T}|x_{t-1}|/\sqrt{T})\check{\beta}$ as in (2), and
$\hat{\omega}^{2}=\frac{1}{T}\sum_{t=1}^{T}\hat{u}_{t}^{2},\qquad\hat{u}_{t}=y_{t}-\hat{\beta}x_{t-1}.$
Here, $\hat{\omega}^{2}$ estimates $\omega^{2}=\int_{0}^{1}\sigma^{2}(r)\,dr$ using OLS residuals. As noted by shephard2020, the Cauchy-based variance estimator performs poorly because the Cauchy estimator converges more slowly and less efficiently than OLS when $(x_{t})$ is heavy-tailed or nearly integrated. Hence, we use OLS residuals to improve efficiency. [^9] A related approach is employed by KMS2015 in the IVX framework of PM.
We assume:
Assumption 3.1.
$\sum_{t=1}^{T}x_{t-1}u_{t}=o_{p}\!\left(\sqrt{T\sum_{t=1}^{T}x_{t-1}^{2}}\right)$.
Assumption 3.1 is very general and holds in many time-series settings. It is weaker than Assumption 3.2 of ibragimov2023new, which requires $\sum_{t=1}^{T}x_{t-1}u_{t}=O_{p}\!\left(\sqrt{T^{p}\sum_{t=1}^{T}x_{t-1}^{2}}\right)$ for $p\in[0,1/16)$. As shown in ibragimov2023new, this holds with $p=0$ when $(x_{t})$ is either (near) unit root or stationary with finite variance; it also applies to certain stationary heavy-tailed processes (see, e.g., Samorodnitsky2007, Samorodnitsky2007).
Under Assumption 3.1,
$|\hat{\beta}-\beta|=o_{p}\!\left(\sqrt{T\big/\sum_{t=1}^{T}x_{t-1}^{2}}\right),\quad\text{and hence}\quad\hat{\omega}^{2}\to_{p}\omega^{2}.$
The asymptotic properties of $\tau(\check{\beta})$ follow.
Proposition 3.4.
Let Assumptions 2.1, 2.2, 2.3, and 3.1 hold.
(a) Under $H_{0}:\beta=0$,
$\tau(\check{\beta})\to_{d}\mathbb{N}(0,1).$
(b) Under $H_{A}:\beta\neq 0$,
$\tau(\check{\beta})=\beta\frac{\sum_{t=1}^{T}|x_{t-1}|}{\omega\sqrt{T}}(1+o_{p}(1))+O_{p}(1),$
so $|\tau(\check{\beta})|\to_{p}\infty$ whenever $\sum_{t=1}^{T}|x_{t-1}|/\sqrt{T}\to_{p}\infty$.
The conclusions of Proposition 3.4 remain valid under weaker conditions. For instance, if Assumptions 2.1 and 3.1 hold and
$\frac{1}{T}\sum_{t=1}^{T}v_{t}^{2}\to_{p}\omega^{2}>0,\qquad\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\text{sign}(x_{t-1})u_{t}\to_{d}\omega Z,$
where $Z\sim\mathbb{N}(0,1)$ is independent of $\omega^{2}$, then $\tau(\check{\beta})$ retains its asymptotic validity. These conditions include stationary volatility with $E[v_{t}^{2}]=\omega^{2}$. Hence, Assumptions 2.2 and 2.3 can be interpreted as primitive sufficient conditions accommodating persistent volatility in predictive regression data.
Remark. Proposition 3.4(a) also holds if $\tau(\check{\beta})$ uses $\bar{\omega}^{2}=T^{-1}\sum_{t=1}^{T}y_{t}^{2}$ instead of $\hat{\omega}^{2}$, since $\beta=0$ under $H_{0}$. Moreover, the corresponding test remains consistent when $(x_{t})$ is stationary with finite variance or follows a generalized local-to-unity process (dou2021generalized). However, it can be inconsistent for heavy-tailed $(x_{t})$. For instance, if $(x_{t})$ is i.i.d. $\alpha$-stable with $\alpha\in(0,2)$ and independent of $(u_{t})$, then
$\bar{\omega}^{2}=\beta^{2}\left(\frac{1}{T}\sum_{t=1}^{T}x_{t-1}^{2}\right)(1+o_{p}(1)),\quad\tau(\check{\beta})=\frac{\sum_{t=1}^{T}|x_{t-1}|}{\sqrt{\sum_{t=1}^{T}x_{t-1}^{2}}}(1+o_{p}(1))=O_{p}(1),$
by the generalized central limit theorem (see feller1971, feller1971; logan1973, logan1973; Davis1983, Davis1983; DavisResnick1986, DavisResnick1986). Thus, the use of $\hat{\omega}^{2}$ (or another consistent estimator under both $H_{0}$ and $H_{A}$) is crucial for the consistency of the hybrid test.
4 Extensions
This section extends the inference methods developed in Section 3 to models with multiple predictors and to regressions including an intercept. Our goal is not to design efficient procedures but to provide simple and robust inference methods that rely on minimal assumptions on the predictors and volatility processes.
4.1 Predictive Regressions with Multiple Predictors
Consider a predictive regression with $K$ predictors $x_{t}=[x_{1,t},\ldots,x_{K,t}]^{\prime}$:
$y_{t}$ $=x_{t-1}^{\prime}B+u_{t}$
$=\beta_{1,K}x_{1,t-1}+\cdots+\beta_{K,K}x_{K,t-1}+u_{t},\qquad B=[\beta_{1,K},\ldots,\beta_{K,K}]^{\prime}.$ (7)
The objective is to test the joint predictability of the covariates, that is,
$H_{0}:\beta_{1,K}=\cdots=\beta_{K,K}=0.$
We construct a testing procedure for $H_{0}$ based on the univariate inference methods in Section 3. Specifically, we estimate $K$ univariate predictive regressions
$y_{t}=\beta_{k}x_{k,t-1}+u_{k,t},\qquad k=1,\ldots,K,$
and test each null hypothesis
$H_{0}^{(k)}:\beta_{k}=0,\qquad k=1,\ldots,K.$
Clearly, $H_{0}$ implies $H_{0}^{(k)}$ for all $k$. The converse also holds under mild regularity conditions, as shown below.
Lemma 4.1.
Consider model (7) and define $z_{t}=[\operatorname{sign}(x_{1,t}),\ldots,\operatorname{sign}(x_{K,t})]^{\prime}$. Suppose that for each $t\in\{1,\dots,T\}$, $\mathbb{E}[z_{t-1}u_{t}]=\mathbf{0}_{K\times 1}$, $0<\mathbb{E}[|x_{k,t-1}|]<\infty$ for all $k$, and that the matrix $\mathbb{E}[z_{t-1}x_{t-1}^{\prime}]$ is invertible. [^10] Even when $\operatorname{sign}(x_{t-1})$ is constant, the univariate Cauchy estimator and associated tests remain well defined. In this case, the estimator simplifies to $\check{\beta}=\sum_{t=1}^{T}y_{t}\big/\sum_{t=1}^{T}x_{t-1}$, implying $\check{\beta}-\beta=\sum_{t=1}^{T}u_{t}\big/\sum_{t=1}^{T}x_{t-1}$, and Proposition 3.4 continues to hold. Hence, the lack of sign variation does not affect the validity of our methods in the univariate case. In the multiple-predictor case, however, the invertibility of $\mathbb{E}[z_{t-1}x_{t-1}^{\prime}]$ imposes mild restrictions on sign changes across predictors. For practical applications, one can induce variation in the sign instrument by recentering the predictor, for example, $\tilde{x}_{t-1}=x_{t-1}-t^{-1}\sum_{s=1}^{t}x_{s-1}$, which preserves both the martingale structure and the asymptotic validity of the estimator. Then, the joint null hypothesis $H_{0}$ holds if and only if $H_{0}^{(k)}$ holds for all $k=1,\ldots,K$.
Lemma 4.1 justifies the use of multiple hypothesis testing based on univariate Cauchy estimators. [^11] See Harvey2015 for an application of the multiple-testing framework in predictive regressions, and KMS2015 for joint-predictability tests in the IVX framework. Note that the IVX approach may lose validity under heavy-tailed predictors or continuous-time data, whereas our method remains robust in such settings. In conjunction with the hybrid test introduced in Section 3.2, we compute the statistic $\tau(\check{\beta}_{k})$ for each parameter $\beta_{k}$, where $\check{\beta}_{k}$ denotes the corresponding Cauchy estimator. Let $p_{k}$ denote its $p$-value. The joint null hypothesis $H_{0}$ is rejected at level $\alpha$ if $\min_{k}p_{k}\leq\alpha/K$, following the Bonferroni correction.
This approach directly extends the univariate robust inference procedure to a multivariate setting and requires only mild conditions for the equivalence between $H_{0}$ and $\{H_{0}^{(k)}\}_{k=1}^{K}$. The Bonferroni correction imposes no assumptions on the joint distribution of the test statistics, which motivates its use here (see Holm1979, Holm1979; BenjaminiHochberg1995, BenjaminiHochberg1995; Shaffer1995, Shaffer1995).
We also note that if one additionally assumes that $\mathbb{E}[z_{t}z_{t}^{\prime}]$ is invertible, the joint hypothesis can be tested directly using a Wald-type statistic:
$W=\left(\sum_{t=1}^{T}z_{t-1}y_{t}\right)^{\prime}\left(\hat{\omega}^{2}\sum_{t=1}^{T}z_{t-1}z_{t-1}^{\prime}\right)^{-1}\left(\sum_{t=1}^{T}z_{t-1}y_{t}\right).$
In particular, under $H_{0}$,
$\left(\hat{\omega}^{2}\sum_{t=1}^{T}z_{t-1}z_{t-1}^{\prime}\right)^{-1/2}\left(\sum_{t=1}^{T}z_{t-1}y_{t}\right)\to_{d}\mathbb{N}(0,I_{K}),$
and hence $W\to_{d}\chi_{K}^{2}$. [^12] As mentioned earlier, $z_{t}$ can be interpreted as an instrument. Therefore, one may use an alternative instrument, as in shephard2020, and construct a Wald-type test accordingly. We leave a systematic comparison between the Bonferroni-type multiple-testing procedure and the Wald-type joint test for future research.
4.2 Predictive Regressions with an Intercept
The analysis in Section 3 assumes that the intercept $\alpha=0$ in (1). When $\alpha\neq 0$, it must be properly accounted for. A natural starting point is the demeaned model
$y_{t}-\bar{y}_{T}=\beta(x_{t-1}-\bar{x}_{T})+u_{t}-\bar{u}_{T},\qquad t=1,\ldots,T,$ (8)
where $\bar{z}_{s}=s^{-1}\sum_{t=1}^{s}z_{t}$ for $z_{t}\in\{y_{t},x_{t-1},u_{t}\}$. However, $(u_{t}-\bar{u}_{T})$ is not an martingale difference sequence (MDS) with respect to $(\mathcal{F}_{t})$, invalidating the martingale CLT used in Sections 2 and 3. Specifically, the Cauchy estimator becomes
$\check{\beta}-\beta=\Biggl(\sum_{t=1}^{T}|x_{t-1}-\bar{x}_{T}|\Biggr)^{-1}\sum_{t=1}^{T}\text{sign}(x_{t-1}-\bar{x}_{T})(u_{t}-\bar{u}_{T}),$
which is problematic because: (i) $u_{t}-\bar{u}_{T}$ is not an MDS, and (ii) $\text{sign}(x_{t-1}-\bar{x}_{T})$ is not $\mathcal{F}_{t-1}$-measurable. Thus, the theory in Section 3 is not directly applicable. [^13] Recursive demeaning using $\bar{y}_{t}$ instead of $\bar{y}_{T}$ does not resolve this issue since $u_{t}-\bar{u}_{t}$ is not an MDS either.
To restore the MDS property, we instead difference the model:
$y_{t}-y_{t-1}=\beta(x_{t-1}-x_{t-2})+(u_{t}-u_{t-1}),$
and estimate this first-differenced (FD) model on alternating subsets of observations. We focus on the even-indexed observations and define the modified Cauchy estimator:
$\check{\beta}_{e}=(D_{T}^{e})^{-1}\sum_{t=2}^{T/2}\text{sign}(x_{2t-2})(y_{2t}-y_{2t-1}),\quad D_{T}^{e}=\sum_{t=2}^{T/2}\text{sign}(x_{2t-2})(x_{2t-1}-x_{2t-2}).$
This estimator has two key properties. First, for even-indexed data, the regression error $u_{t}^{e}=u_{2t}-u_{2t-1}$ forms an MDS with respect to $\mathcal{F}_{t}^{e}:=\mathcal{F}_{2t}$ for $t=1,\ldots,T/2$. [^14] For odd-indexed data, $u_{t}^{o}=u_{2t+1}-u_{2t}$ forms an MDS with respect to $\mathcal{F}_{t}^{o}:=\mathcal{F}_{2t+1}$, yielding an analogous estimator $\check{\beta}_{o}$. Second, $\check{\beta}_{e}$ can again be viewed as an IV estimator, but it uses $\text{sign}(x_{t-2})$, which is $\mathcal{F}_{t-2}$-measurable, as the instrument. [^15] More generally, one may use $\text{sign}\!\left(\sum_{l\leq 2}w_{l}x_{t-l}\right)$ for deterministic weights $\{w_{l}\}$, provided $\mathbb{E}[\text{sign}(\sum_{l\leq 2}w_{l}x_{t-l})(x_{t-1}-x_{t-2})]\neq 0$. Hence, $\text{sign}(x_{2t-2})(u_{2t}-u_{2t-1})$ is an MDS with respect to $(\mathcal{F}_{t}^{e})$.
The inference procedures of Section 3 remain valid for $\check{\beta}_{e}$. In particular, the hybrid test in Section 3.2 can be implemented as
$\tau(\check{\beta}_{e})=\frac{\check{\gamma}_{e}}{\hat{\omega}}$, (9)
where $\check{\gamma}_{e}=D_{T}^{e}\check{\beta}_{e}/\sqrt{T/2}$, and
$\hat{\omega}^{2}=\frac{1}{T}\sum_{t=1}^{T}\hat{u}_{t}^{2},\qquad\hat{u}_{t}=(y_{t}-\bar{y}_{T})-\hat{\beta}(x_{t-1}-\bar{x}_{T}),$
with $\hat{\beta}$ the OLS estimator from the demeaned model (8). Note that $\hat{\omega}^{2}$ is based on the full sample, whereas $\check{\beta}_{e}$ uses only even-indexed data. The asymptotic validity of this hybrid procedure is established next.
Corollary 4.2.
Let Assumptions 2.1, 2.2, and 2.3 hold, and suppose Assumption 3.1 holds with $x_{t-1}$ replaced by $x_{t-1}-\bar{x}_{T}$.
(a) Under $H_{0}:\beta=0$,
$\tau(\check{\beta}_{e})\to_{d}\mathbb{N}(0,1).$
(b) Under $H_{A}:\beta\neq 0$,
$\tau(\check{\beta}_{e})=\beta\,\frac{\sum_{t=1}^{T/2}\text{sign}(x_{2t-2})(x_{2t-1}-x_{2t-2})}{\omega\sqrt{T/2}}(1+o_{p}(1))+O_{p}(1),$
so that $|\tau(\check{\beta}_{e})|\to_{p}\infty$ whenever $\bigl|\sum_{t=1}^{T/2}\text{sign}(x_{2t-2})(x_{2t-1}-x_{2t-2})\bigr|/\sqrt{T/2}\to_{p}\infty$.
Although the odd-indexed estimator $\check{\beta}_{o}$ has analogous properties, $\check{\beta}_{e}$ and $\check{\beta}_{o}$ are typically dependent, with the dependence structure determined by the DGP of $(x_{t})$. Hence, unless additional assumptions are imposed, we restrict attention to a single subset of observations—either with even or odd indices. [^16] Using only half of the data is not uncommon in predictive regressions. See, for example, ZhuCaiPeng2014 and Liu2019, who employ long-lag differencing to eliminate intercepts. In addition, Dufour uses a split-sample approach to address inference problems under a Markovian structure.
Consistency of the hybrid test with an intercept requires
$\frac{1}{\sqrt{T/2}}\sum_{t=1}^{T/2}\operatorname{sign}(x_{2t-2})(x_{2t-1}-x_{2t-2})\to_{p}\infty.$
This holds for most stationary processes $(x_{t})$ if
$\mathbb{E}[\operatorname{sign}(x_{t-1})x_{t}]\neq\mathbb{E}[|x_{t-1}|].$
The condition may fail for certain unit-root processes. For instance, for a random walk $x_{t}=x_{t-1}+\varepsilon_{t}^{x}$, it does not hold. More generally, in the local-to-unity model of Phillips_Magdalinos_2007book,
$x_{t}=\left(1+\frac{c}{T^{\delta}}\right)x_{t-1}+\varepsilon_{t}^{x},\qquad c<0,\quad\delta\in[0,1],$ (10)
with $\varepsilon_{t}^{x}$ satisfying Assumption LP therein, the consistency condition becomes
$\frac{1}{T^{1/2}}\sum_{t=1}^{T/2}\operatorname{sign}(x_{2t-2})(x_{2t-1}-x_{2t-2})$ $=\frac{c}{T^{1/2+\delta}}\sum_{t=1}^{T/2}|x_{2t-2}|+\frac{1}{T^{1/2}}\sum_{t=1}^{T/2}\operatorname{sign}(x_{2t-2})\varepsilon_{2t-1}^{x}.$
The first term diverges if and only if $\delta<1$ (see Lemma 3.2 of Phillips_Magdalinos_2007book, Phillips_Magdalinos_2007book). Since $\varepsilon_{t}^{x}$ and $x_{t-1}$ may be dependent, typically $\mathbb{E}[\operatorname{sign}(x_{2t-2})\varepsilon_{2t-1}^{x}]\neq 0$, causing the second term to diverge as well. Hence, the condition fails only if $\delta=1$ and $\mathbb{E}[\operatorname{sign}(x_{2t-2})\varepsilon_{2t-1}^{x}]=0$. In all other cases ($0\leq\delta<1$ or nonzero covariance), the test remains consistent.
5 Finite Sample Performance
This section investigates the finite-sample performance of the proposed inference methods. Two sets of simulation experiments are conducted. The first set is based on a continuous-time model and compares our robust $t$-statistic–based tests, $t_{q}(\check{\gamma})$ for $q\in\{8,12,16\}$, and the hybrid test $\tau(\check{\beta})$ with the Cauchy RT test of CJP2016 and the Cauchy VC test of ibragimov2023new. The second set is based on a discrete-time predictive regression model with an intercept and compares our procedures with the IVX test of KMS2015.
5.1 Continuous-Time Experiments
5.1.1 Simulation Design
Following CJP2016 and ibragimov2023new, we consider the continuous-time predictive regression model
$dY_{t}$ $=\beta X_{t}\,dt+dU_{t},$ (11)
$dX_{t}$ $=-\frac{\bar{\kappa}}{T}X_{t}\,dt+\sigma_{t}\,dV_{t},\qquad dU_{t}=\sigma_{t}\left(dW_{t}+\int_{\mathbb{R}}x\,\Lambda(dt,dx)\right),$
where $V_{t}$ and $W_{t}$ are Brownian motions with $\mathbb{E}(V_{t}W_{t})=-0.98t$. The constant term in the predictive regression is set to zero without loss of generality, and recursive demeaning is applied. The model is observed at interval $\Delta=1/252$, corresponding to daily observations, so that a sample of length $T$ years contains $252T$ observations.
The volatility process $\sigma_{t}$ follows one of the following specifications:
  • CNST (Constant volatility): $\sigma_{t}^{2}=\sigma_{0}^{2}$, with $\sigma_{0}=1$.
  • SB (Structural break): $\sigma_{t}=\sigma_{0}+(\sigma_{1}-\sigma_{0})1\{t/T\geq 4/5\}$, with $\sigma_{0}=1$ and $\sigma_{1}=4$.
  • RS (Regime switching): $\sigma_{t}=\sigma_{0}(1-s_{t})+\sigma_{1}s_{t}$, where $s_{t}$ is a two-state Markov chain independent of $(Y_{t},X_{t})$, with transition matrix
    $P_{t}=\begin{pmatrix}0.8&0.2\\ 0.8&0.2\end{pmatrix}+\begin{pmatrix}0.2&-0.2\\ -0.8&0.8\end{pmatrix}\exp\!\left(-\frac{\bar{\lambda}}{T}t\right),$
    initialized at its invariant distribution, where $\bar{\lambda}=60$, $\sigma_{0}=1$, and $\sigma_{1}=4$.
  • GBM (Geometric Brownian motion): $d\sigma_{t}^{2}=\frac{1}{2}\frac{\bar{\omega}^{2}}{T}\sigma_{t}^{2}\,dt+\frac{\bar{\omega}^{2}}{\sqrt{T}}\sigma_{t}^{2}\,dZ_{t}$, where $Z_{t}$ is a Brownian motion correlated with $W_{t}$ such that $\mathbb{E}(W_{t}Z_{t})=-0.4t$ and $\bar{\omega}=9$.
We set $T\in\{5,20,50,100\}$ (corresponding to 60, 240, 600, and 1200 monthly observations) and $\bar{\kappa}\in\{0,5,10\}$ for the persistence parameter in (11), and consider a two-sided test of $H_{0}:\beta=0$ against $H_{A}:\beta\neq 0$.
5.1.2 Results
We first assess the empirical size of each test under the null hypothesis $\beta=0$. The results for the four volatility models (CNST, SB, RS, and GBM) and for $T\in\{5,20,50\}$ are reported in Table 1. Overall, both the $t$-statistic–based tests and the hybrid method exhibit satisfactory size performance, closely matching the nominal levels and performing comparably to the Cauchy RT and Cauchy VC tests. Among the $t$-based procedures, moderate partition numbers ($q=12$ or $q=16$) provide the most stable results, whereas smaller $q$ values tend to be mildly undersized. In the GBM case, where volatility is endogenously persistent, the $t$-statistic–based tests become slightly conservative but remain competitive with the Cauchy RT and VC methods.
Next, we analyze the finite-sample power properties of the tests. We consider $\beta\in\{0.004k,k=1,\cdots,5\}$ and $T\in\{20,50,100\}$ under the same volatility specifications. The results are summarized in Tables 25. The proposed tests exhibit power comparable to that of the Cauchy RT and Cauchy VC procedures. For small samples ($T=20$), the Cauchy RT and VC tests occasionally show higher power, but the difference diminishes as $T$ increases. In certain settings, our methods even outperform the existing approaches. For instance, $t_{16}(\check{\gamma})$ dominates under $\beta=0.02$, $\bar{\kappa}=0.5$, and regime-switching volatility (Table 4), whereas the hybrid test $\tau(\check{\beta})$ performs best under $\beta=0.004$, $\bar{\kappa}=20$, $T=20$, and regime-switching volatility.
In summary, all four robust inference procedures—Cauchy RT, Cauchy VC, $t_{q}(\check{\gamma})$, and $\tau(\check{\beta})$—deliver accurate size control and strong discriminatory power under endogenously persistent regressors and persistent volatility. While the Cauchy RT requires high-frequency data and a time transformation, and the Cauchy VC involves nonparametric volatility filtering with a tuning parameter, our proposed $t$-statistic and hybrid methods are much simpler to implement and require neither. Hence, these approaches are best viewed as complementary: the Cauchy RT and Cauchy VC are preferable in high-frequency environments, whereas our procedures provide robust and easily implementable alternatives in more general settings. It is also worth emphasizing that the proposed methods, like the Cauchy VC, are applicable to both continuous- and discrete-time models, whereas the Cauchy RT method is restricted to the continuous-time framework.
5.2 Discrete-Time Experiments
5.2.1 Simulation Design
We now examine the finite-sample performance of the proposed tests in a discrete-time setting with an intercept, comparing them to the IVX test of KMS2015. The data-generating process (DGP) is specified as
$y_{t}$ $=\beta x_{t-1}+\sigma_{t}\varepsilon_{t},$
$x_{t}$ $=\left(1-\frac{\bar{\kappa}}{T}\right)x_{t-1}+\sigma_{t}\eta_{t},$ (12)
for $t=2,\ldots,T$, where $T\in\{60,240,600\}$ corresponds to 5, 20, and 50 years of monthly data. We set $\beta\in\{0.5k:k=0,1,\ldots,5\}$ and $\bar{\kappa}\in\{0,50,100\}$, and consider a one-sided test of $H_{0}\!:\beta=0$ against $H_{A}\!:\beta>0$. [^17] The IVX test of KMS2015 performs well in two-sided testing for a broad class of models. However, as shown in DEMETRESCU2023105271, the IVX method exhibits severe size distortions in one-sided tests when regressors are highly persistent and endogenous. For this reason, we focus on the one-sided case to demonstrate the performance of our methods in this setting.
The innovation process $\eta_{t}$ follows an MA($q$) process:
$\eta_{t}=\sum_{j=1}^{q}C_{j}v_{t-j},$
where $(v_{t},\eta_{t})$ are jointly normal with correlation 0.98-0.98. For the MA(2) case, $C_{1}=C_{2}=1/\sqrt{2}$ and $C_{q}=0$ for $q>3$; for the MA(4) case, $C_{1}=C_{2}=C_{3}=C_{4}=1/2$ and $C_{q}=0$ for $q>5$. The volatility process $\sigma_{t}$ follows the same specifications as in the continuous-time simulations, except that the GBM model is excluded.
We implement the hybrid tests based on the even and odd observations, denoted by $\tau(\check{\beta}_{e})$ and $\tau(\check{\beta}_{o})$, respectively (see (9)), along with their $t$-statistic versions (see (4)). For comparison, we also include the IVX test of KMS2015.
5.2.2 Results
The results, summarized in Tables 611, indicate that the proposed tests exhibit excellent size control under the null hypothesis across all DGPs, whereas the IVX test is substantially oversized, particularly when volatility is nonstationary or exhibits structural breaks. Furthermore, both the hybrid and $t$-statistic approaches demonstrate nontrivial power, even though they are constructed using only half of the observations. Among the size-controlled procedures, the statistic $t_{16}(\tau(\check{\beta}_{o}))$ consistently delivers the strongest performance.
Overall, these findings corroborate the theoretical robustness of our methods. They remain valid under heavy-tailed, endogenous, and persistent regressors, as well as under heteroskedastic and persistent volatility. In contrast, the IVX test performs well only under stationary volatility and light tails. Hence, our proposed robust procedures offer a practical and reliable alternative to existing inference methods for predictive regressions in both continuous- and discrete-time frameworks.
6 Empirical Application
To illustrate the empirical performance of the proposed tests relative to the Cauchy RT and Cauchy VC tests, we reexamine the dataset used by CJP2016 to test the predictability of stock returns using the dividend–price (D/P) and earnings–price (E/P) ratios as predictors. For stock returns, we employ the NYSE/AMEX value-weighted index and the S&P 500 index obtained from the Center for Research in Security Prices (CRSP). The dividend–price ratio is defined as the annual dividend divided by the current total market value. Further details on data construction are provided in Section 6.1 of CJP2016.
Following CJP2016, we estimate two types of predictive regressions: one based on all returns and another based only on returns generated from the diffusive component of stock prices, obtained by first testing for jumps and removing observations corresponding to detected jumps. In all cases, we apply one-sided tests.
[…]
Consistent with our simulation evidence, the Cauchy RT test demonstrates strong finite-sample power but requires high-frequency data due to its reliance on a continuous-time approximation. [^18] For the Cauchy RT test in our simulations, we estimate the discretized time-changed regression using $n$ lower-frequency observations, with $n=12T$ (approximately monthly), generated by the random time-sampling scheme described in Section 5 of CJP2016. The mixed empirical results—where the Cauchy RT test fails to reject the null while both the proposed methods and the Cauchy VC test do reject—may reflect the limited accuracy of the continuous-time approximation when applied to monthly, quarterly, or yearly data. Evaluating the robustness of the continuous-time approximation underlying the Cauchy RT test remains an interesting topic for future research.
7 Conclusion
This paper introduces two robust inference methods for predictive regressions, addressing key econometric challenges commonly encountered in empirical finance, such as endogenously persistent or heavy-tailed regressors and persistent volatility in errors. Building on the Cauchy estimation framework, we develop two simple yet theoretically rigorous procedures: a $t$-statistic–based approach and a hybrid method. Both methods are computationally straightforward and applicable to continuous- and discrete-time models alike.
Simulation evidence demonstrates that the proposed tests perform well in finite samples, maintaining correct size and competitive power under a wide range of data-generating processes, including those characterized by stochastic volatility, structural breaks, and regime switching. Although our procedures require the assumption of asymptotically exogenous volatility, they exhibit excellent robustness and complement existing Cauchy-based methods, including the IVX method of KMS2015, the Cauchy RT test of CJP2016 and the Cauchy VC test of ibragimov2023new.
In an empirical application to stock return predictability, we use the dividend–price and earnings–price ratios as predictors for excess returns on major U.S. equity indices. The results indicate that the dividend–price ratio possesses predictive power, while the earnings–price ratio does not significantly forecast returns. Overall, the proposed inference procedures offer a practical, theoretically sound, and implementable alternative to existing methods for robust inference in predictive regressions.
Appendix: Proofs
Proof of Lemma 3.1.
For $j=1,\dots,q$, we have
$\check{\gamma}_{j}-\zeta_{j}=\sqrt{\frac{q}{T}}\sum_{t=(j-1)[T/q]+1}^{j[T/q]}\operatorname{sign}(x_{t-1})u_{t}=\sqrt{q}\int_{(j-1)/q}^{j/q}\sigma^{T}(r)\,dW^{T}(r).$
The stated result follows immediately from Assumptions 2.2 and 2.3. ∎
Proof of Proposition 3.2.
Part (a) follows from Theorem 1 and the discussion in Section 2.2 of IbragimovMuller2010. For part (b), we deduce from Lemma 3.1 that
$\frac{1}{\sqrt{T/q}}\check{\gamma}_{j}=\frac{1}{\sqrt{T/q}}\zeta_{j}+o_{p}\!\left(\frac{1}{\sqrt{T}}\right)\to_{p}\beta\,\mathbb{E}|x_{t}|$ (A.2)
uniformly in $j$, under $\beta\neq 0$. Recall that
$t_{q}(\check{\gamma})=\frac{\sqrt{q}\,\overline{\gamma}}{s_{\gamma}},\quad\text{with }\overline{\gamma}=\frac{1}{q}\sum_{j=1}^{q}\check{\gamma}_{j},\quad s_{\gamma}^{2}=\frac{1}{q-1}\sum_{j=1}^{q}(\check{\gamma}_{j}-\overline{\gamma})^{2}.$
Hence, the numerator of $t_{q}(\check{\gamma})$ satisfies
$\frac{\sqrt{q}\,\overline{\gamma}}{\sqrt{T}}\to_{p}\beta\,\mathbb{E}|x_{t}|.$ (A.3)
To complete the proof, it suffices to show that $s_{\gamma}^{2}=o_{p}(T)$. Indeed, for $q\geq 2$, we have
$\frac{q(q-1)}{T}s_{\gamma}^{2}=\frac{1}{T/q}\sum_{j=1}^{q}(\check{\gamma}_{j}-\overline{\gamma})^{2}\to_{p}0$
due to (A.2) and (A.3), which completes the proof. ∎
Proof of Corollary 3.3.
We aim to show that
$\frac{\sqrt{q}}{T}\overline{\gamma}\to_{d}\beta\int_{0}^{1}|X(r)|\,dr,$ (A.4)
and
$\frac{q(q-1)}{T^{2}}s_{\gamma}^{2}=\frac{q}{T^{2}}\sum_{j=1}^{q}(\check{\gamma}_{j}-\overline{\gamma})^{2}\to_{d}\beta^{2}\sum_{j=1}^{q}\left(q\int_{(j-1)/q}^{j/q}|X(r)|\,dr-\int_{0}^{1}|X(r)|\,dr\right)^{2}.$ (A.5)
For (A.4), we have
$\frac{\sqrt{q}}{T}\check{\gamma}_{j}$ $=\frac{\sqrt{q}}{T}\zeta_{j}+o_{p}\!\left(\frac{1}{T}\right)=\beta\frac{q}{T^{3/2}}\sum_{t=(j-1)[T/q]+1}^{j[T/q]}|x_{t-1}|+o_{p}(1)\to_{d}\beta q\int_{(j-1)/q}^{j/q}|X(r)|\,dr,$
by Lemma 3.1 and (5), leading to (A.4). Moreover,
$\frac{q}{T^{2}}(\check{\gamma}_{j}-\overline{\gamma})^{2}\to_{d}\beta^{2}\left(q\int_{(j-1)/q}^{j/q}|X(r)|\,dr-\int_{0}^{1}|X(r)|\,dr\right)^{2},$
which yields (A.5).
Combining (A.4) and (A.5), we obtain
$t_{q}(\check{\gamma})=\frac{\sqrt{q}\,\overline{\gamma}}{s_{\gamma}}\to_{d}\operatorname{sign}(\beta)\mathcal{D}_{q},$
where
$\mathcal{D}_{q}=\int_{0}^{1}|X(r)|\,dr\left(\frac{q(q-1)}{\sum_{j=1}^{q}(\int_{0}^{1}|X(r)|\,dr-q\int_{(j-1)/q}^{j/q}|X(r)|\,dr)^{2}}\right)^{1/2},$
as desired. Furthermore, $(q-1)^{1/2}\mathcal{D}_{q}>1$ for $q\geq 2$ with probability one, since
$(q-1)^{1/2}\mathcal{D}_{q}$ $=(q-1)\int_{0}^{1}|X(r)|\,dr\left(\frac{q}{\sum_{j=1}^{q}(\int_{0}^{1}|X(r)|\,dr-q\int_{(j-1)/q}^{j/q}|X(r)|\,dr)^{2}}\right)^{1/2}$
$>\frac{(q-1)\int_{0}^{1}|X(r)|\,dr}{\max_{1\leq j\leq q}\left|\int_{0}^{1}|X(r)|\,dr-q\int_{(j-1)/q}^{j/q}|X(r)|\,dr\right|}>1,$
which completes the proof. ∎
Proof of Proposition 3.4.
It suffices to show that $\hat{\omega}^{2}\to_{p}\omega^{2}$, since this implies
$\tau(\check{\beta})=\frac{\check{\gamma}}{\omega}(1+o_{p}(1))=\left(\beta\frac{\sum_{t=1}^{T}|x_{t-1}|}{\omega\sqrt{T}}+\frac{\sum_{t=1}^{T}\operatorname{sign}(x_{t-1})u_{t}}{\omega\sqrt{T}}\right)(1+o_{p}(1)),$
where, in particular,
$\frac{\sum_{t=1}^{T}\operatorname{sign}(x_{t-1})u_{t}}{\omega\sqrt{T}}\to_{d}\mathbb{N}(0,1)$
under Assumptions 2.1, 2.2, and 2.3.
Let $\omega_{T}^{2}=T^{-1}\sum_{t=1}^{T}u_{t}^{2}$. Then $\omega_{T}^{2}\to_{p}\omega^{2}$ by Assumptions 2.1 and $2.$ Furthermore,
$\hat{\omega}^{2}=\omega_{T}^{2}-\frac{1}{T}\frac{(\sum_{t=1}^{T}x_{t-1}u_{t})^{2}}{\sum_{t=1}^{T}x_{t-1}^{2}}=\omega_{T}^{2}+o_{p}(1)$
by Assumption 3.1, which gives the desired result. ∎
Proof of Lemma 4.1.
We need only show that $H_{0}$ holds if $H_{0}^{(k)}$ holds for all $k$. Let $C=[\beta_{1},\ldots,\beta_{K}]^{\prime}$. By the moment restrictions, $C$ is the solution to
$\mathbb{E}[z_{t-1}y_{t}]=\mathbb{E}\!\left[\operatorname{diag}(|x_{1,t-1}|,\ldots,|x_{K,t-1}|)\right]C.$
Given $0<\mathbb{E}[|x_{k,t-1}|]<\infty$ for all $k$, $H_{0}^{(k)}$ holds for all $k$ if and only if $\mathbb{E}[z_{t-1}y_{t}]=\mathbf{0}_{K\times 1}$.
Moreover,
$\mathbb{E}[z_{t-1}y_{t}]=\mathbb{E}[z_{t-1}x_{t-1}^{\prime}]B,$
and since $\mathbb{E}[z_{t-1}x_{t-1}^{\prime}]$ is assumed invertible, the condition $\mathbb{E}[z_{t-1}y_{t}]=\mathbf{0}_{K\times 1}$ implies $B=0$, completing the proof. ∎
Proof of Corollary 4.2.
Under the modified version of Assumption 3.1 stated in the corollary, it follows from arguments analogous to those in the proof of Proposition 3.4 that $\hat{\omega}^{2}\to_{p}\omega^{2}$. Hence, it suffices to show that
$\frac{1}{\omega\sqrt{T}}\sum_{t=1}^{T/2}\operatorname{sign}(x_{2t-2})(u_{2t}-u_{2t-1})\to_{d}\mathbb{N}(0,1).$
Define $\xi_{t}=\operatorname{sign}(x_{t-2})u_{t}$ for even $t$ and $\xi_{t}=-\operatorname{sign}(x_{t-1})u_{t}$ for odd $t$, so that
$\frac{1}{\omega\sqrt{T}}\sum_{t=1}^{T/2}\operatorname{sign}(x_{2t-2})(u_{2t}-u_{2t-1})=\frac{1}{\omega\sqrt{T}}\sum_{t=1}^{T}\xi_{t}.$
By construction, $(\xi_{t})$ is an MDS with respect to $(\mathcal{F}_{t})$ and satisfies $\mathbb{E}(\xi_{t}^{2}\mid\mathcal{F}_{t-1})=v_{t}^{2}$ under Assumption 2.1. The desired convergence then follows directly from the martingale central limit theorem, given Assumptions 2.2 and 2.3. ∎
Table 1: Size for the continuous time models
$\bar{\kappa}=0$ $\bar{\kappa}=5$ $\bar{\kappa}=20$
T 5 20 50 5 20 50 5 20 50
CNST Cauchy RT 5.3 4.9 5.3 5.2 5.4 4.7 5.5 5.1 5.1
Cauchy VC 5.6 5.0 5.3 5.4 5.0 5.1 5.4 5.0 4.8
$t_{8}(\check{\gamma})$ 4.8 4.6 5.2 4.7 5.0 5.1 4.8 4.9 5.0
$t_{12}(\check{\gamma})$ 4.7 4.7 5.0 4.7 5.1 4.9 5.0 5.2 4.9
$t_{16}(\check{\gamma})$ 4.8 4.8 5.1 4.9 5.0 4.9 4.7 4.9 4.9
$\tau({\check{\beta}})$ 5.2 4.8 5.2 5.0 4.8 5.1 5.2 4.9 4.8
SB Cauchy RT 5.6 5.0 5.1 5.2 5.3 5.0 5.4 5.0 4.9
Cauchy VC 8.0 6.7 6.3 7.8 6.5 6.0 7.9 6.4 6.0
$t_{8}(\check{\gamma})$ 3.6 3.7 3.9 4.0 4.1 3.7 3.8 3.7 3.7
$t_{12}(\check{\gamma})$ 4.7 4.2 4.6 4.3 4.6 4.5 4.2 4.2 4.5
$t_{16}(\check{\gamma})$ 4.7 4.6 4.6 4.8 4.7 4.5 4.3 4.6 4.5
$\tau({\check{\beta}})$ 6.1 5.0 4.9 6.1 5.5 5.1 6.2 5.5 5.0
RS Cauchy RT 5.0 4.8 5.2 4.9 4.9 4.9 5.4 5.1 4.8
Cauchy VC 5.2 5.4 6.1 5.2 5.1 5.8 5.6 5.8 5.8
$t_{8}(\check{\gamma})$ 4.5 4.5 5.1 4.6 4.4 4.9 4.6 5.3 4.6
$t_{12}(\check{\gamma})$ 4.8 4.6 5.0 4.5 4.8 4.8 4.6 4.9 4.7
$t_{16}(\check{\gamma})$ 5.0 4.4 4.9 4.6 4.6 4.9 5.1 5.1 4.5
$\tau({\check{\beta}})$ 5.0 4.9 5.2 5.2 4.7 4.9 5.2 5.3 5.0
GBM Cauchy RT 4.4 4.7 4.4 4.3 4.5 4.4 4.6 4.5 4.5
Cauchy VC 5.4 5.5 6.1 5.7 5.7 5.9 5.7 5.9 6.5
$t_{8}(\check{\gamma})$ 3.2 3.1 3.2 3.6 3.8 3.4 3.6 3.8 3.9
$t_{12}(\check{\gamma})$ 3.8 3.6 3.7 4.2 4.3 3.6 4.2 4.2 4.3
$t_{16}(\check{\gamma})$ 3.7 3.8 4.1 4.4 4.2 4.0 4.6 4.5 4.6
$\tau({\check{\beta}})$ 5.3 4.6 5.0 5.2 4.8 4.9 5.6 4.9 5.2
CNST, SB, GBM, and RS denote constant volatility, structural break, geometric Brownian motion, and regime switching, respectively.
Table 2: Power for the continuous time models, constant volatility
$\bar{\kappa}=0$ $\bar{\kappa}=5$ $\bar{\kappa}=20$
T 20 50 100 20 50 100 20 50 100
$\beta=0.004$ Cauchy RT 8.8 25.2 85.4 6.1 9.9 22.7 6.3 13.2 27.9
Cauchy VC 9.1 25.6 86.2 6.5 10.0 23.5 8.2 13.4 26.8
$t_{8}(\check{\gamma})$ 8.0 21.0 79.5 6.7 11.2 24.6 5.1 14.6 30.6
$t_{12}(\check{\gamma})$ 9.5 21.4 82.0 6.2 10.5 25.4 5.9 14.3 31.0
$t_{16}(\check{\gamma})$ 8.8 22.8 84.0 5.5 10.0 23.3 5.5 14.5 29.9
$\tau(\check{\beta})$ 6.4 23.5 66.6 5.3 8.6 17.5 5.9 12.9 24.3
$\beta=0.008$ Cauchy RT 15.5 85.5 100.0 9.2 22.1 83.9 6.6 27.4 85.6
Cauchy VC 17.2 86.2 100.0 8.8 23.1 84.3 11.9 27.7 84.2
$t_{8}(\check{\gamma})$ 13.8 79.9 100.0 9.1 23.5 86.7 6.0 32.5 89.2
$t_{12}(\check{\gamma})$ 15.0 83.4 100.0 8.9 25.5 86.8 6.5 31.7 92.0
$t_{16}(\check{\gamma})$ 15.0 83.7 100.0 8.3 23.7 87.0 6.1 31.9 91.7
$\tau(\check{\beta})$ 14.0 66.7 96.6 7.3 18.8 57.6 8.5 27.6 70.0
$\beta=0.012$ Cauchy RT 37.3 99.2 100.0 12.2 50.1 100.0 7.7 52.2 100.0
Cauchy VC 40.2 99.3 100.0 12.6 51.9 100.0 16.5 56.3 100.0
$t_{8}(\check{\gamma})$ 30.3 98.0 100.0 12.9 55.5 99.8 6.9 60.6 100.0
$t_{12}(\check{\gamma})$ 33.1 99.5 100.0 12.0 56.3 100.0 7.2 61.6 100.0
$t_{16}(\check{\gamma})$ 35.3 98.9 100.0 12.3 54.1 100.0 6.6 59.2 100.0
$\tau(\check{\beta})$ 32.4 87.7 99.1 9.7 34.0 93.5 12.1 48.9 99.0
$\beta=0.016$ Cauchy RT 67.0 100.0 100.0 17.0 84.4 100.0 8.4 83.2 100.0
Cauchy VC 68.8 100.0 100.0 16.6 84.7 100.0 22.4 83.3 100.0
$t_{8}(\check{\gamma})$ 58.8 99.7 100.0 18.1 88.7 100.0 7.3 88.7 100.0
$t_{12}(\check{\gamma})$ 60.7 100.0 100.0 16.8 86.6 100.0 7.6 89.7 100.0
$t_{16}(\check{\gamma})$ 62.4 99.9 100.0 16.7 86.9 100.0 7.0 89.2 100.0
$\tau(\check{\beta})$ 51.2 95.8 99.9 12.4 58.5 99.0 17.8 72.9 100.0
$\beta=0.02$ Cauchy RT 86.7 100.0 100.0 23.2 98.0 100.0 9.3 98.2 100.0
Cauchy VC 87.9 100.0 100.0 24.4 98.5 100.0 29.4 97.4 100.0
$t_{8}(\check{\gamma})$ 79.3 100.0 100.0 24.6 97.8 100.0 8.5 98.3 100.0
$t_{12}(\check{\gamma})$ 81.9 100.0 100.0 23.8 99.1 100.0 8.5 99.1 100.0
$t_{16}(\check{\gamma})$ 82.3 100.0 100.0 23.2 98.3 100.0 7.7 99.2 100.0
$\tau(\check{\beta})$ 65.2 97.5 100.0 16.3 81.3 99.6 23.8 92.3 100.0
Table 3: Power for the continuous time models, structural break in volatility
$\bar{\kappa}=0$ $\bar{\kappa}=5$ $\bar{\kappa}=20$
T 20 50 100 20 50 100 20 50 100
$\beta=0.004$ Cauchy RT 8.7 20.0 67.7 8.3 11.6 35.9 8.7 17.2 52.4
Cauchy VC 8.5 20.6 57.6 8.9 10.5 19.2 8.2 14.6 25.3
$t_{8}(\check{\gamma})$ 5.6 11.3 45.6 4.4 8.7 15.7 6.4 8.4 23.3
$t_{12}(\check{\gamma})$ 6.6 14.8 45.3 6.7 9.1 17.5 6.2 10.5 22.7
$t_{16}(\check{\gamma})$ 8.2 14.6 44.1 6.9 8.9 19.5 6.6 10.1 22.9
$\tau(\check{\beta})$ 5.8 13.7 31.4 7.1 8.1 13.5 5.8 10.2 16.7
$\beta=0.008$ Cauchy RT 14.4 65.1 96.6 11.7 35.7 88.4 14.5 52.2 99.6
Cauchy VC 14.9 60.6 95.6 10.8 21.6 55.8 11.5 30.1 72.9
$t_{8}(\check{\gamma})$ 9.5 44.1 88.5 6.8 16.3 57.4 9.0 21.0 68.3
$t_{12}(\check{\gamma})$ 11.4 45.6 89.8 8.9 17.9 55.9 9.2 23.6 69.4
$t_{16}(\check{\gamma})$ 12.7 42.3 92.1 8.4 18.3 55.9 8.8 22.2 72.2
$\tau(\check{\beta})$ 8.9 32.4 81.9 8.9 14.0 34.3 7.9 16.6 49.7
$\beta=0.012$ Cauchy RT 32.1 87.0 99.3 15.5 68.6 98.8 24.2 90.1 100.0
Cauchy VC 26.1 83.1 99.2 13.5 36.0 91.1 15.5 50.3 98.8
$t_{8}(\check{\gamma})$ 16.9 72.9 97.0 9.2 32.7 89.0 12.0 43.1 95.2
$t_{12}(\check{\gamma})$ 20.7 74.5 99.1 11.9 34.5 90.0 12.2 41.7 98.0
$t_{16}(\check{\gamma})$ 20.0 74.3 99.6 10.0 31.9 90.7 12.6 41.4 98.8
$\tau(\check{\beta})$ 14.7 60.2 98.1 10.8 23.4 73.7 10.8 33.2 92.1
$\beta=0.016$ Cauchy RT 50.2 95.6 99.7 22.8 89.8 100.0 37.7 99.7 100.0
Cauchy VC 42.9 95.1 99.6 17.0 58.9 99.5 20.7 74.5 100.0
$t_{8}(\check{\gamma})$ 30.2 88.8 98.5 12.6 56.2 96.1 16.6 67.4 99.8
$t_{12}(\check{\gamma})$ 31.7 90.4 99.7 14.7 53.8 98.4 17.0 65.0 100.0
$t_{16}(\check{\gamma})$ 31.8 91.1 100.0 13.0 54.4 99.0 18.5 66.6 100.0
$\tau(\check{\beta})$ 23.2 81.4 99.8 13.0 35.6 96.9 14.3 51.8 100.0
$\beta=0.02$ Cauchy RT 65.4 98.6 99.9 34.6 97.4 100.0 55.1 100.0 100.0
Cauchy VC 58.8 98.5 100.0 21.4 78.7 99.9 27.1 91.5 100.0
$t_{8}(\check{\gamma})$ 45.5 95.3 98.8 17.6 74.0 98.2 21.7 86.2 100.0
$t_{12}(\check{\gamma})$ 46.1 96.4 99.8 18.0 75.7 99.7 22.7 87.5 100.0
$t_{16}(\check{\gamma})$ 44.7 97.9 100.0 18.0 73.9 100.0 24.9 87.9 100.0
$\tau(\check{\beta})$ 32.2 92.4 100.0 15.7 56.6 99.8 18.2 73.4 100.0
Table 4: Power for the continuous time models, regime switching in volatility
$\bar{\kappa}=0$ $\bar{\kappa}=5$ $\bar{\kappa}=20$
T 20 50 100 20 50 100 20 50 100
$\beta=0.004$ Cauchy RT 6.6 23.8 64.2 5.9 10.3 21.0 8.1 12.8 32.9
Cauchy VC 11.0 49.9 85.9 7.8 12.9 33.0 7.4 13.0 34.5
$t_{8}(\check{\gamma})$ 7.7 22.1 77.7 6.1 11.8 22.2 7.5 10.4 28.4
$t_{12}(\check{\gamma})$ 8.0 23.2 83.6 6.2 10.9 22.6 7.7 11.0 28.9
$t_{16}(\check{\gamma})$ 7.9 22.8 84.0 6.5 11.8 22.9 7.8 10.6 30.8
$\tau(\check{\beta})$ 8.2 22.3 64.8 7.2 7.6 15.0 8.3 9.3 20.2
$\beta=0.008$ Cauchy RT 11.9 69.3 96.0 8.0 20.3 65.5 12.2 32.2 90.2
Cauchy VC 25.7 84.6 97.8 10.0 30.2 84.1 10.8 28.2 86.6
$t_{8}(\check{\gamma})$ 15.9 77.2 99.8 9.6 24.3 79.8 10.5 26.9 81.9
$t_{12}(\check{\gamma})$ 15.2 80.5 99.9 9.2 23.8 81.3 11.0 27.0 82.9
$t_{16}(\check{\gamma})$ 16.0 82.1 100.0 9.5 24.9 81.4 11.5 26.6 83.7
$\tau(\check{\beta})$ 16.7 67.7 94.4 9.1 13.8 50.9 10.3 19.2 60.0
$\beta=0.012$ Cauchy RT 31.6 87.0 98.9 10.7 39.8 95.5 17.9 63.4 99.5
Cauchy VC 47.3 94.5 98.5 12.1 58.3 95.4 14.9 53.5 98.7
$t_{8}(\check{\gamma})$ 33.0 96.7 100.0 13.9 50.8 99.1 15.4 51.7 99.4
$t_{12}(\check{\gamma})$ 35.1 98.2 100.0 13.3 52.2 99.5 15.1 55.5 99.8
$t_{16}(\check{\gamma})$ 34.8 98.3 100.0 13.2 51.2 99.1 15.4 53.5 99.8
$\tau(\check{\beta})$ 32.3 86.6 98.2 11.1 29.0 87.6 13.1 37.8 95.0
$\beta=0.016$ Cauchy RT 52.7 95.2 99.7 14.2 65.3 99.3 26.2 90.6 100.0
Cauchy VC 65.1 97.2 99.0 17.3 81.7 97.4 19.3 78.2 99.6
$t_{8}(\check{\gamma})$ 52.5 99.7 100.0 18.0 82.1 99.9 21.8 80.0 100.0
$t_{12}(\check{\gamma})$ 59.7 99.6 100.0 18.2 83.6 100.0 20.9 82.5 100.0
$t_{16}(\check{\gamma})$ 61.3 99.9 100.0 19.4 81.8 100.0 21.6 81.0 100.0
$\tau(\check{\beta})$ 51.6 94.3 98.9 13.5 54.9 96.6 17.4 59.5 99.5
$\beta=0.02$ Cauchy RT 66.4 97.7 100.0 20.9 86.1 99.9 36.2 98.2 100.0
Cauchy VC 76.8 98.0 99.2 21.5 91.1 98.7 25.0 92.1 99.7
$t_{8}(\check{\gamma})$ 74.3 100.0 100.0 23.7 95.4 99.9 27.8 94.7 100.0
$t_{12}(\check{\gamma})$ 80.4 99.9 100.0 23.6 96.3 100.0 28.4 95.8 100.0
$t_{16}(\check{\gamma})$ 81.1 100.0 100.0 25.7 96.8 100.0 29.1 96.3 100.0
$\tau(\check{\beta})$ 67.3 97.0 99.1 17.3 77.2 98.7 22.7 81.8 99.8
Table 5: Power for the continuous time models, geometric Brownian motion
$\bar{\kappa}=0$ $\bar{\kappa}=5$ $\bar{\kappa}=20$
T 20 50 100 20 50 100 20 50 100
$\beta=0.004$ Cauchy RT 7.8 19.2 59.9 6.0 9.6 24.6 8.5 14.0 41.8
Cauchy VC 42.4 59.7 80.4 6.7 13.0 29.5 7.2 10.8 19.6
$t_{8}(\check{\gamma})$ 11.0 27.1 67.7 4.4 7.5 21.4 5.7 8.4 18.8
$t_{12}(\check{\gamma})$ 10.3 28.5 72.0 5.4 8.8 21.7 4.9 8.1 19.2
$t_{16}(\check{\gamma})$ 10.2 29.2 73.2 5.5 9.3 21.2 5.8 9.2 19.9
$\tau(\check{\beta})$ 8.8 25.5 61.8 5.6 7.5 13.3 6.2 7.0 14.7
$\beta=0.008$ Cauchy RT 13.7 59.1 92.4 8.2 24.7 73.4 13.2 40.9 92.8
Cauchy VC 55.3 78.9 91.2 9.4 28.0 53.6 8.8 20.5 51.2
$t_{8}(\check{\gamma})$ 21.6 65.5 94.6 7.0 18.8 68.3 8.4 16.3 57.5
$t_{12}(\check{\gamma})$ 23.9 69.7 95.5 7.4 21.2 70.9 7.0 18.6 63.2
$t_{16}(\check{\gamma})$ 22.4 71.3 96.5 7.9 21.1 72.2 8.2 19.3 64.7
$\tau(\check{\beta})$ 21.2 59.2 88.0 6.7 12.0 35.0 8.0 14.5 40.5
$\beta=0.012$ Cauchy RT 29.0 80.8 97.9 11.2 48.3 95.0 18.7 72.2 99.2
Cauchy VC 65.6 88.3 92.2 13.8 42.9 65.6 12.1 36.1 69.9
$t_{8}(\check{\gamma})$ 38.9 85.5 97.3 10.7 43.0 89.6 10.4 35.2 81.2
$t_{12}(\check{\gamma})$ 40.2 88.4 98.5 11.0 43.6 93.0 10.5 39.8 88.0
$t_{16}(\check{\gamma})$ 42.0 89.0 99.1 10.0 44.4 92.8 11.1 39.6 91.5
$\tau(\check{\beta})$ 38.9 77.0 92.1 8.3 22.3 64.9 10.8 26.3 69.7
$\beta=0.016$ Cauchy RT 47.9 91.7 99.1 16.0 70.5 98.5 27.8 91.5 99.7
Cauchy VC 73.4 91.6 93.4 17.6 56.2 70.4 15.0 52.5 77.3
$t_{8}(\check{\gamma})$ 53.4 91.8 98.4 14.6 68.5 94.4 13.9 57.8 83.9
$t_{12}(\check{\gamma})$ 57.4 94.9 99.3 14.9 71.3 97.1 15.1 63.9 93.9
$t_{16}(\check{\gamma})$ 58.9 95.2 99.6 14.4 72.6 98.2 15.9 62.9 96.6
$\tau(\check{\beta})$ 51.2 86.3 93.5 10.3 37.4 74.8 13.1 42.1 82.3
$\beta=0.02$ Cauchy RT 60.1 96.6 99.5 23.7 87.6 98.9 40.3 97.9 99.9
Cauchy VC 79.7 92.6 93.9 22.3 64.9 72.9 19.4 63.6 80.7
$t_{8}(\check{\gamma})$ 67.5 94.8 99.0 20.2 81.4 96.0 18.3 71.0 85.5
$t_{12}(\check{\gamma})$ 69.9 97.2 99.7 19.7 85.7 98.6 20.0 80.7 95.0
$t_{16}(\check{\gamma})$ 72.9 97.2 99.9 20.2 87.5 99.1 21.2 79.9 98.2
$\tau(\check{\beta})$ 59.9 90.7 94.4 12.8 54.6 80.2 16.8 57.5 89.5
Table 6: Size and power for the discrete time models, constant volatility, MA(2) innovation
$\bar{\kappa}=0$ $\bar{\kappa}=50$ $\bar{\kappa}=100$
T 20 50 100 20 50 100 20 50 100
$\beta=0$ $\tau(\check{\beta}_{o})$ 5.0 4.9 4.8 5.2 5.3 4.8 4.5 4.7 5.2
$\tau(\check{\beta}_{e})$ 4.8 4.7 5.2 5.0 4.8 4.8 5.1 5.0 4.9
$t_{8}(\tau(\check{\beta}_{o}))$ 4.7 4.5 4.8 3.4 4.1 4.5 3.7 5.0 4.6
$t_{12}(\tau(\check{\beta}_{o}))$ 5.1 4.7 4.5 3.1 3.8 4.4 2.8 4.0 4.2
$t_{16}(\tau(\check{\beta}_{o}))$ 4.9 4.9 5.1 2.8 4.0 4.2 2.5 3.5 3.9
IVX 14.2 13.5 12.3 10.2 10.5 10.3 9.5 10.3 10.5
$\beta=0.5$ $\tau(\check{\beta}_{o})$ 8.4 7.9 8.4 17.0 17.5 16.6 24.9 24.9 25.8
$\tau(\check{\beta}_{e})$ 8.2 8.0 8.4 17.4 16.0 17.1 25.4 25.7 26.0
$t_{8}(\tau(\check{\beta}_{o}))$ 15.4 15.6 15.9 12.7 14.2 15.4 18.7 20.5 21.9
$t_{12}(\tau(\check{\beta}_{o}))$ 20.3 20.9 20.2 13.8 15.1 15.8 18.2 21.4 21.6
$t_{16}(\tau(\check{\beta}_{o}))$ 24.8 24.9 26.0 15.0 16.7 18.4 18.3 20.9 22.3
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1$ $\tau(\check{\beta}_{o})$ 13.4 13.4 13.9 39.3 41.1 39.7 60.5 61.1 62.0
$\tau(\check{\beta}_{e})$ 13.7 13.5 13.7 39.6 39.1 39.6 62.0 62.8 62.5
$t_{8}(\tau(\check{\beta}_{o}))$ 33.8 33.5 33.6 29.7 32.4 33.0 45.1 48.9 50.7
$t_{12}(\tau(\check{\beta}_{o}))$ 47.1 47.4 47.1 35.0 36.4 37.0 47.6 51.9 52.6
$t_{16}(\tau(\check{\beta}_{o}))$ 58.8 58.6 59.4 42.2 42.0 43.0 52.3 53.8 56.0
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1.5$ $\tau(\check{\beta}_{o})$ 20.2 20.1 20.8 63.3 65.2 65.0 85.9 87.6 87.7
$\tau(\check{\beta}_{e})$ 20.7 20.4 20.6 63.3 64.0 64.3 86.5 87.8 87.8
$t_{8}(\tau(\check{\beta}_{o}))$ 52.8 52.0 52.0 49.0 52.4 53.0 68.4 74.0 75.4
$t_{12}(\tau(\check{\beta}_{o}))$ 71.3 70.8 70.7 58.8 59.8 59.6 73.6 77.4 79.2
$t_{16}(\tau(\check{\beta}_{o}))$ 83.7 83.2 83.4 69.8 68.5 68.5 80.0 80.8 82.7
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2$ $\tau(\check{\beta}_{o})$ 27.7 27.3 28.3 80.3 82.1 82.9 95.4 96.6 97.2
$\tau(\check{\beta}_{e})$ 28.0 27.3 28.4 80.1 81.4 82.6 95.9 97.0 97.1
$t_{8}(\tau(\check{\beta}_{o}))$ 68.1 67.0 66.7 65.5 68.5 69.1 83.7 88.5 89.4
$t_{12}(\tau(\check{\beta}_{o}))$ 86.3 85.0 85.1 76.8 77.4 77.7 88.8 91.1 92.6
$t_{16}(\tau(\check{\beta}_{o}))$ 94.3 93.9 94.1 86.2 85.3 85.6 93.2 94.0 94.8
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2.5$ $\tau(\check{\beta}_{o})$ 34.3 33.7 34.7 89.5 91.7 92.1 98.6 99.2 99.3
$\tau(\check{\beta}_{e})$ 34.8 33.8 35.4 89.1 91.2 92.0 98.7 99.1 99.3
$t_{8}(\tau(\check{\beta}_{o}))$ 78.5 77.8 77.4 76.7 79.5 80.5 90.9 94.7 95.7
$t_{12}(\tau(\check{\beta}_{o}))$ 93.5 92.9 92.9 87.7 87.8 88.7 95.4 96.6 97.5
$t_{16}(\tau(\check{\beta}_{o}))$ 98.2 98.1 98.0 94.8 93.9 94.2 97.7 98.1 98.6
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
Table 7: Size and power for the discrete time models, constant volatility, MA(4) innovation
$\bar{\kappa}=0$ $\bar{\kappa}=50$ $\bar{\kappa}=100$
T 20 50 100 20 50 100 20 50 100
$\beta=0$ $\tau(\check{\beta}_{o})$ 5.0 4.9 4.8 5.2 5.3 4.8 4.5 4.7 5.2
$\tau(\check{\beta}_{e})$ 4.8 4.7 5.2 5.0 4.8 4.8 5.1 5.0 4.9
$t_{8}(\tau(\check{\beta}_{o}))$ 4.7 4.5 4.8 3.4 4.1 4.5 3.7 5.0 4.6
$t_{12}(\tau(\check{\beta}_{o}))$ 5.1 4.7 4.5 3.1 3.8 4.4 2.8 4.0 4.2
$t_{16}(\tau(\check{\beta}_{o}))$ 4.9 4.9 5.1 2.8 4.0 4.2 2.5 3.5 3.9
IVX 14.2 13.5 12.3 10.2 10.5 10.3 9.5 10.3 10.5
$\beta=0.5$ $\tau(\check{\beta}_{o})$ 7.1 6.7 7.0 12.3 12.5 12.2 16.5 16.6 17.2
$\tau(\check{\beta}_{e})$ 7.0 7.0 7.2 12.4 11.5 12.3 16.8 17.2 17.5
$t_{8}(\tau(\check{\beta}_{o}))$ 11.5 11.2 12.0 9.1 10.6 11.4 13.0 14.1 15.0
$t_{12}(\tau(\check{\beta}_{o}))$ 14.4 14.4 14.1 9.6 10.9 11.7 11.6 14.1 14.5
$t_{16}(\tau(\check{\beta}_{o}))$ 16.7 16.5 17.6 10.1 11.7 13.0 11.3 13.6 14.6
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1$ $\tau(\check{\beta}_{o})$ 10.1 9.8 10.3 25.4 25.8 24.6 39.4 39.4 40.5
$\tau(\check{\beta}_{e})$ 10.3 10.0 10.4 25.4 24.2 25.4 40.3 40.3 40.8
$t_{8}(\tau(\check{\beta}_{o}))$ 22.5 22.6 23.0 18.7 21.1 22.0 28.7 31.9 33.6
$t_{12}(\tau(\check{\beta}_{o}))$ 31.2 31.4 30.4 21.5 23.2 24.1 29.0 33.4 33.6
$t_{16}(\tau(\check{\beta}_{o}))$ 38.7 38.8 39.2 25.1 26.4 28.0 31.0 33.4 34.9
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1.5$ $\tau(\check{\beta}_{o})$ 14.2 14.2 14.7 42.4 44.1 43.1 64.4 65.1 66.4
$\tau(\check{\beta}_{e})$ 14.4 14.4 14.4 42.8 42.5 42.8 65.9 66.9 66.5
$t_{8}(\tau(\check{\beta}_{o}))$ 36.4 36.0 36.0 32.2 34.7 35.2 48.3 52.5 54.5
$t_{12}(\tau(\check{\beta}_{o}))$ 50.3 50.5 50.3 37.8 39.1 40.0 51.2 55.2 56.5
$t_{16}(\tau(\check{\beta}_{o}))$ 62.5 62.3 63.3 46.0 45.4 46.1 56.9 57.9 60.0
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2$ $\tau(\check{\beta}_{o})$ 19.0 18.9 19.6 59.7 61.4 61.6 82.9 84.3 84.6
$\tau(\check{\beta}_{e})$ 19.4 19.1 19.1 59.5 60.2 60.6 83.4 84.7 84.8
$t_{8}(\tau(\check{\beta}_{o}))$ 49.9 49.0 49.3 45.7 48.9 49.7 65.1 70.4 72.2
$t_{12}(\tau(\check{\beta}_{o}))$ 67.9 67.5 67.1 55.1 56.2 55.9 69.9 73.8 75.9
$t_{16}(\tau(\check{\beta}_{o}))$ 80.4 80.3 80.5 65.8 64.5 64.3 76.5 77.2 79.4
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2.5$ $\tau(\check{\beta}_{o})$ 24.1 24.1 24.9 73.1 75.7 75.7 92.5 93.8 94.4
$\tau(\check{\beta}_{e})$ 24.5 24.0 24.6 73.7 74.3 75.5 92.8 94.1 94.1
$t_{8}(\tau(\check{\beta}_{o}))$ 61.6 60.4 60.7 58.5 61.6 62.0 77.8 83.2 84.4
$t_{12}(\tau(\check{\beta}_{o}))$ 80.4 79.2 79.4 69.3 70.5 70.1 83.1 86.1 87.8
$t_{16}(\tau(\check{\beta}_{o}))$ 90.6 90.1 90.3 79.9 78.8 79.0 88.7 89.3 90.8
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
Table 8: Size and power for the discrete time models, structural break in volatility, MA(2) innovation
$\bar{\kappa}=0$ $\bar{\kappa}=50$ $\bar{\kappa}=100$
T 20 50 100 20 50 100 20 50 100
$\beta=0$ $\tau(\check{\beta}_{o})$ 4.5 4.9 5.1 5.3 5.2 4.8 5.1 4.7 4.7
$\tau(\check{\beta}_{e})$ 4.8 4.8 4.9 5.0 4.5 4.8 5.0 4.9 4.9
$t_{8}(\tau(\check{\beta}_{o}))$ 3.7 3.7 3.7 2.4 3.4 3.5 2.8 3.5 3.4
$t_{12}(\tau(\check{\beta}_{o}))$ 4.5 4.3 3.8 3.0 3.6 3.8 2.7 3.3 3.6
$t_{16}(\tau(\check{\beta}_{o}))$ 4.3 4.8 4.6 2.7 3.4 4.1 2.5 3.1 3.7
IVX 30.7 31.2 30.1 32.8 34.2 34.3 33.9 34.5 36.0
$\beta=0.5$ $\tau(\check{\beta}_{o})$ 7.7 8.1 8.7 13.7 13.7 13.2 18.5 18.7 19.3
$\tau(\check{\beta}_{e})$ 8.0 8.2 8.4 13.6 13.0 12.8 19.3 19.7 19.7
$t_{8}(\tau(\check{\beta}_{o}))$ 11.5 11.6 11.7 9.2 10.5 10.8 12.9 15.1 15.6
$t_{12}(\tau(\check{\beta}_{o}))$ 16.0 16.0 16.0 10.8 12.1 12.4 13.3 16.1 16.6
$t_{16}(\tau(\check{\beta}_{o}))$ 19.6 20.2 20.5 12.2 13.8 14.6 14.6 16.1 17.6
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1$ $\tau(\check{\beta}_{o})$ 13.5 13.5 14.3 29.2 29.3 28.5 44.3 45.0 46.0
$\tau(\check{\beta}_{e})$ 14.0 13.8 13.9 28.2 28.1 27.3 46.1 46.6 46.1
$t_{8}(\tau(\check{\beta}_{o}))$ 24.8 24.7 24.8 21.1 23.1 23.4 32.0 35.3 36.3
$t_{12}(\tau(\check{\beta}_{o}))$ 36.3 36.6 36.4 27.3 28.2 28.0 35.0 39.3 39.6
$t_{16}(\tau(\check{\beta}_{o}))$ 44.8 45.7 46.1 31.5 33.6 34.2 39.5 41.0 43.0
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1.5$ $\tau(\check{\beta}_{o})$ 20.4 20.4 21.5 46.8 48.1 47.6 68.9 70.8 72.2
$\tau(\check{\beta}_{e})$ 21.3 20.8 21.0 46.9 46.3 47.5 71.2 71.5 72.5
$t_{8}(\tau(\check{\beta}_{o}))$ 39.4 38.6 39.3 35.6 37.6 38.0 51.1 56.1 57.6
$t_{12}(\tau(\check{\beta}_{o}))$ 56.3 57.2 56.4 45.7 46.4 45.9 57.6 62.3 63.3
$t_{16}(\tau(\check{\beta}_{o}))$ 67.5 69.2 69.3 53.9 55.3 55.1 63.8 65.9 67.7
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2$ $\tau(\check{\beta}_{o})$ 27.6 27.7 28.9 62.7 64.9 64.7 84.4 87.3 88.0
$\tau(\check{\beta}_{e})$ 28.3 28.0 28.0 63.2 63.2 64.5 85.7 87.3 87.9
$t_{8}(\tau(\check{\beta}_{o}))$ 51.7 51.4 51.6 48.5 50.0 51.8 65.0 71.3 72.7
$t_{12}(\tau(\check{\beta}_{o}))$ 71.7 72.5 71.4 62.0 62.2 62.1 74.0 77.3 79.1
$t_{16}(\tau(\check{\beta}_{o}))$ 82.3 83.1 83.1 71.6 72.3 72.0 81.1 82.0 83.9
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2.5$ $\tau(\check{\beta}_{o})$ 34.4 34.3 35.5 74.0 76.9 77.2 92.1 94.6 95.4
$\tau(\check{\beta}_{e})$ 35.1 34.6 34.3 74.7 75.9 77.3 93.0 94.6 95.3
q=4 23.8 23.3 23.4 34.1 37.0 38.4 46.8 50.2 51.4
$t_{8}(\tau(\check{\beta}_{o}))$ 61.7 61.1 61.2 58.6 60.4 62.5 74.1 80.5 82.4
$t_{12}(\tau(\check{\beta}_{o}))$ 81.7 82.3 81.3 73.6 73.7 74.4 84.0 86.5 88.2
$t_{16}(\tau(\check{\beta}_{o}))$ 90.6 91.1 91.4 83.2 83.1 83.0 90.1 91.1 91.9
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
Table 9: Size and power for the discrete time models, structural break in volatility, MA(4) innovation
$\bar{\kappa}=0$ $\bar{\kappa}=50$ $\bar{\kappa}=100$
T 20 50 100 20 50 100 20 50 100
$\beta=0$ $\tau(\check{\beta}_{o})$ 4.6 4.9 5.0 4.5 5.0 5.0 4.3 5.0 5.0
$\tau(\check{\beta}_{e})$ 4.8 4.7 4.8 4.6 4.6 4.8 4.9 5.1 5.4
q=4 3.3 3.1 3.0 2.2 2.5 3.0 2.4 2.7 3.0
$t_{8}(\tau(\check{\beta}_{o}))$ 3.4 3.7 3.5 2.2 2.7 3.0 2.0 2.7 3.2
$t_{12}(\tau(\check{\beta}_{o}))$ 4.2 4.1 3.7 2.2 2.8 3.2 1.8 2.6 3.0
$t_{16}(\tau(\check{\beta}_{o}))$ 4.3 4.4 4.4 2.5 2.7 3.3 1.6 2.3 3.0
IVX 32.7 32.5 31.2 32.1 33.5 34.3 32.4 34.1 35.5
$\beta=0.5$ $\tau(\check{\beta}_{o})$ 9.8 10.1 11.0 7.9 8.9 9.0 9.0 11.4 12.0
$\tau(\check{\beta}_{e})$ 10.2 10.2 10.4 8.0 8.3 8.8 10.4 11.8 12.4
$t_{8}(\tau(\check{\beta}_{o}))$ 15.1 15.3 15.5 5.5 6.7 7.5 5.8 8.1 9.0
$t_{12}(\tau(\check{\beta}_{o}))$ 21.6 22.5 22.1 7.1 9.0 9.3 6.1 8.9 9.9
$t_{16}(\tau(\check{\beta}_{o}))$ 26.3 28.5 28.5 9.0 10.3 12.1 6.6 9.6 10.7
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1$ $\tau(\check{\beta}_{o})$ 18.9 19.2 20.8 13.9 16.0 15.7 18.1 23.3 25.3
$\tau(\check{\beta}_{e})$ 19.3 20.0 20.0 14.0 14.7 15.8 20.0 24.4 26.2
$t_{8}(\tau(\check{\beta}_{o}))$ 34.2 34.0 34.5 11.9 14.1 15.1 13.3 18.6 19.8
$t_{12}(\tau(\check{\beta}_{o}))$ 48.3 50.4 50.2 18.2 20.2 21.7 16.1 21.9 23.3
$t_{16}(\tau(\check{\beta}_{o}))$ 59.6 61.5 62.9 22.8 26.5 28.3 18.2 24.6 26.9
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1.5$ $\tau(\check{\beta}_{o})$ 29.3 29.7 30.9 22.0 25.4 25.5 31.0 39.5 43.1
$\tau(\check{\beta}_{e})$ 29.4 29.7 30.0 21.9 24.4 25.8 33.9 40.5 43.3
$t_{8}(\tau(\check{\beta}_{o}))$ 50.5 50.7 50.9 21.5 24.0 25.3 24.4 31.9 33.6
$t_{12}(\tau(\check{\beta}_{o}))$ 69.6 71.3 71.2 33.1 35.7 37.3 30.8 38.3 40.3
$t_{16}(\tau(\check{\beta}_{o}))$ 81.2 83.1 83.2 41.5 46.4 47.7 36.5 44.0 47.4
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2$ $\tau(\check{\beta}_{o})$ 38.4 38.1 39.8 31.5 36.6 36.8 45.5 56.6 60.8
$\tau(\check{\beta}_{e})$ 38.5 38.7 38.8 31.8 35.8 37.5 49.0 58.0 60.4
$t_{8}(\tau(\check{\beta}_{o}))$ 63.2 63.5 63.0 31.9 34.5 36.6 36.1 45.2 47.7
$t_{12}(\tau(\check{\beta}_{o}))$ 82.2 83.8 83.8 49.2 51.5 52.3 46.4 54.8 56.8
$t_{16}(\tau(\check{\beta}_{o}))$ 91.6 92.8 92.7 60.3 64.8 65.0 54.6 62.8 65.2
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2.5$ $\tau(\check{\beta}_{o})$ 45.9 45.2 46.9 41.8 47.2 48.6 59.1 70.3 74.9
$\tau(\check{\beta}_{e})$ 45.5 45.7 46.3 41.9 47.2 48.4 62.4 71.8 74.2
$t_{8}(\tau(\check{\beta}_{o}))$ 71.5 71.7 70.9 41.7 45.3 46.4 46.5 56.7 59.3
$t_{12}(\tau(\check{\beta}_{o}))$ 89.5 90.3 90.3 62.4 64.9 64.7 60.0 68.1 69.8
$t_{16}(\tau(\check{\beta}_{o}))$ 96.2 96.8 96.6 74.1 77.3 78.2 70.3 76.7 78.3
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
Table 10: Size and power for the discrete time models, regime switching in volatility, MA(2) innovation
$\bar{\kappa}=0$ $\bar{\kappa}=50$ $\bar{\kappa}=100$
T 20 50 100 20 50 100 20 50 100
$\beta=0$ $\tau(\check{\beta}_{o})$ 3.2 4.0 4.3 4.7 5.0 5.0 5.2 4.9 4.8
$\tau(\check{\beta}_{e})$ 3.8 3.9 4.2 5.0 4.9 4.9 4.9 5.1 4.8
$t_{8}(\tau(\check{\beta}_{o}))$ 2.6 2.9 3.5 2.7 2.5 3.0 3.4 3.4 4.0
$t_{12}(\tau(\check{\beta}_{o}))$ 2.4 2.3 3.2 1.8 2.0 2.5 2.3 2.6 3.2
$t_{16}(\tau(\check{\beta}_{o}))$ 2.3 2.1 2.8 1.7 1.8 2.3 2.1 2.0 2.4
IVX 13.5 13.2 12.4 9.7 11.1 11.0 10.3 11.2 12.4
$\beta=0.5$ $\tau(\check{\beta}_{o})$ 5.8 6.8 7.4 16.1 17.1 16.6 23.4 24.1 24.4
$\tau(\check{\beta}_{e})$ 6.4 6.7 6.9 16.8 16.8 16.6 23.9 24.9 23.6
$t_{8}(\tau(\check{\beta}_{o}))$ 10.2 10.8 12.1 11.0 11.5 12.3 16.2 17.4 19.0
$t_{12}(\tau(\check{\beta}_{o}))$ 13.3 13.8 15.0 11.5 10.7 11.8 17.0 17.1 18.3
$t_{16}(\tau(\check{\beta}_{o}))$ 16.1 16.0 18.8 12.9 12.0 12.9 17.7 15.4 18.0
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1$ $\tau(\check{\beta}_{o})$ 10.4 11.3 12.0 39.6 40.9 39.3 56.7 60.4 60.6
$\tau(\check{\beta}_{e})$ 11.0 11.7 11.8 39.7 39.6 39.8 57.3 60.9 59.6
$t_{8}(\tau(\check{\beta}_{o}))$ 26.2 27.6 28.7 28.6 30.4 30.1 41.8 45.1 47.2
$t_{12}(\tau(\check{\beta}_{o}))$ 37.5 39.1 40.4 34.4 33.2 33.4 46.9 47.9 48.9
$t_{16}(\tau(\check{\beta}_{o}))$ 47.6 49.3 50.2 41.3 38.2 38.3 53.6 49.7 50.2
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1.5$ $\tau(\check{\beta}_{o})$ 17.3 18.2 19.0 64.3 67.1 63.9 81.5 86.0 87.0
$\tau(\check{\beta}_{e})$ 17.9 18.6 18.7 63.8 65.3 64.4 82.2 86.8 86.2
$t_{8}(\tau(\check{\beta}_{o}))$ 45.8 46.7 47.9 49.7 51.9 50.3 65.8 70.3 72.3
$t_{12}(\tau(\check{\beta}_{o}))$ 63.9 65.3 66.4 60.3 58.8 58.2 74.7 75.4 76.2
$t_{16}(\tau(\check{\beta}_{o}))$ 76.6 77.6 78.2 70.9 67.5 65.9 82.3 80.3 78.4
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2$ $\tau(\check{\beta}_{o})$ 25.0 25.4 26.5 80.2 83.1 81.7 92.0 95.5 96.4
$\tau(\check{\beta}_{e})$ 25.6 26.2 25.9 79.4 82.0 82.1 92.5 95.5 96.1
$t_{8}(\tau(\check{\beta}_{o}))$ 62.3 62.9 63.6 66.5 68.0 67.3 80.4 84.3 86.0
$t_{12}(\tau(\check{\beta}_{o}))$ 81.2 82.7 83.6 79.0 78.2 76.9 89.0 90.0 90.2
$t_{16}(\tau(\check{\beta}_{o}))$ 91.1 91.9 92.4 88.4 85.9 84.8 94.6 93.5 93.3
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2.5$ $\tau(\check{\beta}_{o})$ 32.6 32.5 33.9 88.9 91.4 90.7 96.5 98.6 99.1
$\tau(\check{\beta}_{e})$ 33.4 33.6 32.7 88.2 90.9 90.6 96.8 98.6 98.9
$t_{8}(\tau(\check{\beta}_{o}))$ 73.9 75.2 75.6 78.3 79.5 78.8 88.7 91.8 92.9
$t_{12}(\tau(\check{\beta}_{o}))$ 91.0 91.9 92.2 89.7 89.0 88.2 95.7 95.9 96.1
$t_{16}(\tau(\check{\beta}_{o}))$ 96.9 97.1 97.4 95.7 94.6 94.2 98.3 98.0 97.9
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
Table 11: Size and power for the discrete time models, regime switching in volatility, MA(4) innovation
$\bar{\kappa}=0$ $\bar{\kappa}=50$ $\bar{\kappa}=100$
T 20 50 100 20 50 100 20 50 100
$\beta=0$ $\tau(\check{\beta}_{o})$ 3.8 3.9 4.3 4.3 5.1 5.3 4.6 4.8 4.9
$\tau(\check{\beta}_{e})$ 4.0 4.2 4.3 4.8 4.9 4.7 4.9 4.8 4.5
$t_{8}(\tau(\check{\beta}_{o}))$ 3.5 3.5 3.5 2.3 2.3 2.5 2.3 2.4 3.2
$t_{12}(\tau(\check{\beta}_{o}))$ 3.1 3.4 3.2 1.8 2.0 2.6 2.0 1.9 2.6
$t_{16}(\tau(\check{\beta}_{o}))$ 3.5 3.2 3.4 1.6 1.8 2.1 1.6 1.8 2.0
IVX 15.6 13.6 13.4 9.8 10.5 11.1 9.6 11.2 12.4
$\beta=0.5$ $\tau(\check{\beta}_{o})$ 7.8 7.9 8.8 10.2 11.6 11.5 12.4 14.5 15.8
$\tau(\check{\beta}_{e})$ 8.5 8.7 8.7 10.5 11.4 11.9 12.7 14.4 14.9
$t_{8}(\tau(\check{\beta}_{o}))$ 16.0 15.5 17.4 8.2 7.9 8.0 8.8 9.8 11.2
$t_{12}(\tau(\check{\beta}_{o}))$ 20.7 21.9 23.2 9.3 8.9 9.2 9.7 10.4 11.3
$t_{16}(\tau(\check{\beta}_{o}))$ 26.5 26.4 28.5 10.8 9.8 11.1 10.5 10.1 10.7
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1$ $\tau(\check{\beta}_{o})$ 15.3 15.4 17.0 22.7 24.9 24.1 27.2 34.7 36.4
$\tau(\check{\beta}_{e})$ 16.2 16.8 16.9 22.9 24.8 24.0 28.5 34.4 35.7
$t_{8}(\tau(\check{\beta}_{o}))$ 37.7 38.7 40.8 21.0 20.2 19.5 22.9 25.2 28.1
$t_{12}(\tau(\check{\beta}_{o}))$ 53.2 53.8 56.8 26.3 25.9 24.4 28.0 29.0 29.3
$t_{16}(\tau(\check{\beta}_{o}))$ 65.2 66.9 69.0 34.5 31.2 32.7 33.5 31.7 32.3
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1.5$ $\tau(\check{\beta}_{o})$ 24.1 24.2 26.8 39.3 42.3 41.9 47.5 58.9 61.5
$\tau(\check{\beta}_{e})$ 25.4 25.9 26.3 38.3 41.5 41.5 47.6 59.1 60.6
$t_{8}(\tau(\check{\beta}_{o}))$ 59.4 59.6 61.8 37.5 36.3 34.6 42.3 44.5 47.6
$t_{12}(\tau(\check{\beta}_{o}))$ 78.1 78.7 81.1 49.4 48.4 45.9 52.1 53.0 53.6
$t_{16}(\tau(\check{\beta}_{o}))$ 88.3 89.8 90.7 62.2 58.6 58.4 62.0 59.9 59.4
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2$ $\tau(\check{\beta}_{o})$ 32.9 33.4 35.4 55.1 59.0 59.1 64.5 77.4 80.3
$\tau(\check{\beta}_{e})$ 34.3 35.1 35.0 53.1 58.4 58.5 64.7 77.3 79.3
$t_{8}(\tau(\check{\beta}_{o}))$ 74.1 74.3 76.0 53.0 52.1 50.5 59.3 62.3 65.2
$t_{12}(\tau(\check{\beta}_{o}))$ 90.3 91.1 92.0 69.3 68.4 65.7 72.1 72.5 73.0
$t_{16}(\tau(\check{\beta}_{o}))$ 96.6 97.2 97.6 82.1 80.2 78.8 82.8 80.5 79.4
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2.5$ $\tau(\check{\beta}_{o})$ 40.9 41.0 42.4 67.3 71.8 72.4 76.2 87.7 91.3
$\tau(\check{\beta}_{e})$ 42.3 42.4 42.3 64.5 71.5 71.8 76.0 87.6 89.9
$t_{8}(\tau(\check{\beta}_{o}))$ 83.8 83.7 84.8 66.6 65.6 63.7 71.9 75.0 77.3
$t_{12}(\tau(\check{\beta}_{o}))$ 95.8 96.1 96.7 82.8 81.9 79.9 85.1 85.1 85.0
$t_{16}(\tau(\check{\beta}_{o}))$ 99.0 99.3 99.4 92.7 91.5 90.0 93.0 91.2 90.8
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
Table 12: Empirical results
Series Frequency $\beta$ Cauchy RT Cauchy VC $t_{12}(\check{\gamma})$ $t_{16}(\check{\gamma})$ $\tau(\check{\beta})$
Panel A: D/P as predictor for the period of 1927-2011
CRSP Monthly 0.005 1.06 0.85 0.42 0.44 0.43
Quarterly 0.007 0.48 1.26 1.29 1.34 1.09
Yearly 0.063 0.99 $2.15^{**}$ $2.04^{**}$ $2.35^{**}$ $1.82^{**}$
S&P500 Monthly 0.003 0.74 1.14 0.89 0.82 0.72
Quarterly 0.008 0.65 $1.78^{**}$ $1.75^{*}$ $1.73^{*}$ $1.31^{*}$
Yearly 0.042 0.82 $2.24^{**}$ $2.33^{**}$ $2.00^{**}$ $1.95^{**}$
Panel B: D/P as predictor for the period of 1927-2011 with jumps removed
CRSP Monthly 0.001 0.22 $50^{*} 1.1$5 1.06 0.90
Quarterly 0.015 0.93 $2.02^{**}$ $2.45^{**}$ $2.40^{**}$ $1.46^{*}$
Yearly 0.005 0.05 0.56 0.56 0.53 0.45
S&P500 Monthly 0.002 0.48 $1.66^{*}$ 1.35 1.14 1.02
Quarterly 0.017 1.21 $1.49^{*}$ 1.70 1.61 1.20
Yearly 0.008 0.11 $1.51^{*}$ $2.45^{**}$ $2.56^{**}$ $1.80^{**}$
Panel C: E/P as predictor for the period of 1950-2011
S&P500 Monthly 0.000 -0.05 0.32 0.26 0.22 0.30
Quarterly 0.007 0.39 0.39 0.38 0.35 0.35
Yearly 0.059 0.83 0.76 1.04 0.23 0.84
Panel D: E/P as predictor for the period of 1950-2011 with jumps removed
S&P500 Monthly 0.001 0.16 0.01 0.14 0.18 0.11
Quarterly 0.000 0.02 0.71 0.78 0.61 0.48
Yearly -0.038 -0.38 0.69 -0.05 0.55 0.44
Test results for return predictability of the NYSE/AMEX value-weighted index (CRSP) and S&P 500 using the Cauchy RT, Cauchy VC, $t_{q}$ ($q=12,16$), and $\tau(\check{\beta})$ tests across different regression frequencies. Panels A–B use the dividend–price ratio (D/P), and Panels C–D use the earnings–price ratio (E/P) as predictors. Significance at the 5% and 1% levels is denoted by ‘‘’’ and ‘‘∗∗’’, respectively.