### v3

# Rustam Ibragimova,b    Jihyun Kimd    Anton Skrobotove a Imperial College Business School b New Economic School

Abstract
This paper develops robust inference methods for predictive regressions that address key challenges posed by endogenously persistent or heavy-tailed regressors, as well as persistent volatility in errors. Building on the Cauchy estimation framework, we propose two novel tests: one based on $t$-statistic group inference and the other employing a hybrid approach that combines Cauchy and OLS estimation. These methods effectively mitigate size distortions that commonly arise in standard inference procedures under endogeneity, near nonstationarity, heavy tails, and persistent volatility. The proposed tests are simple to implement and applicable to both continuous- and discrete-time models. Extensive simulation experiments demonstrate favorable finite-sample performance across a range of realistic settings. An empirical application examines the predictability of excess stock returns using the dividend–price and earnings–price ratios as predictors. The results suggest that the dividend–price ratio possesses predictive power, whereas the earnings–price ratio does not significantly forecast returns.
Keywords: predictive regressions, robust inference, near nonstationarity, heterogeneity, heavy tails, persistent volatility, endogeneity.
JEL Codes: C12, C22
1 Introduction
Predictive regressions play a central role in empirical finance, providing a framework for assessing whether financial or macroeconomic variables can forecast future returns. Prominent applications include the forecasting of equity and aggregate returns (see, among others, Campbell and Yogo, 2006; Goyal and Welch, 2006; Campbell and Thompson, 2008; Hirshleifer et al., 2009; Kelly and Jiang, 2014; Rapach et al., 2016; Møller and Rangvid, 2015; Goyal et al., 2024, and references therein) and tests of market efficiency (e.g., Fama, 1970, 1984, 1991; the review in Martin and Nagel, 2022, and references therein). A large body of work has examined the econometric properties of predictive regressions (see Phillips, 2015 for a review), highlighting several statistical challenges that complicate inference on return predictability.
A key difficulty arises from the persistence and endogeneity of commonly used predictors. Variables such as the dividend–price and earnings–price ratios typically exhibit near–unit–root behavior, and their innovations are correlated with future returns over long horizons. This combination induces substantial biases in conventional hypothesis tests (see, e.g., Stambaugh, 1999; Kim and Park, 2017). In addition, stock return volatility is stochastic and highly persistent (Jacquier et al., 2004; Hansen and Lunde, 2014), and Cavaliere (2004) shows that such volatility can lead to severe size distortions in tests that rely on stationarity. Predictive regression data also often exhibit heavy tails, jumps, structural breaks, and regime switching, further undermining standard inference methods.
A large literature has proposed procedures to address persistent endogeneity in predictive regressions. Notably, Campbell and Yogo (2006), Chen and Deo (2009), Phillips and Magdalinos (2009), and Kostakis et al. (2015), among others, develop inference methods that perform well in the presence of persistent and endogenous regressors. However, these approaches typically rely on assumptions such as homoskedasticity or weak dependence in volatility and may not remain valid under empirically relevant features such as persistent or nonstationary volatility, structural breaks, or regime switching. For example, Ibragimov et al. (2024) show that standard tests can suffer from severe size distortions in the presence of persistent volatility.
To address these limitations, Choi et al. (2016) propose an inference method (the Cauchy RT) based on the Cauchy estimator and a time-change transformation in a continuous-time framework. [^2] See also Bu et al. (2023) for an alternative method robust to endogenously persistent or heavy-tailed regressors and persistent volatility in continuous time. Ibragimov et al. (2024) introduce another approach (the Cauchy VC), also based on the Cauchy estimator but with a nonparametric volatility correction, which applies to both continuous- and discrete-time models.
This paper proposes two practical tests that serve as robust alternatives to these methods. The proposed procedures are robust to heterogeneous and persistent volatility, as well as to endogenous, persistent, and heavy-tailed regressors. Both methods employ Cauchy estimation, as in Choi et al. (2016) and Ibragimov et al. (2024), to address endogeneity, persistence, and heavy tails. The two approaches differ in their treatment of heterogeneous volatility: the first extends the $t$-statistic–based group inference of Ibragimov and Müller (2010) to asymptotically normal Cauchy estimators, while the second is a hybrid procedure that combines Cauchy and OLS estimation by using the Cauchy estimator for the coefficient and OLS residuals for variance estimation.
The proposed methods are simple to implement and avoid the technical complexities associated with existing approaches, such as the time-change transformation in Choi et al. (2016) or the nonparametric volatility correction in Ibragimov et al. (2024). Although they rely on an asymptotic exogeneity condition on volatility, simulation results show that they perform well in finite samples, even under mild violations of this condition. Moreover, the proposed methods apply to both continuous- and discrete-time models. Overall, our procedures complement existing Cauchy-based methods and provide a practical alternative for robust inference under a broad range of empirically relevant environments.
In addition to the continuous-time framework, we evaluate the performance of our methods in discrete-time predictive regressions and compare them with the IVX approach of Kostakis et al. (2015). The IVX method is specifically designed to address persistence and endogeneity and performs well under homoskedastic or weakly dependent volatility. Our results highlight that the two approaches are complementary. While IVX delivers strong performance under standard conditions, it may suffer from size distortions in the presence of nonstationary or persistent volatility. By contrast, our methods are designed to remain robust to such features, including heavy tails and time-varying volatility. This comparison underscores the advantages of each approach across empirically relevant settings.
The remainder of the paper is organized as follows. Section 2 introduces the predictive regression model and the Cauchy estimator. Section 3 develops the proposed inference procedures and establishes their theoretical properties. Section 4 extends the analysis to models with multiple predictors and intercepts. Sections 5 and 6 present simulation results and an empirical application. Section 7 concludes. All proofs are provided in the Appendix, and additional theoretical analysis, as well as supplementary simulation and empirical results, are reported in the accompanying Online Appendix.
2 Predictive Regressions and the Cauchy Estimator
2.1 Model and Issues
Throughout the paper, we consider $(\mathcal{F}_{t})$-adapted processes defined on a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t}),P)$ equipped with an increasing filtration $(\mathcal{F}_{t})$ of sub-$\sigma$-fields of $\mathcal{F}$. Our objective is to test the (un)predictability of the process $(y_{t})$ (e.g., the time series of excess stock returns) based on a covariate process $(x_{t})$ (e.g., the time series of price–dividend ratios). As usual, we consider the linear predictive regression model
$y_{t}$ $=\alpha+\beta x_{t-1}+u_{t},\quad t=1,\dots,T.$ (1)
Following the standard specification for a volatility model, we assume that
$u_{t}=v_{t}\varepsilon_{t},$
where $(v_{t})$ is a volatility process and $(\varepsilon_{t})$ is a martingale difference sequence (MDS) with respect to $(\mathcal{F}_{t})$. We impose the following regularity conditions on $(v_{t},\varepsilon_{t})$.
Assumption 2.1.
(a) $E(\varepsilon_{t}^{2}|\mathcal{F}_{t-1})=1$; (b) $(v_{t})$ is $(\mathcal{F}_{t-1})$-adapted and nonnegative; and (c) $T^{-1}\sum_{t=1}^{T}E\!\left[\varepsilon_{t}^{2}1\{|\varepsilon_{t}|\geq\delta\sqrt{T}\}\middle|\mathcal{F}_{t-1}\right]\!\to_{p}0$ for any $\delta>0$.
Conditions (a) and (b) are standard and ensure that the conditional variance of $u_{t}$ is identified: $E(u_{t}^{2}|\mathcal{F}_{t-1})=v_{t}^{2}$. Condition (c) is a conditional Lindeberg condition, which holds, for example, if $\sup_{1\leq t\leq T}E(|\varepsilon_{t}|^{2+\delta}|\mathcal{F}_{t-1})$ is bounded for some $\delta>0$ with probability one. See Ibragimov et al. (2024) and references therein for further discussion and examples of processes satisfying Assumption 2.1.
The hypothesis of unpredictability of $(y_{t})$ corresponds to $H_{0}:\beta=0$ in regression (1). It is well-known that standard OLS $t$-statistic inference is not robust to many empirically relevant features of financial data. For instance, the OLS estimator of $\beta$ is not asymptotically Gaussian under $H_{0}$ if $(x_{t})$ is endogenous and (nearly) nonstationary (see Elliott and Stock, 1994; Phillips, 1987; Giraitis and Phillips, 2006; Phillips and Magdalinos, 2007a; Kostakis et al., 2015), or even if $(x_{t})$ is stationary but has infinite variance (e.g., Granger and Orr, 1972; Embrechts et al., 1997; Ibragimov et al., 2015). This non-Gaussianity persists even when the errors are homoskedastic with $v_{t}^{2}=\sigma^{2}$ for all tt. [^3] As usual, the endogeneity of $(x_{t-1})$ refers to the existence of nonzero long-run covariance between the innovations of $(y_{t})$ and $(x_{t-1})$. Furthermore, stock return data exhibit time-varying and stochastically persistent volatility, which causes the distribution of the OLS $t$-statistic to deviate from standard normality, leading to size distortions in conventional tests (see Choi et al., 2016; Ibragimov et al., 2024).
2.2 The Cauchy Estimator
Both inference methods proposed in this paper build upon the following Cauchy estimator of $\beta$ (assuming no intercept, i.e., $\alpha=0$):
$\check{\beta}=\left(\sum_{t=1}^{T}|x_{t-1}|\right)^{-1}\sum_{t=1}^{T}\text{sign}(x_{t-1})\,y_{t},$
where $\text{sign}(\cdot)$ denotes the sign function, $\text{sign}(x)=1$ for $x\geq 0$ and $\text{sign}(x)=-1$ for $x<0$. The estimator $\check{\beta}$ can be interpreted as an instrumental variable (IV) estimator using $\text{sign}(x_{t-1})$ as an instrument for $x_{t-1}$ (see, e.g., So and Shin, 1999; Breitung and Demetrescu, 2015; Kim and Meddahi, 2020; Shephard, 2020).
Under Assumption 2.1, $\text{sign}(x_{t-1})\varepsilon_{t}$ (denoted by $\xi_{t}$) is a unit-variance MDS with respect to $(\mathcal{F}_{t})$. Define the continuous-time partial sum process $(W^{T}(r),0\leq r\leq 1)$ by
$W^{T}(r)=T^{-1/2}\sum_{t=1}^{[Tr]}\xi_{t},$
which takes values in $\mathbf{D}_{\mathbb{R}}[0,1]$, the space of càdlàg functions on $[0,1]$ with values in $\mathbb{R}^{d}$. By the functional central limit theorem for martingales (Theorem 18.2 of Billingsley, 1986), we have $W^{T}\Rightarrow W$ in $\mathbf{D}_{\mathbb{R}}[0,1]$, where $W$ is a standard Brownian motion.
For the volatility process $(v_{t})$, define $\sigma^{T}(r)=v_{[Tr]}$ on $\mathbf{D}_{\mathbb{R}^{+}}[0,1]$. Then the Cauchy estimator can be expressed in terms of $\sigma^{T}$ and $W^{T}$ as
$\left(\sum_{t=1}^{T}|x_{t-1}|/\sqrt{T}\right)\!\bigl(\check{\beta}-\beta\bigr)$ $=T^{-1/2}\sum_{t=1}^{T}\text{sign}(x_{t-1})v_{t}\varepsilon_{t}=\int_{0}^{1}\sigma^{T}(r)\,dW^{T}(r).$
Following Ibragimov et al. (2024), we assume that the volatility process $\sigma^{T}$ is persistent in the sense that it admits a limiting process $\sigma$ defined on $[0,1]$ such that $(W^{T},\sigma^{T})\Rightarrow(W,\sigma)$ jointly.
Assumption 2.2.
There exists a nonnegative process $\sigma$ on $\mathbf{D}_{\mathbb{R}^{+}}[0,1]$ such that
$0<\int_{0}^{1}\sigma^{2}(r)\,dr<\infty,\quad\text{and}\quad(W^{T},\sigma^{T})\Rightarrow(W,\sigma)\quad\text{in}\quad\mathbf{D}_{\mathbb{R}\times\mathbb{R}^{+}}[0,1],$
where $W$ is a standard Brownian motion adapted to the same filtration as $\sigma$.
Assumption 2.2 encompasses a wide class of models, including those with nonstationary volatility, regime switching, or structural breaks. [^4] Assumptions 2.1 and 2.2 exclude some globally homoskedastic processes, such as stationary GARCH models. However, the hybrid testing procedure proposed later remains valid under $T^{-1}\sum_{t=1}^{T}v_{t}^{2}\to_{p}\omega^{2}>0$, which includes conditionally heteroskedastic but globally homoskedastic processes, such as stationary GARCH models (see also Section 4 of Ibragimov et al., 2024). It also covers cases with deterministic volatility $v_{t}=\sigma(t/T)$, as in Cavaliere and Taylor (2007, 2008), Xu and Phillips (2008), and Harvey et al. (2019), among others. [^5] Assumption 2.2 is a simplified version of the condition $v_{[Tr]}/a_{T}\Rightarrow\sigma_{r}$ in Assumption 2 of Cavaliere and Taylor (2009). We focus on stochastically bounded volatilities with $a_{T}=1$, excluding explosive volatility settings ($a_{T}\to\infty$) for simplicity. It further includes nonstationary volatility processes such as those in Hansen (1995) and Chung and Park (2007), where $v_{t}^{2}$ is a smooth positive transformation of a (near) unit root process. Overall, Assumptions 2.1 and 2.2 are general enough to allow for stochastic and discontinuous volatility—features commonly observed in financial returns.
Under Assumptions 2.1 and 2.2, the properly normalized Cauchy estimator satisfies
$\left(\sum_{t=1}^{T}|x_{t-1}|/\sqrt{T}\right)\!(\check{\beta}-\beta)\Rightarrow\int_{0}^{1}\sigma(r)\,dW(r),$
by standard results on the convergence of stochastic integrals (see Hansen, 1992; Kurtz and Protter, 1991; Ibragimov et al., 2024). The limit $\int_{0}^{1}\sigma(r)\,dW(r)$ is in general a non-Gaussian martingale, becoming Gaussian only if $W$ and $\sigma$ are independent. In that case, $\int_{0}^{1}\sigma(r)\,dW(r)$ is a scale mixture of normals with variance $\int_{0}^{1}\sigma^{2}(r)\,dr$, denoted
$\int_{0}^{1}\sigma(r)\,dW(r)\;=_{d}\;\mathbb{MN}\!\left(0,\int_{0}^{1}\sigma^{2}(r)\,dr\right).$
We formalize the independence assumption as follows.
Assumption 2.3.
The processes $\sigma$ and $W$ in Assumption 2.2 are independent.
Assumption 2.3 requires the volatility process $\sigma^{T}$ to be asymptotically independent of the martingale $W^{T}$, but does not preclude finite-sample dependence. For example, consider
$\sigma^{T}\!\left(t/T\right)=T^{-\delta}f(x_{t-1},\varepsilon_{t})+\sigma_{0}^{T}\!\left(t/T\right),\quad\delta>0,$
where $f:\mathbb{R}^{2}\to\mathbb{R}^{+}$ is bounded and $\sigma_{0}^{T}$ is independent of $W^{T}$ with $(W^{T},\sigma_{0}^{T})\Rightarrow(W,\sigma)$, where $W$ and $\sigma$ are independent. For any $\delta>0$, the volatility process $\sigma^{T}$ in this example satisfies Assumption 2.3, even though $\sigma^{T}$ and $W^{T}$ may be dependent for any fixed $T>0$.
In the following sections, we develop inference methods based on the Cauchy estimator. Section 3 focuses on predictive regressions with a single predictor and no intercept, while Section 4 extends the analysis to models with multiple predictors and an intercept.
3 Robust Inference for Predictive Regressions
3.1 Robust $t$-Statistic Approach
The first approach relies on $t$-statistic-based inference using group estimates of $\beta$, as proposed by Ibragimov and Müller (2010) (see also Ibragimov and Müller, 2016; Section 3.3 of Ibragimov et al., 2015). The method is based on normalized Cauchy estimators—specifically, the numerator of the Cauchy estimator divided by $\sqrt{T}$ in the full-sample case:
$\check{\gamma}=\frac{\sum_{t=1}^{T}|x_{t-1}|}{\sqrt{T}}\check{\beta}=\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\text{sign}(x_{t-1})y_{t}.$ (2)
Following the $t$-statistic approach, we partition the sample into a fixed number $q\geq 2$ of approximately equal groups of consecutive observations. The observation $(y_{t},x_{t-1})$ at time $t$ belongs to the $j$th group $\mathcal{G}_{j}$ if
$t\in\mathcal{G}_{j}=\{s:(j-1)[T/q]<s\leq j[T/q]\},\quad j=1,\dots,q.$
We compute the normalized Cauchy statistic in (2) within each group:
$\check{\gamma}_{j}=\sqrt{\frac{q}{T}}\sum_{t=(j-1)[T/q]+1}^{j[T/q]}\text{sign}(x_{t-1})y_{t}.$ (3)
The $t$-statistic based on the $q$ group statistics $\{\check{\gamma}_{j}\}_{j=1}^{q}$ is given by
$t_{q}(\check{\gamma})=\sqrt{q}\,\frac{\bar{\gamma}}{s_{\gamma}},$ (4)
where
$\bar{\gamma}=q^{-1}\sum_{j=1}^{q}\check{\gamma}_{j},\qquad s_{\gamma}^{2}=(q-1)^{-1}\sum_{j=1}^{q}(\check{\gamma}_{j}-\bar{\gamma})^{2}.$
Under the null hypothesis $H_{0}:\beta=0$, the test rejects $H_{0}$ in favor of $H_{A}:\beta\neq 0$ if $|t_{q}(\check{\gamma})|>cv_{q}(\alpha)$, where $cv_{q}(\alpha)$ denotes the two-sided $t$-critical value at level $\alpha$, i.e. $P(|T_{q-1}|>cv_{q}(\alpha))=\alpha$ for $T_{q-1}\sim t_{q-1}$ (one-sided tests are analogous).
To study the asymptotic behavior of $\{\check{\gamma}_{j}\}_{j=1}^{q}$, we decompose
$\check{\gamma}_{j}=\zeta_{j}+\psi_{j},$
where
$\zeta_{j}=\beta\sqrt{\frac{q}{T}}\sum_{t\in\mathcal{G}_{j}}|x_{t-1}|,\qquad\psi_{j}=\sqrt{\frac{q}{T}}\sum_{t\in\mathcal{G}_{j}}\text{sign}(x_{t-1})u_{t}.$
Under Assumption 2.1, $\{\psi_{j}\}_{j=1}^{q}$ forms a sequence of martingale differences uncorrelated across groups, yielding the following asymptotic characterization.
Lemma 3.1.
Let Assumptions 2.1, 2.2, and 2.3 hold. For any fixed $q\geq 2$ and $\beta\in\mathbb{R}$,
$(\check{\gamma}_{1}-\zeta_{1},\dots,\check{\gamma}_{q}-\zeta_{q})^{\prime}\to_{d}\mathbb{MN}\bigl(0,\mathrm{diag}(q\omega_{1}^{2},\dots,q\omega_{q}^{2})\bigr),$
where $\omega_{j}^{2}=\int_{(j-1)/q}^{j/q}\sigma^{2}(r)\,dr$ for $j=1,\dots,q$.
The statistics $\{\check{\gamma}_{j}\}_{j=1}^{q}$ do not satisfy the standard condition in Ibragimov and Müller (2010), which requires estimators $\{\tilde{\beta}_{j}\}_{j=1}^{q}$ such that
$\{m_{T}(\tilde{\beta}_{j}-\beta)\}_{j=1}^{q}\to_{d}\{V_{j}Z_{j}\}_{j=1}^{q},$
for some $m_{T}\to\infty$, $Z_{j}\stackrel{{\scriptstyle iid}}{{\sim}}\mathbb{N}(0,1)$, and $\{V_{j}\}$ independent of $\{Z_{j}\}$. By contrast, Lemma 3.1 shows that $\{\check{\gamma}_{j}\}_{j=1}^{q}$ lack such a diverging normalization. Consequently, as shown in Proposition 3.2, the $t$-statistic approach yields correct asymptotic size but is consistent only for a restricted class of covariates, excluding (near) unit-root processes. This inconsistency arises precisely because the asymptotics of $\{\check{\gamma}_{j}\}_{j=1}^{q}$ do not involve a diverging sequence (see proofs of Proposition 3.2 and Corollary 3.3).
Nevertheless, with additional regularity conditions, if $\{c_{T}^{-1}\sum_{t\in\mathcal{G}_{j}}|x_{t-1}|\}_{j=1}^{q}\to_{d}\{D_{j}\}_{j=1}^{q}$ for positive random variables $\{D_{j}\}$ and a sequence $c_{T}/\sqrt{T}\to\infty$, then the Cauchy estimator $\check{\beta}_{j}$ computed within each group satisfies
$\{m_{T}(\check{\beta}_{j}-\beta)\}_{j=1}^{q}\to_{d}\{P_{j}\}_{j=1}^{q},$
for $m_{T}=c_{T}\sqrt{q/T}$. In general, however, $\{P_{j}\}_{j=1}^{q}$ are non-Gaussian, especially when $(x_{t})$ is (near) unit root and endogenous. Applying the $t$-statistic approach to $\{\check{\beta}_{j}\}_{j=1}^{q}$ thus yields consistency for broader classes of covariates but may incur size distortions due to non-Gaussianity.
Proposition 3.2.
Let Assumptions 2.1, 2.2, and 2.3 hold, with fixed $q\geq 2$ and $\alpha\leq 0.83$.
(a) Under $H_{0}:\beta=0$,
$\lim_{T\to\infty}\mathbb{P}(|t_{q}(\check{\gamma})|>cv_{q}(\alpha)\mid H_{0})\leq\alpha.$
(b) Under $H_{A}:\beta\neq 0$, suppose $(x_{t})$ is stationary with $E|x_{t}|<\infty$ and satisfies
$\sup_{1\leq s\leq T-T/q}\Biggl|E|x_{t}|-\frac{1}{T/q}\sum_{t=s}^{s+T/q}|x_{t}|\Biggr|\to_{p}0.$
Then
$\lim_{T\to\infty}\mathbb{P}(|t_{q}(\check{\gamma})|>cv_{q}(\alpha)\mid H_{A})=1.$
Proposition 3.2 shows that the $t$-statistic approach is conservative under $H_{0}$ and consistent under $H_{A}$ when $(x_{t})$ is stationary with a finite first moment. It is thus valid and robust to persistent heteroskedasticity and endogenously heavy-tailed covariates. However, it becomes inconsistent for highly persistent covariates, such as (near) unit-root processes. To illustrate, consider the generalized local-to-unity framework of Dou and Müller (2021), where $X^{T}(r)=x_{[Tr]}$ for $r\in[0,1]$ and
$T^{-1/2}\bigl(X^{T}(\cdot)-X^{T}(0)\bigr)\to_{d}X(\cdot)-X(0),$ (5)
with $X$ a stationary continuous-time Gaussian ARMA process. [^6] See Dou and Müller (2021) for a detailed discussion.
Corollary 3.3.
Let Assumptions 2.12.3 hold and suppose $(x_{t})$ satisfies (5). Under $\beta\neq 0$, $t_{q}(\check{\gamma})\to_{d}\text{sign}(\beta)\mathcal{D}_{q}$ for $q\geq 2$, where
$\mathcal{D}_{q}=\int_{0}^{1}|X(r)|dr\left(\frac{q(q-1)}{\sum_{j=1}^{q}\bigl(\int_{0}^{1}|X(r)|dr-q\int_{(j-1)/q}^{j/q}|X(r)|dr\bigr)^{2}}\right)^{1/2},$
and $\mathcal{D}_{q}>(q-1)^{-1/2}$ with probability one.
When $(x_{t})$ is highly persistent, $t_{q}(\check{\gamma})$ converges to $\mathcal{D}_{q}$ rather than diverging, with lower bound $(q-1)^{-1/2}$. Simulations in Section 5 confirm that rejection probabilities remain high even when $t_{q}(\check{\gamma})$ is asymptotically bounded. For $q=2$,
$\mathcal{D}_{2}=\frac{\int_{0}^{1}|X(r)|dr}{\bigl|\int_{0}^{1/2}|X(r)|dr-\int_{1/2}^{1}|X(r)|dr\bigr|}>1.$ (6)
The ratio form in (6) implies large realizations of $\mathcal{D}_{2}$ in finite samples, producing high rejection rates even under inconsistency. Figure 1 plots the simulated density of $\mathcal{D}_{2}$ when $X$ is Brownian motion. [^7] Based on 100,000 simulated draws. The minimum simulated value is 1.15, and $\mathbb{P}(|\mathcal{D}_{2}|>cv_{2}(0.05))=0.15$ with $cv_{2}(0.05)=4.303$.
Figure 1: Simulated density of $\mathcal{D}_{2}$ in (6).
3.2 A Hybrid Test
We now propose a simple hybrid test that remains consistent for a broad class of covariates. Under Assumptions 2.12.3,
$\frac{\sum_{t=1}^{T}|x_{t-1}|}{\sqrt{T}}(\check{\beta}-\beta)=\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\text{sign}(x_{t-1})u_{t}\to_{d}\int_{0}^{1}\sigma(r)\,dW(r)=\omega Z,$
where $Z\sim\mathbb{N}(0,1)$ and $\omega^{2}=\int_{0}^{1}\sigma^{2}(r)\,dr$.
A key feature of the Cauchy estimator $\check{\beta}$ is that its properly normalized limit distribution is invariant to the data-generating process of $(x_{t})$. By contrast, the OLS estimator’s variance depends on both $(u_{t})$ and $(x_{t})$, complicating variance estimation even under homoskedasticity. For $\check{\beta}$, the asymptotic variance depends solely on $u_{t}$, requiring only heteroskedasticity-robust adjustments. [^8] See Shephard (2020), Section 4.3, for related discussion.
We define the hybrid test statistic as
$\tau(\check{\beta})=\frac{\check{\gamma}}{\hat{\omega}}$,
where $\check{\gamma}=(\sum_{t=1}^{T}|x_{t-1}|/\sqrt{T})\check{\beta}$ as in (2), and
$\hat{\omega}^{2}=\frac{1}{T}\sum_{t=1}^{T}\hat{u}_{t}^{2},\qquad\hat{u}_{t}=y_{t}-\hat{\beta}x_{t-1}.$
Here, $\hat{\omega}^{2}$ estimates $\omega^{2}=\int_{0}^{1}\sigma^{2}(r)\,dr$ using OLS residuals. As noted by Shephard (2020), the Cauchy-based variance estimator performs poorly because the Cauchy estimator converges more slowly and less efficiently than OLS when $(x_{t})$ is heavy-tailed or nearly integrated. Hence, we use OLS residuals to improve efficiency. [^9] A related approach is employed by Kostakis et al. (2015) in the IVX framework of Phillips and Magdalinos (2007a).
We assume:
Assumption 3.1.
$\sum_{t=1}^{T}x_{t-1}u_{t}=o_{p}\!\left(\sqrt{T\sum_{t=1}^{T}x_{t-1}^{2}}\right)$.
Assumption 3.1 is very general and holds in many time-series settings. It is weaker than Assumption 3.2 of Ibragimov et al. (2024), which requires $\sum_{t=1}^{T}x_{t-1}u_{t}=O_{p}\!\left(\sqrt{T^{p}\sum_{t=1}^{T}x_{t-1}^{2}}\right)$ for $p\in[0,1/16)$. As shown in Ibragimov et al. (2024), this holds with $p=0$ when $(x_{t})$ is either (near) unit root or stationary with finite variance; it also applies to certain stationary heavy-tailed processes (see, e.g., Samorodnitsky et al., 2007).
Under Assumption 3.1,
$|\hat{\beta}-\beta|=o_{p}\!\left(\sqrt{T\big/\sum_{t=1}^{T}x_{t-1}^{2}}\right),\quad\text{and hence}\quad\hat{\omega}^{2}\to_{p}\omega^{2}.$
The asymptotic properties of $\tau(\check{\beta})$ follow.
Proposition 3.4.
Let Assumptions 2.1, 2.2, 2.3, and 3.1 hold.
(a) Under $H_{0}:\beta=0$,
$\tau(\check{\beta})\to_{d}\mathbb{N}(0,1).$
(b) Under $H_{A}:\beta\neq 0$,
$\tau(\check{\beta})=\beta\frac{\sum_{t=1}^{T}|x_{t-1}|}{\omega\sqrt{T}}(1+o_{p}(1))+O_{p}(1),$
so $|\tau(\check{\beta})|\to_{p}\infty$ whenever $\sum_{t=1}^{T}|x_{t-1}|/\sqrt{T}\to_{p}\infty$.
The conclusions of Proposition 3.4 remain valid under weaker conditions. For instance, if Assumptions 2.1 and 3.1 hold and
$\frac{1}{T}\sum_{t=1}^{T}v_{t}^{2}\to_{p}\omega^{2}>0,\qquad\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\text{sign}(x_{t-1})u_{t}\to_{d}\omega Z,$
where $Z\sim\mathbb{N}(0,1)$ is independent of $\omega^{2}$, then $\tau(\check{\beta})$ retains its asymptotic validity. These conditions include stationary volatility with $E[v_{t}^{2}]=\omega^{2}$. Hence, Assumptions 2.2 and 2.3 can be interpreted as primitive sufficient conditions accommodating persistent volatility in predictive regression data.
Remark. Proposition 3.4(a) also holds if $\tau(\check{\beta})$ uses $\bar{\omega}^{2}=T^{-1}\sum_{t=1}^{T}y_{t}^{2}$ instead of $\hat{\omega}^{2}$, since $\beta=0$ under $H_{0}$. Moreover, the corresponding test remains consistent when $(x_{t})$ is stationary with finite variance or follows a generalized local-to-unity process (Dou and Müller, 2021). However, it can be inconsistent for heavy-tailed $(x_{t})$. For instance, if $(x_{t})$ is i.i.d. $\alpha$-stable with $\alpha\in(0,2)$ and independent of $(u_{t})$, then
$\bar{\omega}^{2}=\beta^{2}\left(\frac{1}{T}\sum_{t=1}^{T}x_{t-1}^{2}\right)(1+o_{p}(1)),\quad\tau(\check{\beta})=\frac{\sum_{t=1}^{T}|x_{t-1}|}{\sqrt{\sum_{t=1}^{T}x_{t-1}^{2}}}(1+o_{p}(1))=O_{p}(1),$
by the generalized central limit theorem (see Feller, 1971; Logan et al., 1973; Davis, 1983; Davis and Resnick, 1986). Thus, the use of $\hat{\omega}^{2}$ (or another consistent estimator under both $H_{0}$ and $H_{A}$) is crucial for the consistency of the hybrid test.
4 Extensions
This section extends the inference methods developed in Section 3 to models with multiple predictors and to regressions including an intercept. Our goal is not to design efficient procedures but to provide simple and robust inference methods that rely on minimal assumptions on the predictors and volatility processes.
4.1 Predictive Regressions with Multiple Predictors
Consider a predictive regression with $K$ predictors $x_{t}=[x_{1,t},\ldots,x_{K,t}]^{\prime}$:
$y_{t}$ $=x_{t-1}^{\prime}B+u_{t}$
$=\beta_{1,K}x_{1,t-1}+\cdots+\beta_{K,K}x_{K,t-1}+u_{t},\qquad B=[\beta_{1,K},\ldots,\beta_{K,K}]^{\prime}.$ (7)
The objective is to test the joint predictability of the covariates, that is,
$H_{0}:\beta_{1,K}=\cdots=\beta_{K,K}=0.$
We construct a testing procedure for $H_{0}$ based on the univariate inference methods in Section 3. Specifically, we estimate $K$ univariate predictive regressions
$y_{t}=\beta_{k}x_{k,t-1}+u_{k,t},\qquad k=1,\ldots,K,$
and test each null hypothesis
$H_{0}^{(k)}:\beta_{k}=0,\qquad k=1,\ldots,K.$
Clearly, $H_{0}$ implies $H_{0}^{(k)}$ for all $k$. The converse also holds under mild regularity conditions, as shown below.
Lemma 4.1.
Consider model (S.1) and define $z_{t}=[\operatorname{sign}(x_{1,t}),\ldots,\operatorname{sign}(x_{K,t})]^{\prime}$. Suppose that for each $t\in\{1,\dots,T\}$, $E[z_{t-1}u_{t}]=\mathbf{0}_{K\times 1}$, $0<E[|x_{k,t-1}|]<\infty$ for all $k$, and that the matrix $E[z_{t-1}x_{t-1}^{\prime}]$ is invertible. [^10] Even when $\operatorname{sign}(x_{t-1})$ is constant, the univariate Cauchy estimator and associated tests remain well defined. In this case, the estimator simplifies to $\check{\beta}=\sum_{t=1}^{T}y_{t}\big/\sum_{t=1}^{T}x_{t-1}$, implying $\check{\beta}-\beta=\sum_{t=1}^{T}u_{t}\big/\sum_{t=1}^{T}x_{t-1}$, and Proposition 3.4 continues to hold. Hence, the lack of sign variation does not affect the validity of our methods in the univariate case. In the multiple-predictor case, however, the invertibility of $E[z_{t-1}x_{t-1}^{\prime}]$ imposes mild restrictions on sign changes across predictors. For practical applications, one can induce variation in the sign instrument by recentering the predictor, for example, $\tilde{x}_{t-1}=x_{t-1}-t^{-1}\sum_{s=1}^{t}x_{s-1}$, which preserves both the martingale structure and the asymptotic validity of the estimator. Then, the joint null hypothesis $H_{0}$ holds if and only if $H_{0}^{(k)}$ holds for all $k=1,\ldots,K$.
Lemma 4.1 justifies the use of multiple hypothesis testing based on univariate Cauchy estimators. [^11] See Harvey et al. (2015) for an application of the multiple-testing framework in predictive regressions, and Kostakis et al. (2015) for joint-predictability tests in the IVX framework. Note that the IVX approach may lose validity under heavy-tailed predictors or continuous-time data, whereas our method remains robust in such settings. In conjunction with the hybrid test introduced in Section 3.2, we compute the statistic $\tau(\check{\beta}_{k})$ for each parameter $\beta_{k}$, where $\check{\beta}_{k}$ denotes the corresponding Cauchy estimator. Let $p_{k}$ denote its $p$-value. The joint null hypothesis $H_{0}$ is rejected at level $\alpha$ if $\min_{k}p_{k}\leq\alpha/K$, following the Bonferroni correction.
This approach directly extends the univariate robust inference procedure to a multivariate setting and requires only mild conditions for the equivalence between $H_{0}$ and $\{H_{0}^{(k)}\}_{k=1}^{K}$. The Bonferroni correction imposes no assumptions on the joint distribution of the test statistics, which motivates its use here (see Holm, 1979; Benjamini and Hochberg, 1995; Shaffer, 1995).
We also note that under some additional regularity conditions, the joint hypothesis can be tested directly using a Wald-type statistic:
$W=\left(\sum_{t=1}^{T}z_{t-1}y_{t}\right)^{\prime}\left(\hat{\omega}^{2}\sum_{t=1}^{T}z_{t-1}z_{t-1}^{\prime}\right)^{-1}\left(\sum_{t=1}^{T}z_{t-1}y_{t}\right).$
In particular, under $H_{0}$,
$\left(\hat{\omega}^{2}\sum_{t=1}^{T}z_{t-1}z_{t-1}^{\prime}\right)^{-1/2}\left(\sum_{t=1}^{T}z_{t-1}y_{t}\right)\to_{d}\mathbb{N}(0,I_{K}),$
and hence $W\to_{d}\chi_{K}^{2}$. [^12] As mentioned earlier, $z_{t}$ can be interpreted as an instrument. Therefore, one may use an alternative instrument, as in Shephard (2020), and construct a Wald-type test accordingly. We provide a detailed analysis of the Wald-type test in the Online Appendix, Section A.1.
While both the Wald test and the Bonferroni-adjusted marginal tests evaluate the same null hypothesis, they are not equivalent procedures except under special cases such as asymptotic orthogonality of the regressors. The Wald test exploits the full covariance structure and yields an elliptical rejection region, whereas the Bonferroni procedure yields a rectangular region and controls family-wise error. Consequently, the procedures may differ in finite-sample and asymptotic power when regressors are correlated. [^13] Again, our goal is not to design efficient procedures but to provide simple and robust methods that rely on minimal assumptions. We leave a systematic comparison between the Bonferroni-type multiple-testing procedure and the Wald-type joint test for future research.
4.2 Predictive Regressions with an Intercept
The analysis in Section 3 assumes that the intercept $\alpha=0$ in (1). When $\alpha\neq 0$, it must be properly accounted for. A natural starting point is the demeaned model
$y_{t}-\bar{y}_{T}=\beta(x_{t-1}-\bar{x}_{T})+u_{t}-\bar{u}_{T},\qquad t=1,\ldots,T,$ (8)
where $\bar{z}_{s}=s^{-1}\sum_{t=1}^{s}z_{t}$ for $z_{t}\in\{y_{t},x_{t-1},u_{t}\}$. However, $(u_{t}-\bar{u}_{T})$ is not a martingale difference sequence (MDS) with respect to $(\mathcal{F}_{t})$, invalidating the martingale CLT used in Sections 2 and 3. Specifically, the Cauchy estimator becomes
$\check{\beta}-\beta=\Biggl(\sum_{t=1}^{T}|x_{t-1}-\bar{x}_{T}|\Biggr)^{-1}\sum_{t=1}^{T}\text{sign}(x_{t-1}-\bar{x}_{T})(u_{t}-\bar{u}_{T}),$
which is problematic because: (i) $u_{t}-\bar{u}_{T}$ is not an MDS, and (ii) $\text{sign}(x_{t-1}-\bar{x}_{T})$ is not $\mathcal{F}_{t-1}$-measurable. Thus, the theory in Section 3 is not directly applicable. [^14] Recursive demeaning using $\bar{y}_{t}$ instead of $\bar{y}_{T}$ does not resolve this issue since $u_{t}-\bar{u}_{t}$ is not an MDS either.
To restore the MDS property, we instead difference the model:
$y_{t}-y_{t-1}=\beta(x_{t-1}-x_{t-2})+(u_{t}-u_{t-1}),$
and estimate this first-differenced (FD) model on alternating subsets of observations. [^15] Differencing $y_{t}$ to eliminate the nonzero mean is a widely used technique in the analysis of dynamic panel data with individual heterogeneity. See, among others, Anderson and Hsiao (1981); Anderson and Hsiao (1982); Arellano and Bond (1991). We focus on the even-indexed observations and define the modified Cauchy estimator:
$\check{\beta}_{e}=(D_{T}^{e})^{-1}\sum_{t=2}^{T/2}\text{sign}(x_{2t-2})(y_{2t}-y_{2t-1}),\quad D_{T}^{e}=\sum_{t=2}^{T/2}\text{sign}(x_{2t-2})(x_{2t-1}-x_{2t-2}).$
This estimator has two key properties. First, for even-indexed data, the regression error $u_{t}^{e}=u_{2t}-u_{2t-1}$ forms an MDS with respect to $\mathcal{F}_{t}^{e}:=\mathcal{F}_{2t}$ for $t=1,\ldots,T/2$. [^16] For odd-indexed data, $u_{t}^{o}=u_{2t+1}-u_{2t}$ forms an MDS with respect to $\mathcal{F}_{t}^{o}:=\mathcal{F}_{2t+1}$, yielding an analogous estimator $\check{\beta}_{o}$. Second, $\check{\beta}_{e}$ can again be viewed as an IV estimator, but it uses $\text{sign}(x_{t-2})$, which is $\mathcal{F}_{t-2}$-measurable, as the instrument. [^17] More generally, one may use $\text{sign}\!\left(\sum_{l\leq 2}w_{l}x_{t-l}\right)$ for deterministic weights $\{w_{l}\}$, provided $E[\text{sign}(\sum_{l\leq 2}w_{l}x_{t-l})(x_{t-1}-x_{t-2})]\neq 0$. Hence, $\text{sign}(x_{2t-2})(u_{2t}-u_{2t-1})$ is an MDS with respect to $(\mathcal{F}_{t}^{e})$.
The inference procedures of Section 3 remain valid for $\check{\beta}_{e}$. In particular, the hybrid test in Section 3.2 can be implemented as
$\tau(\check{\beta}_{e})=\frac{\check{\gamma}_{e}}{\hat{\omega}}$, (9)
where $\check{\gamma}_{e}=D_{T}^{e}\check{\beta}_{e}/\sqrt{T/2}$, and
$\hat{\omega}^{2}=\frac{1}{T}\sum_{t=1}^{T}\hat{u}_{t}^{2},\qquad\hat{u}_{t}=(y_{t}-\bar{y}_{T})-\hat{\beta}(x_{t-1}-\bar{x}_{T}),$
with $\hat{\beta}$ the OLS estimator from the demeaned model (8). Note that $\hat{\omega}^{2}$ is based on the full sample, whereas $\check{\beta}_{e}$ uses only even-indexed data. The asymptotic validity of this hybrid procedure is established next.
Corollary 4.2.
Let Assumptions 2.1, 2.2, and 2.3 hold, and suppose Assumption 3.1 holds with $x_{t-1}$ replaced by $x_{t-1}-\bar{x}_{T}$.
(a) Under $H_{0}:\beta=0$,
$\tau(\check{\beta}_{e})\to_{d}\mathbb{N}(0,1).$
(b) Under $H_{A}:\beta\neq 0$,
$\tau(\check{\beta}_{e})=\beta\,\frac{\sum_{t=1}^{T/2}\text{sign}(x_{2t-2})(x_{2t-1}-x_{2t-2})}{\omega\sqrt{T/2}}(1+o_{p}(1))+O_{p}(1),$
so that $|\tau(\check{\beta}_{e})|\to_{p}\infty$ whenever $\bigl|\sum_{t=1}^{T/2}\text{sign}(x_{2t-2})(x_{2t-1}-x_{2t-2})\bigr|/\sqrt{T/2}\to_{p}\infty$.
Although the odd-indexed estimator $\check{\beta}_{o}$ has analogous properties, $\check{\beta}_{e}$ and $\check{\beta}_{o}$ are typically dependent, with the dependence structure determined by the DGP of $(x_{t})$. Hence, unless additional assumptions are imposed, we restrict attention to a single subset of observations—either with even or odd indices. [^18] Using only half of the data is not uncommon in predictive regressions. See, for example, Zhu et al. (2014) and Liu et al. (2019), who employ long-lag differencing to eliminate intercepts. In addition, Dufour and Torrès (2000) uses a split-sample approach to address inference problems under a Markovian structure.
Consistency of the hybrid test with an intercept requires
$\frac{1}{\sqrt{T/2}}\sum_{t=1}^{T/2}\operatorname{sign}(x_{2t-2})(x_{2t-1}-x_{2t-2})\to_{p}\infty.$
This holds for most stationary processes $(x_{t})$ if
$E[\operatorname{sign}(x_{t-1})x_{t}]\neq E[|x_{t-1}|].$
The condition may fail for certain unit-root processes. For instance, for a random walk $x_{t}=x_{t-1}+\varepsilon_{t}^{x}$, it does not hold. More generally, in the local-to-unity model of Phillips and Magdalinos (2007b),
$x_{t}=\left(1+\frac{c}{T^{\delta}}\right)x_{t-1}+\varepsilon_{t}^{x},\qquad c<0,\quad\delta\in[0,1],$ (10)
with $\varepsilon_{t}^{x}$ satisfying Assumption LP therein, the consistency condition becomes
$\frac{1}{T^{1/2}}\sum_{t=1}^{T/2}\operatorname{sign}(x_{2t-2})(x_{2t-1}-x_{2t-2})$ $=\frac{c}{T^{1/2+\delta}}\sum_{t=1}^{T/2}|x_{2t-2}|+\frac{1}{T^{1/2}}\sum_{t=1}^{T/2}\operatorname{sign}(x_{2t-2})\varepsilon_{2t-1}^{x}.$
The first term diverges if and only if $\delta<1$ (see Lemma 3.2 of Phillips and Magdalinos, 2007b). Since $\varepsilon_{t}^{x}$ and $x_{t-1}$ may be dependent, it is generally the case that $E[\operatorname{sign}(x_{2t-2})\varepsilon_{2t-1}^{x}]\neq 0$, implying that the second term also diverges. Consequently, inconsistency arises only when $\delta=1$ and $E[\operatorname{sign}(x_{2t-2})\varepsilon_{2t-1}^{x}]=0$. In all other cases (i.e., $0\leq\delta<1$ or when the expectation is nonzero), the test remains consistent.
Moreover, the robustness to a nonzero intercept is not without cost. Specifically, the test based on $\tau(\check{\beta}_{e})$ is less powerful than that based on $\tau(\check{\beta})$ when the intercept is zero. A detailed power comparison is provided in Section A.2 of the Online Appendix.
5 Finite Sample Performance
This section investigates the finite-sample performance of the proposed methods. Two sets of simulation experiments are conducted. The first set is based on a continuous-time model and compares our robust $t$-statistic–based tests, $t_{q}(\check{\gamma})$ for $q\in\{8,12,16\}$, and the hybrid test $\tau(\check{\beta})$ with the Cauchy RT test of Choi et al. (2016) and the Cauchy VC test of Ibragimov et al. (2024). In this setting, the intercept is suppressed in the test statistics. The second set is based on a discrete-time predictive regression model and compares our procedures, $\tau(\check{\beta}_{e})$ and $\tau(\check{\beta}_{o})$, with the IVX test of Kostakis et al. (2015). In this setting, the statistics are computed with the intercept included in the estimation. [^19] The number of replications is 10,000 for each setting.
5.1 Continuous-Time Experiments
5.1.1 Simulation Design
Following Choi et al. (2016) and Ibragimov et al. (2024), we consider the continuous-time predictive regression model
$dY_{t}$ $=\beta X_{t}\,dt+dU_{t},$ (11)
$dX_{t}$ $=-\frac{\bar{\kappa}}{T}X_{t}\,dt+\sigma_{t}\,dV_{t},\qquad dU_{t}=\sigma_{t}\left(dW_{t}+\int_{\mathbb{R}}x\,\Lambda(dt,dx)\right),$
where $V_{t}$ and $W_{t}$ are Brownian motions with $E(V_{t}W_{t})=-0.98t$. The constant term in the predictive regression is set to zero, and recursive demeaning is applied as Choi et al. (2016). The model is observed at interval $\Delta=1/252$, corresponding to daily observations, so that a sample of length $T$ years contains $252T$ observations.
The volatility process $\sigma_{t}$ follows one of the following specifications:
  • CNST (Constant vvolatility): $\sigma_{t}^{2}=\sigma_{0}^{2}$, with $\sigma_{0}=1$.
  • SB (Structural Break): $\sigma_{t}=\sigma_{0}+(\sigma_{1}-\sigma_{0})1\{t/T\geq 4/5\}$, with $\sigma_{0}=1$ and $\sigma_{1}=4$.
  • RS (Regime Switching): $\sigma_{t}=\sigma_{0}(1-s_{t})+\sigma_{1}s_{t}$, where $s_{t}$ is a two-state Markov chain independent of $(Y_{t},X_{t})$, with transition matrix
    $P_{t}=\begin{pmatrix}0.8&0.2\\ 0.8&0.2\end{pmatrix}+\begin{pmatrix}0.2&-0.2\\ -0.8&0.8\end{pmatrix}\exp\!\left(-\frac{\bar{\lambda}}{T}t\right),$
    initialized at its invariant distribution, where $\bar{\lambda}=60$, $\sigma_{0}=1$, and $\sigma_{1}=4$.
  • GBM (Geometric Brownian Motion): $d\sigma_{t}^{2}=\frac{1}{2}\frac{\bar{\omega}^{2}}{T}\sigma_{t}^{2}\,dt+\frac{\bar{\omega}^{2}}{\sqrt{T}}\sigma_{t}^{2}\,dZ_{t}$, where $Z_{t}$ is a Brownian motion correlated with $W_{t}$ such that $E(W_{t}Z_{t})=-0.4t$ and $\bar{\omega}=9$.
Following Choi et al. (2016), we set $T\in\{20,50,100\}$ (corresponding to 240, 600, and 1200 monthly observations) and $\bar{\kappa}\in\{0,5,10\}$ for the persistence parameter in (11). We consider a two-sided test of $H_{0}:\beta=0$ against $H_{A}:\beta\neq 0$.
5.1.2 Results
We first assess the empirical size of each test under $H_{0}:\beta=0$. The results for the four volatility models (CNST, SB, RS, and GBM) are reported in Table 1. Overall, both the $t$-statistic–based tests and the hybrid method exhibit satisfactory size performance, closely matching the nominal levels and performing comparably to the Cauchy RT and Cauchy VC tests. Among the $t$-based procedures, moderate partition numbers ($q=12$ or $16$) provide the most stable results, whereas smaller $q$ values tend to be mildly undersized. In the GBM case, where volatility is endogenously persistent, the $t$-statistic–based tests become slightly conservative but remain competitive with the Cauchy RT and VC methods.
Next, we analyze the finite-sample power properties of the tests. We consider $\beta\in\{0.004k,k=1,\cdots,5\}$ under the same volatility specifications. The results are summarized in Tables 2S.8. The proposed tests exhibit power comparable to that of the Cauchy RT and Cauchy VC procedures. For small samples ($T=20$), the Cauchy RT and VC tests occasionally show higher power, but the difference diminishes as $T$ increases. In certain settings, our methods even outperform the existing approaches. For instance, $t_{16}(\check{\gamma})$ dominates under $\beta=0.02$, $\bar{\kappa}=0.5$, and regime switching volatility (Table 4), whereas the hybrid test $\tau(\check{\beta})$ performs best under $\beta=0.004$, $\bar{\kappa}=20$, $T=20$, and regime switching volatility.
In summary, all four robust inference procedures—Cauchy RT, Cauchy VC, $t_{q}(\check{\gamma})$, and $\tau(\check{\beta})$—deliver accurate size control and strong discriminatory power under endogenously persistent regressors and persistent volatility. While the Cauchy RT requires high-frequency data and a time transformation, and the Cauchy VC involves nonparametric volatility filtering with a tuning parameter, our proposed $t$-statistic and hybrid methods are much simpler to implement and require neither. Hence, these approaches are best viewed as complementary: the Cauchy RT and Cauchy VC are preferable in high-frequency environments, whereas our procedures provide robust and easily implementable alternatives in more general settings. It is also worth emphasizing that the proposed methods, like the Cauchy VC, are applicable to both continuous- and discrete-time models, whereas the Cauchy RT method is restricted to the continuous-time framework.
5.2 Discrete-Time Experiments
5.2.1 Simulation Design
We now examine the finite-sample performance of the proposed tests in a discrete-time setting with an intercept, comparing them to the IVX test of Kostakis et al. (2015). The DGP is specified as
$y_{t}$ $=\beta x_{t-1}+\sigma_{t}\varepsilon_{t},$ (12)
$x_{t}$ $=\left(1-\frac{\bar{\kappa}}{T}\right)x_{t-1}+\sigma_{t}\eta_{t},$
for $t=2,\ldots,12T$, where $T\in\{20,50,100\}$ corresponds to 20, 50, and 100 years of monthly data (i.e., 240, 600, and 1200 monthly observations, respectively). The constant term in the predictive regression is set to zero. We set $\beta\in\{0.5k:k=0,1,\ldots,5\}$ and $\bar{\kappa}\in\{0,50,100\}$, and consider a one-sided test of $H_{0}\!:\beta=0$ against $H_{A}\!:\beta>0$. [^20] The IVX test of Kostakis et al. (2015) performs well in two-sided testing for a broad class of models with constant volatility. However, as shown in Demetrescu et al. (2023), the IVX method exhibits severe size distortions in one-sided tests when regressors are highly persistent and endogenous. For this reason, we focus on the one-sided case to highlight the performance of our methods in this setting. Simulations for two-sided and joint hypothesis tests are also reported in Section B of the Online Appendix. Notably, while the IVX approach demonstrates reliable size control under constant volatility, it exhibits significant size distortions in the presence of non-constant volatility. In contrast, our methods remain robust to such heteroskedasticity.
The innovation process $\eta_{t}$ follows an MA($q$) process:
$\eta_{t}=\sum_{j=0}^{q}C_{j}v_{t-j},$ (13)
where $(\varepsilon_{t},v_{t})$ are jointly normal with correlation 0.98-0.98. In our simulations, we consider an MA(1) process with $C_{0}=C_{1}=1/\sqrt{2}$. [^21] Additional simulation results with an MA(3) process are reported in the Online Appendix, Section B.1. Moreover, as discussed in Section 4.2, the intercept-robust version of our test, $\tau(\check{\beta}_{e})$, becomes inconsistent when $x_{t}$ follows the local-to-unity process (10) with $\delta=1$ and $\eta_{t}$ is serially independent. For this reason, we do not consider i.i.d. $\eta_{t}$ in our simulations. See the Online Appendix, Section A.2, for a detailed theoretical power analysis of the test statistic $\tau(\check{\beta}_{e})$. The volatility process $\sigma_{t}$ follows the same specifications as in the continuous-time simulations, except that the GBM model is excluded.
We implement the hybrid tests based on the even and odd observations, denoted by $\tau(\check{\beta}_{e})$ and $\tau(\check{\beta}_{o})$, respectively (see (9)). For comparison, we also include the IVX test of Kostakis et al. (2015), in which the intercept is included in the estimation.
5.2.2 Results
The results, summarized in Tables S.98, indicate that the proposed tests exhibit excellent size control under the null hypothesis across all DGPs, whereas the IVX test is substantially oversized, particularly when volatility is nonstationary or exhibits structural breaks. Furthermore, the hybrid approaches demonstrate nontrivial power, even though they are constructed using only half of the observations.
Overall, these findings corroborate the theoretical robustness of our methods. They remain valid under heavy-tailed, endogenous, and persistent regressors, as well as under heteroskedastic and persistent volatility. In contrast, the IVX test exhibits severe size distortions in one-sided tests across all cases, including the constant volatility setting, as documented by Demetrescu et al. (2023).
We also conduct two-sided tests and report the results in the Online Appendix (Section B.2). In this case, the IVX method exhibits good size performance under constant volatility but suffers from severe size distortions in settings with time-varying volatility, whereas our methods maintain accurate size control across all specifications.
Taken together, these results suggest that the proposed robust procedures provide a practical and reliable alternative to existing inference methods for predictive regressions in both continuous- and discrete-time frameworks.
6 Empirical Application
To illustrate the empirical performance of the proposed tests relative to the Cauchy RT and Cauchy VC tests, we reexamine the dataset used by Choi et al. (2016) to test the predictability of stock returns using the dividend–price (D/P) and earnings–price (E/P) ratios as predictors. [^22] Further empirical results, including intercept-robust estimations and Wald-type tests compared to IVX (Kostakis et al., 2015), are detailed in the Online Appendix C. For stock returns, we employ the NYSE/AMEX value-weighted index and the S&P 500 index obtained from the Center for Research in Security Prices (CRSP). The dividend–price ratio is defined as the annual dividend divided by the current total market value. Further details on data construction are provided in Section 6.1 of Choi et al. (2016).
Following Choi et al. (2016), we estimate two types of predictive regressions: one based on all returns and another based only on returns generated from the diffusive component of stock prices, obtained by first testing for jumps and removing observations corresponding to detected jumps. In all cases, we apply one-sided tests.
[…]
Consistent with our simulation evidence, the Cauchy RT test demonstrates strong finite-sample power but requires high-frequency data due to its reliance on a continuous-time approximation. [^23] For the Cauchy RT test in our simulations, we estimate the discretized time-changed regression using $n$ lower-frequency observations, with $n=12T$ (approximately monthly), generated by the random time-sampling scheme described in Section 5 of Choi et al. (2016). The mixed empirical results—where the Cauchy RT test fails to reject the null while both the proposed methods and the Cauchy VC test do reject—may reflect the limited accuracy of the continuous-time approximation when applied to monthly, quarterly, or yearly data. Evaluating the robustness of the continuous-time approximation underlying the Cauchy RT test remains an interesting topic for future research.
7 Conclusion
This paper introduces two robust inference methods for predictive regressions, addressing key econometric challenges commonly encountered in empirical finance, such as endogenously persistent or heavy-tailed regressors and persistent volatility in errors. Building on the Cauchy estimation framework, we develop two simple yet theoretically rigorous procedures: a $t$-statistic–based approach and a hybrid method. Both methods are computationally straightforward and applicable to continuous- and discrete-time models alike.
Simulation evidence demonstrates that the proposed tests perform well in finite samples, maintaining correct size and competitive power under a wide range of data-generating processes, including those characterized by stochastic volatility, structural breaks, and regime switching. Although our procedures require the assumption of asymptotically exogenous volatility, they exhibit excellent robustness and complement existing methods, including the IVX method of Kostakis et al. (2015), the Cauchy RT test of Choi et al. (2016), and the Cauchy VC test of Ibragimov et al. (2024).
In an empirical application to stock return predictability, we use the dividend–price and earnings–price ratios as predictors for excess returns on major U.S. equity indices. The results indicate that the dividend–price ratio possesses predictive power, while the earnings–price ratio does not significantly forecast returns. Overall, the proposed inference procedures offer a practical, theoretically sound, and implementable alternative to existing methods for robust inference in predictive regressions.
Appendix: Proofs
Proof of Lemma 3.1.
For $j=1,\dots,q$, we have
$\check{\gamma}_{j}-\zeta_{j}=\sqrt{\frac{q}{T}}\sum_{t=(j-1)[T/q]+1}^{j[T/q]}\operatorname{sign}(x_{t-1})u_{t}=\sqrt{q}\int_{(j-1)/q}^{j/q}\sigma^{T}(r)\,dW^{T}(r).$
The stated result follows immediately from Assumptions 2.2 and 2.3. ∎
Proof of Proposition 3.2.
Part (a) follows from Theorem 1 and the discussion in Section 2.2 of Ibragimov and Müller (2010). For part (b), we deduce from Lemma 3.1 that
$\frac{1}{\sqrt{T/q}}\check{\gamma}_{j}=\frac{1}{\sqrt{T/q}}\zeta_{j}+o_{p}\!\left(\frac{1}{\sqrt{T}}\right)\to_{p}\beta\,E|x_{t}|$ (A.2)
uniformly in $j$, under $\beta\neq 0$. Recall that
$t_{q}(\check{\gamma})=\frac{\sqrt{q}\,\overline{\gamma}}{s_{\gamma}},\quad\text{with }\overline{\gamma}=\frac{1}{q}\sum_{j=1}^{q}\check{\gamma}_{j},\quad s_{\gamma}^{2}=\frac{1}{q-1}\sum_{j=1}^{q}(\check{\gamma}_{j}-\overline{\gamma})^{2}.$
Hence, the numerator of $t_{q}(\check{\gamma})$ satisfies
$\frac{\sqrt{q}\,\overline{\gamma}}{\sqrt{T}}\to_{p}\beta\,E|x_{t}|.$ (A.3)
To complete the proof, it suffices to show that $s_{\gamma}^{2}=o_{p}(T)$. Indeed, for $q\geq 2$, we have
$\frac{q(q-1)}{T}s_{\gamma}^{2}=\frac{1}{T/q}\sum_{j=1}^{q}(\check{\gamma}_{j}-\overline{\gamma})^{2}\to_{p}0$
due to (A.2) and (A.3), which completes the proof. ∎
Proof of Corollary 3.3.
We aim to show that
$\frac{\sqrt{q}}{T}\overline{\gamma}\to_{d}\beta\int_{0}^{1}|X(r)|\,dr,$ (A.4)
and
$\frac{q(q-1)}{T^{2}}s_{\gamma}^{2}=\frac{q}{T^{2}}\sum_{j=1}^{q}(\check{\gamma}_{j}-\overline{\gamma})^{2}\to_{d}\beta^{2}\sum_{j=1}^{q}\left(q\int_{(j-1)/q}^{j/q}|X(r)|\,dr-\int_{0}^{1}|X(r)|\,dr\right)^{2}.$ (A.5)
For (A.4), we have
$\frac{\sqrt{q}}{T}\check{\gamma}_{j}$ $=\frac{\sqrt{q}}{T}\zeta_{j}+o_{p}\!\left(\frac{1}{T}\right)=\beta\frac{q}{T^{3/2}}\sum_{t=(j-1)[T/q]+1}^{j[T/q]}|x_{t-1}|+o_{p}(1)\to_{d}\beta q\int_{(j-1)/q}^{j/q}|X(r)|\,dr,$
by Lemma 3.1 and (5), leading to (A.4). Moreover,
$\frac{q}{T^{2}}(\check{\gamma}_{j}-\overline{\gamma})^{2}\to_{d}\beta^{2}\left(q\int_{(j-1)/q}^{j/q}|X(r)|\,dr-\int_{0}^{1}|X(r)|\,dr\right)^{2},$
which yields (A.5).
Combining (A.4) and (A.5), we obtain
$t_{q}(\check{\gamma})=\frac{\sqrt{q}\,\overline{\gamma}}{s_{\gamma}}\to_{d}\operatorname{sign}(\beta)\mathcal{D}_{q},$
where
$\mathcal{D}_{q}=\int_{0}^{1}|X(r)|\,dr\left(\frac{q(q-1)}{\sum_{j=1}^{q}(\int_{0}^{1}|X(r)|\,dr-q\int_{(j-1)/q}^{j/q}|X(r)|\,dr)^{2}}\right)^{1/2},$
as desired. Furthermore, $(q-1)^{1/2}\mathcal{D}_{q}>1$ for $q\geq 2$ with probability one, since
$(q-1)^{1/2}\mathcal{D}_{q}$ $=(q-1)\int_{0}^{1}|X(r)|\,dr\left(\frac{q}{\sum_{j=1}^{q}(\int_{0}^{1}|X(r)|\,dr-q\int_{(j-1)/q}^{j/q}|X(r)|\,dr)^{2}}\right)^{1/2}$
$>\frac{(q-1)\int_{0}^{1}|X(r)|\,dr}{\max_{1\leq j\leq q}\left|\int_{0}^{1}|X(r)|\,dr-q\int_{(j-1)/q}^{j/q}|X(r)|\,dr\right|}>1,$
which completes the proof. ∎
Proof of Proposition 3.4.
It suffices to show that $\hat{\omega}^{2}\to_{p}\omega^{2}$, since this implies
$\tau(\check{\beta})=\frac{\check{\gamma}}{\omega}(1+o_{p}(1))=\left(\beta\frac{\sum_{t=1}^{T}|x_{t-1}|}{\omega\sqrt{T}}+\frac{\sum_{t=1}^{T}\operatorname{sign}(x_{t-1})u_{t}}{\omega\sqrt{T}}\right)(1+o_{p}(1)),$
where, in particular,
$\frac{\sum_{t=1}^{T}\operatorname{sign}(x_{t-1})u_{t}}{\omega\sqrt{T}}\to_{d}\mathbb{N}(0,1)$
under Assumptions 2.1, 2.2, and 2.3.
Let $\omega_{T}^{2}=T^{-1}\sum_{t=1}^{T}u_{t}^{2}$. Then $\omega_{T}^{2}\to_{p}\omega^{2}$ by Assumptions 2.1 and $2.$ Furthermore,
$\hat{\omega}^{2}=\omega_{T}^{2}-\frac{1}{T}\frac{(\sum_{t=1}^{T}x_{t-1}u_{t})^{2}}{\sum_{t=1}^{T}x_{t-1}^{2}}=\omega_{T}^{2}+o_{p}(1)$
by Assumption 3.1, which gives the desired result. ∎
Proof of Lemma 4.1.
We need only show that $H_{0}$ holds if $H_{0}^{(k)}$ holds for all $k$. Let $C=[\beta_{1},\ldots,\beta_{K}]^{\prime}$. By the moment restrictions, $C$ is the solution to
$E[z_{t-1}y_{t}]=E\!\left[\operatorname{diag}(|x_{1,t-1}|,\ldots,|x_{K,t-1}|)\right]C.$
Given $0<E[|x_{k,t-1}|]<\infty$ for all $k$, $H_{0}^{(k)}$ holds for all $k$ if and only if $E[z_{t-1}y_{t}]=\mathbf{0}_{K\times 1}$.
Moreover,
$E[z_{t-1}y_{t}]=E[z_{t-1}x_{t-1}^{\prime}]B,$
and since $E[z_{t-1}x_{t-1}^{\prime}]$ is assumed invertible, the condition $E[z_{t-1}y_{t}]=\mathbf{0}_{K\times 1}$ implies $B=0$, completing the proof. ∎
Proof of Corollary 4.2.
Under the modified version of Assumption 3.1 stated in the corollary, it follows from arguments analogous to those in the proof of Proposition 3.4 that $\hat{\omega}^{2}\to_{p}\omega^{2}$. Hence, it suffices to show that
$\frac{1}{\omega\sqrt{T}}\sum_{t=1}^{T/2}\operatorname{sign}(x_{2t-2})(u_{2t}-u_{2t-1})\to_{d}\mathbb{N}(0,1).$
Define $\xi_{t}=\operatorname{sign}(x_{t-2})u_{t}$ for even $t$ and $\xi_{t}=-\operatorname{sign}(x_{t-1})u_{t}$ for odd $t$, so that
$\frac{1}{\omega\sqrt{T}}\sum_{t=1}^{T/2}\operatorname{sign}(x_{2t-2})(u_{2t}-u_{2t-1})=\frac{1}{\omega\sqrt{T}}\sum_{t=1}^{T}\xi_{t}.$
By construction, $(\xi_{t})$ is an MDS with respect to $(\mathcal{F}_{t})$ and satisfies $E(\xi_{t}^{2}\mid\mathcal{F}_{t-1})=v_{t}^{2}$ under Assumption 2.1. The desired convergence then follows directly from the martingale central limit theorem, given Assumptions 2.2 and 2.3. ∎
Table 1: Size for Continuous-Time Models
$\bar{\kappa}=0$ $\bar{\kappa}=5$ $\bar{\kappa}=20$
T 20 50 100 20 50 100 20 50 100
CNST Cauchy RT 4.9 5.3 5.0 5.4 4.7 5.0 5.1 5.1 5.1
Cauchy VC 5.0 5.3 4.9 5.0 5.1 5.0 5.0 4.8 5.0
$t_{8}(\check{\gamma})$ 4.6 5.2 4.9 5.0 5.1 5.4 4.9 5.0 5.2
$t_{12}(\check{\gamma})$ 4.7 5.0 5.1 5.1 4.9 5.4 5.2 4.9 5.0
$t_{16}(\check{\gamma})$ 4.8 5.1 5.0 5.0 4.9 5.3 4.9 4.9 5.1
$\tau(\check{\beta})$ 4.5 5.2 5.0 4.8 5.1 5.2 4.9 4.8 5.2
SB Cauchy RT 5.0 5.1 5.8 5.3 5.0 5.9 5.0 4.9 5.9
Cauchy VC 6.7 6.3 4.7 6.5 6.0 4.9 6.4 6.0 5.0
$t_{8}(\check{\gamma})$ 3.7 3.9 3.6 4.1 3.7 4.1 3.7 3.7 3.9
$t_{12}(\check{\gamma})$ 4.2 4.6 4.2 4.6 4.5 4.3 4.2 4.5 4.1
$t_{16}(\check{\gamma})$ 4.6 4.6 4.6 4.7 4.5 4.7 4.6 4.5 4.8
$\tau(\check{\beta})$ 5.0 4.9 4.9 5.5 5.1 5.0 5.5 5.0 4.8
RS Cauchy RT 4.8 5.2 6.6 4.9 4.9 6.5 5.1 4.8 6.2
Cauchy VC 5.4 6.1 5.1 5.1 5.8 5.2 5.8 5.8 4.7
$t_{8}(\check{\gamma})$ 4.5 5.1 4.6 4.4 4.9 5.1 5.3 4.6 4.5
$t_{12}(\check{\gamma})$ 4.6 5.0 4.7 4.8 4.8 5.3 4.9 4.7 5.1
$t_{16}(\check{\gamma})$ 4.4 4.9 4.7 4.6 4.9 5.2 5.1 4.5 5.0
$\tau(\check{\beta})$ 4.9 5.2 4.9 4.7 4.9 5.1 5.3 5.0 5.3
GBM Cauchy RT 4.7 4.4 6.6 4.5 4.4 6.5 4.5 4.5 6.4
Cauchy VC 5.5 6.1 4.4 5.7 5.9 4.8 5.9 6.5 5.0
$t_{8}(\check{\gamma})$ 3.1 3.2 3.2 3.8 3.4 3.5 3.6 3.9 3.9
$t_{12}(\check{\gamma})$ 3.3 3.7 3.8 4.3 3.6 4.0 4.2 4.3 4.4
$t_{16}(\check{\gamma})$ 3.8 4.1 4.1 4.2 4.0 4.2 4.5 4.6 4.3
$\tau(\check{\beta})$ 4.6 5.0 4.5 4.8 4.9 4.8 4.9 5.2 5.1
This table reports the rejection rates of the Cauchy RT, Cauchy VC, $t_{q}(\check{\gamma})$ for $q\in\{8,12,16\}$, and $\tau(\check{\beta})$ tests based on simulated data under $\beta=0$ (two-sided tests). The data are generated from the continuous-time model (11) with persistence parameter $\bar{\kappa}\in\{0,5,20\}$. The model is sampled at $\Delta=1/252$ (daily), so that a sample of length $T$ years contains $252T$ observations, with $T\in\{20,50,100\}$. CNST, SB, GBM, and RS denote constant volatility, structural breaks, geometric Brownian motion, and regime switching, respectively.
Table 2: Power for Continuous-Time Models with Constant Volatility
$\bar{\kappa}=0$ $\bar{\kappa}=5$ $\bar{\kappa}=20$
T 20 50 100 20 50 100 20 50 100
$\beta=0.004$ Cauchy RT 8.8 25.2 85.4 6.1 9.9 22.7 6.3 13.2 27.9
Cauchy VC 9.1 25.6 86.2 6.5 10.0 23.5 8.2 13.4 26.8
$t_{8}(\check{\gamma})$ 8.0 21.0 79.5 6.7 11.2 24.6 5.1 14.6 30.6
$t_{12}(\check{\gamma})$ 9.5 21.4 82.0 6.2 10.5 25.4 5.9 14.3 31.0
$t_{16}(\check{\gamma})$ 8.8 22.8 84.0 5.5 10.0 23.3 5.5 14.5 29.9
$\tau(\check{\beta})$ 6.4 23.5 66.6 5.3 8.6 17.5 5.9 12.9 24.3
$\beta=0.008$ Cauchy RT 15.5 85.5 100.0 9.2 22.1 83.9 6.6 27.4 85.6
Cauchy VC 17.2 86.2 100.0 8.8 23.1 84.3 11.9 27.7 84.2
$t_{8}(\check{\gamma})$ 13.8 79.9 100.0 9.1 23.5 86.7 6.0 32.5 89.2
$t_{12}(\check{\gamma})$ 15.0 83.4 100.0 8.9 25.5 86.8 6.5 31.7 92.0
$t_{16}(\check{\gamma})$ 15.0 83.7 100.0 8.3 23.7 87.0 6.1 31.9 91.7
$\tau(\check{\beta})$ 14.0 66.7 96.6 7.3 18.8 57.6 8.5 27.6 70.0
$\beta=0.012$ Cauchy RT 37.3 99.2 100.0 12.2 50.1 100.0 7.7 52.2 100.0
Cauchy VC 40.2 99.3 100.0 12.6 51.9 100.0 16.5 56.3 100.0
$t_{8}(\check{\gamma})$ 30.3 98.0 100.0 12.9 55.5 99.8 6.9 60.6 100.0
$t_{12}(\check{\gamma})$ 33.1 99.5 100.0 12.0 56.3 100.0 7.2 61.6 100.0
$t_{16}(\check{\gamma})$ 35.3 98.9 100.0 12.3 54.1 100.0 6.6 59.2 100.0
$\tau(\check{\beta})$ 32.4 87.7 99.1 9.7 34.0 93.5 12.1 48.9 99.0
$\beta=0.016$ Cauchy RT 67.0 100.0 100.0 17.0 84.4 100.0 8.4 83.2 100.0
Cauchy VC 68.8 100.0 100.0 16.6 84.7 100.0 22.4 83.3 100.0
$t_{8}(\check{\gamma})$ 58.8 99.7 100.0 18.1 88.7 100.0 7.3 88.7 100.0
$t_{12}(\check{\gamma})$ 60.7 100.0 100.0 16.8 86.6 100.0 7.6 89.7 100.0
$t_{16}(\check{\gamma})$ 62.4 99.9 100.0 16.7 86.9 100.0 7.0 89.2 100.0
$\tau(\check{\beta})$ 51.2 95.8 99.9 12.4 58.5 99.0 17.8 72.9 100.0
$\beta=0.02$ Cauchy RT 86.7 100.0 100.0 23.2 98.0 100.0 9.3 98.2 100.0
Cauchy VC 87.9 100.0 100.0 24.4 98.5 100.0 29.4 97.4 100.0
$t_{8}(\check{\gamma})$ 79.3 100.0 100.0 24.6 97.8 100.0 8.5 98.3 100.0
$t_{12}(\check{\gamma})$ 81.9 100.0 100.0 23.8 99.1 100.0 8.5 99.1 100.0
$t_{16}(\check{\gamma})$ 82.3 100.0 100.0 23.2 98.3 100.0 7.7 99.2 100.0
$\tau(\check{\beta})$ 65.2 97.5 100.0 16.3 81.3 99.6 23.8 92.3 100.0
This table reports the rejection rates of the Cauchy RT, Cauchy VC, $t_{q}(\check{\gamma})$ for $q\in\{8,12,16\}$, and $\tau(\check{\beta})$ tests based on simulated data under $\beta\in\{0.004,0.008,0.012,0,016,0.02\}$ (two-sided tests). The data are generated from the continuous-time model (11) with persistence parameter $\bar{\kappa}\in\{0,5,20\}$. The model is sampled at $\Delta=1/252$ (daily), so that a sample of length $T$ years contains $252T$ observations, with $T\in\{20,50,100\}$. The volatility is constant.
Table 3: Power for Continuous-Time Models with Structural Breaks in Volatility
$\bar{\kappa}=0$ $\bar{\kappa}=5$ $\bar{\kappa}=20$
T 20 50 100 20 50 100 20 50 100
$\beta=0.004$ Cauchy RT 8.7 20.0 67.7 8.3 11.6 35.9 8.7 17.2 52.4
Cauchy VC 8.5 20.6 57.6 8.9 10.5 19.2 8.2 14.6 25.3
$t_{8}(\check{\gamma})$ 5.6 11.3 45.6 4.4 8.7 15.7 6.4 8.4 23.3
$t_{12}(\check{\gamma})$ 6.6 14.8 45.3 6.7 9.1 17.5 6.2 10.5 22.7
$t_{16}(\check{\gamma})$ 8.2 14.6 44.1 6.9 8.9 19.5 6.6 10.1 22.9
$\tau(\check{\beta})$ 5.8 13.7 31.4 7.1 8.1 13.5 5.8 10.2 16.7
$\beta=0.008$ Cauchy RT 14.4 65.1 96.6 11.7 35.7 88.4 14.5 52.2 99.6
Cauchy VC 14.9 60.6 95.6 10.8 21.6 55.8 11.5 30.1 72.9
$t_{8}(\check{\gamma})$ 9.5 44.1 88.5 6.8 16.3 57.4 9.0 21.0 68.3
$t_{12}(\check{\gamma})$ 11.4 45.6 89.8 8.9 17.9 55.9 9.2 23.6 69.4
$t_{16}(\check{\gamma})$ 12.7 42.3 92.1 8.4 18.3 55.9 8.8 22.2 72.2
$\tau(\check{\beta})$ 8.9 32.4 81.9 8.9 14.0 34.3 7.9 16.6 49.7
$\beta=0.012$ Cauchy RT 32.1 87.0 99.3 15.5 68.6 98.8 24.2 90.1 100.0
Cauchy VC 26.1 83.1 99.2 13.5 36.0 91.1 15.5 50.3 98.8
$t_{8}(\check{\gamma})$ 16.9 72.9 97.0 9.2 32.7 89.0 12.0 43.1 95.2
$t_{12}(\check{\gamma})$ 20.7 74.5 99.1 11.9 34.5 90.0 12.2 41.7 98.0
$t_{16}(\check{\gamma})$ 20.0 74.3 99.6 10.0 31.9 90.7 12.6 41.4 98.8
$\tau(\check{\beta})$ 14.7 60.2 98.1 10.8 23.4 73.7 10.8 33.2 92.1
$\beta=0.016$ Cauchy RT 50.2 95.6 99.7 22.8 89.8 100.0 37.7 99.7 100.0
Cauchy VC 42.9 95.1 99.6 17.0 58.9 99.5 20.7 74.5 100.0
$t_{8}(\check{\gamma})$ 30.2 88.8 98.5 12.6 56.2 96.1 16.6 67.4 99.8
$t_{12}(\check{\gamma})$ 31.7 90.4 99.7 14.7 53.8 98.4 17.0 65.0 100.0
$t_{16}(\check{\gamma})$ 31.8 91.1 100.0 13.0 54.4 99.0 18.5 66.6 100.0
$\tau(\check{\beta})$ 23.2 81.4 99.8 13.0 35.6 96.9 14.3 51.8 100.0
$\beta=0.02$ Cauchy RT 65.4 98.6 99.9 34.6 97.4 100.0 55.1 100.0 100.0
Cauchy VC 58.8 98.5 100.0 21.4 78.7 99.9 27.1 91.5 100.0
$t_{8}(\check{\gamma})$ 45.5 95.3 98.8 17.6 74.0 98.2 21.7 86.2 100.0
$t_{12}(\check{\gamma})$ 46.1 96.4 99.8 18.0 75.7 99.7 22.7 87.5 100.0
$t_{16}(\check{\gamma})$ 44.7 97.9 100.0 18.0 73.9 100.0 24.9 87.9 100.0
$\tau(\check{\beta})$ 32.2 92.4 100.0 15.7 56.6 99.8 18.2 73.4 100.0
This table reports the rejection rates of the Cauchy RT, Cauchy VC, $t_{q}(\check{\gamma})$ for $q\in\{8,12,16\}$, and $\tau(\check{\beta})$ tests based on simulated data under $\beta\in\{0.004,0.008,0.012,0,016,0.02\}$ (two-sided tests). The data are generated from the continuous-time model (11) with persistence parameter $\bar{\kappa}\in\{0,5,20\}$. The model is sampled at $\Delta=1/252$ (daily), so that a sample of length $T$ years contains $252T$ observations, with $T\in\{20,50,100\}$. The volatility exhibits structural breaks.
Table 4: Power for Continuous-Time Models with Regime Switching in Volatility
$\bar{\kappa}=0$ $\bar{\kappa}=5$ $\bar{\kappa}=20$
T 20 50 100 20 50 100 20 50 100
$\beta=0.004$ Cauchy RT 6.6 23.8 64.2 5.9 10.3 21.0 8.1 12.8 32.9
Cauchy VC 11.0 49.9 85.9 7.8 12.9 33.0 7.4 13.0 34.5
$t_{8}(\check{\gamma})$ 7.7 22.1 77.7 6.1 11.8 22.2 7.5 10.4 28.4
$t_{12}(\check{\gamma})$ 8.0 23.2 83.6 6.2 10.9 22.6 7.7 11.0 28.9
$t_{16}(\check{\gamma})$ 7.9 22.8 84.0 6.5 11.8 22.9 7.8 10.6 30.8
$\tau(\check{\beta})$ 8.2 22.3 64.8 7.2 7.6 15.0 8.3 9.3 20.2
$\beta=0.008$ Cauchy RT 11.9 69.3 96.0 8.0 20.3 65.5 12.2 32.2 90.2
Cauchy VC 25.7 84.6 97.8 10.0 30.2 84.1 10.8 28.2 86.6
$t_{8}(\check{\gamma})$ 15.9 77.2 99.8 9.6 24.3 79.8 10.5 26.9 81.9
$t_{12}(\check{\gamma})$ 15.2 80.5 99.9 9.2 23.8 81.3 11.0 27.0 82.9
$t_{16}(\check{\gamma})$ 16.0 82.1 100.0 9.5 24.9 81.4 11.5 26.6 83.7
$\tau(\check{\beta})$ 16.7 67.7 94.4 9.1 13.8 50.9 10.3 19.2 60.0
$\beta=0.012$ Cauchy RT 31.6 87.0 98.9 10.7 39.8 95.5 17.9 63.4 99.5
Cauchy VC 47.3 94.5 98.5 12.1 58.3 95.4 14.9 53.5 98.7
$t_{8}(\check{\gamma})$ 33.0 96.7 100.0 13.9 50.8 99.1 15.4 51.7 99.4
$t_{12}(\check{\gamma})$ 35.1 98.2 100.0 13.3 52.2 99.5 15.1 55.5 99.8
$t_{16}(\check{\gamma})$ 34.8 98.3 100.0 13.2 51.2 99.1 15.4 53.5 99.8
$\tau(\check{\beta})$ 32.3 86.6 98.2 11.1 29.0 87.6 13.1 37.8 95.0
$\beta=0.016$ Cauchy RT 52.7 95.2 99.7 14.2 65.3 99.3 26.2 90.6 100.0
Cauchy VC 65.1 97.2 99.0 17.3 81.7 97.4 19.3 78.2 99.6
$t_{8}(\check{\gamma})$ 52.5 99.7 100.0 18.0 82.1 99.9 21.8 80.0 100.0
$t_{12}(\check{\gamma})$ 59.7 99.6 100.0 18.2 83.6 100.0 20.9 82.5 100.0
$t_{16}(\check{\gamma})$ 61.3 99.9 100.0 19.4 81.8 100.0 21.6 81.0 100.0
$\tau(\check{\beta})$ 51.6 94.3 98.9 13.5 54.9 96.6 17.4 59.5 99.5
$\beta=0.02$ Cauchy RT 66.4 97.7 100.0 20.9 86.1 99.9 36.2 98.2 100.0
Cauchy VC 76.8 98.0 99.2 21.5 91.1 98.7 25.0 92.1 99.7
$t_{8}(\check{\gamma})$ 74.3 100.0 100.0 23.7 95.4 99.9 27.8 94.7 100.0
$t_{12}(\check{\gamma})$ 80.4 99.9 100.0 23.6 96.3 100.0 28.4 95.8 100.0
$t_{16}(\check{\gamma})$ 81.1 100.0 100.0 25.7 96.8 100.0 29.1 96.3 100.0
$\tau(\check{\beta})$ 67.3 97.0 99.1 17.3 77.2 98.7 22.7 81.8 99.8
This table reports the rejection rates of the Cauchy RT, Cauchy VC, $t_{q}(\check{\gamma})$ for $q\in\{8,12,16\}$, and $\tau(\check{\beta})$ tests based on simulated data under $\beta\in\{0.004,0.008,0.012,0,016,0.02\}$ (two-sided tests). The data are generated from the continuous-time model (11) with persistence parameter $\bar{\kappa}\in\{0,5,20\}$. The model is sampled at $\Delta=1/252$ (daily), so that a sample of length $T$ years contains $252T$ observations, with $T\in\{20,50,100\}$. The volatility exhibits regime switching.
Table 5: Power for Continuous-Time Models with a Geometric Brownian Motion
$\bar{\kappa}=0$ $\bar{\kappa}=5$ $\bar{\kappa}=20$
T 20 50 100 20 50 100 20 50 100
$\beta=0.004$ Cauchy RT 7.8 19.2 59.9 6.0 9.6 24.6 8.5 14.0 41.8
Cauchy VC 42.4 59.7 80.4 6.7 13.0 29.5 7.2 10.8 19.6
$t_{8}(\check{\gamma})$ 11.0 27.1 67.7 4.4 7.5 21.4 5.7 8.4 18.8
$t_{12}(\check{\gamma})$ 10.3 28.5 72.0 5.4 8.8 21.7 4.9 8.1 19.2
$t_{16}(\check{\gamma})$ 10.2 29.2 73.2 5.5 9.3 21.2 5.8 9.2 19.9
$\tau(\check{\beta})$ 8.8 25.5 61.8 5.6 7.5 13.3 6.2 7.0 14.7
$\beta=0.008$ Cauchy RT 13.7 59.1 92.4 8.2 24.7 73.4 13.2 40.9 92.8
Cauchy VC 55.3 78.9 91.2 9.4 28.0 53.6 8.8 20.5 51.2
$t_{8}(\check{\gamma})$ 21.6 65.5 94.6 7.0 18.8 68.3 8.4 16.3 57.5
$t_{12}(\check{\gamma})$ 23.9 69.7 95.5 7.4 21.2 70.9 7.0 18.6 63.2
$t_{16}(\check{\gamma})$ 22.4 71.3 96.5 7.9 21.1 72.2 8.2 19.3 64.7
$\tau(\check{\beta})$ 21.2 59.2 88.0 6.7 12.0 35.0 8.0 14.5 40.5
$\beta=0.012$ Cauchy RT 29.0 80.8 97.9 11.2 48.3 95.0 18.7 72.2 99.2
Cauchy VC 65.6 88.3 92.2 13.8 42.9 65.6 12.1 36.1 69.9
$t_{8}(\check{\gamma})$ 38.9 85.5 97.3 10.7 43.0 89.6 10.4 35.2 81.2
$t_{12}(\check{\gamma})$ 40.2 88.4 98.5 11.0 43.6 93.0 10.5 39.8 88.0
$t_{16}(\check{\gamma})$ 42.0 89.0 99.1 10.0 44.4 92.8 11.1 39.6 91.5
$\tau(\check{\beta})$ 38.9 77.0 92.1 8.3 22.3 64.9 10.8 26.3 69.7
$\beta=0.016$ Cauchy RT 47.9 91.7 99.1 16.0 70.5 98.5 27.8 91.5 99.7
Cauchy VC 73.4 91.6 93.4 17.6 56.2 70.4 15.0 52.5 77.3
$t_{8}(\check{\gamma})$ 53.4 91.8 98.4 14.6 68.5 94.4 13.9 57.8 83.9
$t_{12}(\check{\gamma})$ 57.4 94.9 99.3 14.9 71.3 97.1 15.1 63.9 93.9
$t_{16}(\check{\gamma})$ 58.9 95.2 99.6 14.4 72.6 98.2 15.9 62.9 96.6
$\tau(\check{\beta})$ 51.2 86.3 93.5 10.3 37.4 74.8 13.1 42.1 82.3
$\beta=0.02$ Cauchy RT 60.1 96.6 99.5 23.7 87.6 98.9 40.3 97.9 99.9
Cauchy VC 79.7 92.6 93.9 22.3 64.9 72.9 19.4 63.6 80.7
$t_{8}(\check{\gamma})$ 67.5 94.8 99.0 20.2 81.4 96.0 18.3 71.0 85.5
$t_{12}(\check{\gamma})$ 69.9 97.2 99.7 19.7 85.7 98.6 20.0 80.7 95.0
$t_{16}(\check{\gamma})$ 72.9 97.2 99.9 20.2 87.5 99.1 21.2 79.9 98.2
$\tau(\check{\beta})$ 59.9 90.7 94.4 12.8 54.6 80.2 16.8 57.5 89.5
This table reports the rejection rates of the Cauchy RT, Cauchy VC, $t_{q}(\check{\gamma})$ for $q\in\{8,12,16\}$, and $\tau(\check{\beta})$ tests based on simulated data under $\beta\in\{0.004,0.008,0.012,0,016,0.02\}$ (two-sided tests). The data are generated from the continuous-time model (11) with persistence parameter $\bar{\kappa}\in\{0,5,20\}$. The model is sampled at $\Delta=1/252$ (daily), so that a sample of length $T$ years contains $252T$ observations, with $T\in\{20,50,100\}$. The volatility follows a geometric Brownian motion.
Table 6: Size and Power for Discrete-Time Models with Constant Volatility and MA(1) Innovations
$\bar{\kappa}=0$ $\bar{\kappa}=50$ $\bar{\kappa}=100$
T 20 50 100 20 50 100 20 50 100
$\beta=0$ $\tau(\check{\beta}_{o})$ 5.0 4.9 4.8 5.2 5.3 4.8 4.5 4.7 5.2
$\tau(\check{\beta}_{e})$ 4.8 4.7 5.2 5.0 4.8 4.8 5.1 5.0 4.9
IVX 14.2 13.5 12.3 10.2 10.5 10.3 9.5 10.3 10.5
$\beta=0.5$ $\tau(\check{\beta}_{o})$ 8.4 7.9 8.4 17.0 17.5 16.6 24.9 24.9 25.8
$\tau(\check{\beta}_{e})$ 8.2 8.0 8.4 17.4 16.0 17.1 25.4 25.7 26.0
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1$ $\tau(\check{\beta}_{o})$ 13.4 13.4 13.9 39.3 41.1 39.7 60.5 61.1 62.0
$\tau(\check{\beta}_{e})$ 13.7 13.5 13.7 39.6 39.1 39.6 62.0 62.8 62.5
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1.5$ $\tau(\check{\beta}_{o})$ 20.2 20.1 20.8 63.3 65.2 65.0 85.9 87.6 87.7
$\tau(\check{\beta}_{e})$ 20.7 20.4 20.6 63.3 64.0 64.3 86.5 87.8 87.8
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2$ $\tau(\check{\beta}_{o})$ 27.7 27.3 28.3 80.3 82.1 82.9 95.4 96.6 97.2
$\tau(\check{\beta}_{e})$ 28.0 27.3 28.4 80.1 81.4 82.6 95.9 97.0 97.1
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2.5$ $\tau(\check{\beta}_{o})$ 34.3 33.7 34.7 89.5 91.7 92.1 98.6 99.2 99.3
$\tau(\check{\beta}_{e})$ 34.8 33.8 35.4 89.1 91.2 92.0 98.7 99.1 99.3
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
This table reports the rejection rates of $\tau(\check{\beta}_{o})$ and $\tau(\check{\beta}_{e})$, as well as the IVX tests, based on simulated data under $\beta\in\{0,0.5,1,1.5,2,2.5\}$ (one-sided tests). The data are generated from the discrete-time model (S.2)–(S.3) with persistence parameter $\bar{\kappa}\in\{0,50,100\}$ and MA(1) innovations. The model is sampled at a monthly frequency, so that a sample of length $T$ years contains $12T$ observations, with $T\in\{20,50,100\}$. The volatility is constant.
Table 7: Size and Power for Discrete-Time Models with Structural Breaks in Volatility and MA(1) Innovations
$\bar{\kappa}=0$ $\bar{\kappa}=50$ $\bar{\kappa}=100$
T 20 50 100 20 50 100 20 50 100
$\beta=0$ $\tau(\check{\beta}_{o})$ 4.5 4.9 5.1 5.3 5.2 4.8 5.1 4.7 4.7
$\tau(\check{\beta}_{e})$ 4.8 4.8 4.9 5.0 4.5 4.8 5.0 4.9 4.9
IVX 30.7 31.2 30.1 32.8 34.2 34.3 33.9 34.5 36.0
$\beta=0.5$ $\tau(\check{\beta}_{o})$ 7.7 8.1 8.7 13.7 13.7 13.2 18.5 18.7 19.3
$\tau(\check{\beta}_{e})$ 8.0 8.2 8.4 13.6 13.0 12.8 19.3 19.7 19.7
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1$ $\tau(\check{\beta}_{o})$ 13.5 13.5 14.3 29.2 29.3 28.5 44.3 45.0 46.0
$\tau(\check{\beta}_{e})$ 14.0 13.8 13.9 28.2 28.1 27.3 46.1 46.6 46.1
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1.5$ $\tau(\check{\beta}_{o})$ 20.4 20.4 21.5 46.8 48.1 47.6 68.9 70.8 72.2
$\tau(\check{\beta}_{e})$ 21.3 20.8 21.0 46.9 46.3 47.5 71.2 71.5 72.5
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2$ $\tau(\check{\beta}_{o})$ 27.6 27.7 28.9 62.7 64.9 64.7 84.4 87.3 88.0
$\tau(\check{\beta}_{e})$ 28.3 28.0 28.0 63.2 63.2 64.5 85.7 87.3 87.9
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2.5$ $\tau(\check{\beta}_{o})$ 34.4 34.3 35.5 74.0 76.9 77.2 92.1 94.6 95.4
$\tau(\check{\beta}_{e})$ 35.1 34.6 34.3 74.7 75.9 77.3 93.0 94.6 95.3
q=4 23.8 23.3 23.4 34.1 37.0 38.4 46.8 50.2 51.4
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
This table reports the rejection rates of $\tau(\check{\beta}_{o})$ and $\tau(\check{\beta}_{e})$, as well as the IVX tests, based on simulated data under $\beta\in\{0,0.5,1,1.5,2,2.5\}$ (one-sided tests). The data are generated from the discrete-time model (S.2)–(S.3) with persistence parameter $\bar{\kappa}\in\{0,50,100\}$ and MA(1) innovations. The model is sampled at a monthly frequency, so that a sample of length $T$ years contains $12T$ observations, with $T\in\{20,50,100\}$. The volatility exhibits structural breaks.
Table 8: Size and Power for Discrete-Time Models with Regime Switching in Volatility and MA(1) Innovations
$\bar{\kappa}=0$ $\bar{\kappa}=50$ $\bar{\kappa}=100$
T 20 50 100 20 50 100 20 50 100
$\beta=0$ $\tau(\check{\beta}_{o})$ 3.2 4.0 4.3 4.7 5.0 5.0 5.2 4.9 4.8
$\tau(\check{\beta}_{e})$ 3.8 3.9 4.2 5.0 4.9 4.9 4.9 5.1 4.8
IVX 13.5 13.2 12.4 9.7 11.1 11.0 10.3 11.2 12.4
$\beta=0.5$ $\tau(\check{\beta}_{o})$ 5.8 6.8 7.4 16.1 17.1 16.6 23.4 24.1 24.4
$\tau(\check{\beta}_{e})$ 6.4 6.7 6.9 16.8 16.8 16.6 23.9 24.9 23.6
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1$ $\tau(\check{\beta}_{o})$ 10.4 11.3 12.0 39.6 40.9 39.3 56.7 60.4 60.6
$\tau(\check{\beta}_{e})$ 11.0 11.7 11.8 39.7 39.6 39.8 57.3 60.9 59.6
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1.5$ $\tau(\check{\beta}_{o})$ 17.3 18.2 19.0 64.3 67.1 63.9 81.5 86.0 87.0
$\tau(\check{\beta}_{e})$ 17.9 18.6 18.7 63.8 65.3 64.4 82.2 86.8 86.2
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2$ $\tau(\check{\beta}_{o})$ 25.0 25.4 26.5 80.2 83.1 81.7 92.0 95.5 96.4
$\tau(\check{\beta}_{e})$ 25.6 26.2 25.9 79.4 82.0 82.1 92.5 95.5 96.1
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2.5$ $\tau(\check{\beta}_{o})$ 32.6 32.5 33.9 88.9 91.4 90.7 96.5 98.6 99.1
$\tau(\check{\beta}_{e})$ 33.4 33.6 32.7 88.2 90.9 90.6 96.8 98.6 98.9
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
This table reports the rejection rates of $\tau(\check{\beta}_{o})$ and $\tau(\check{\beta}_{e})$, as well as the IVX tests, based on simulated data under $\beta\in\{0,0.5,1,1.5,2,2.5\}$ (one-sided tests). The data are generated from the discrete-time model (S.2)–(S.3) with persistence parameter $\bar{\kappa}\in\{0,50,100\}$ and MA(1) innovations. The model is sampled at a monthly frequency, so that a sample of length $T$ years contains $12T$ observations, with $T\in\{20,50,100\}$. The volatility exhibits regime switching.
Table 9: Empirical Results on Stock Return Predictability (Two-Sided Tests)
Series Frequency $\beta$ Cauchy RT Cauchy VC $t_{12}(\check{\gamma})$ $t_{16}(\check{\gamma})$ $\tau(\check{\beta})$
Panel A: D/P as predictor for the period of 1927-2011
CRSP Monthly 0.005 1.06 0.85 0.42 0.44 0.43
Quarterly 0.007 0.48 1.26 1.29 1.34 1.09
Yearly 0.063 0.99 $2.15^{**}$ $2.04^{**}$ $2.35^{**}$ $1.82^{**}$
S&P500 Monthly 0.003 0.74 1.14 0.89 0.82 0.72
Quarterly 0.008 0.65 $1.78^{**}$ $1.75^{*}$ $1.73^{*}$ $1.31^{*}$
Yearly 0.042 0.82 $2.24^{**}$ $2.33^{**}$ $2.00^{**}$ $1.95^{**}$
Panel B: D/P as predictor for the period of 1927-2011 with jumps removed
CRSP Monthly 0.001 0.22 $50^{*} 1.1$5 1.06 0.90
Quarterly 0.015 0.93 $2.02^{**}$ $2.45^{**}$ $2.40^{**}$ $1.46^{*}$
Yearly 0.005 0.05 0.56 0.56 0.53 0.45
S&P500 Monthly 0.002 0.48 $1.66^{*}$ 1.35 1.14 1.02
Quarterly 0.017 1.21 $1.49^{*}$ 1.70 1.61 1.20
Yearly 0.008 0.11 $1.51^{*}$ $2.45^{**}$ $2.56^{**}$ $1.80^{**}$
Panel C: E/P as predictor for the period of 1950-2011
S&P500 Monthly 0.000 -0.05 0.32 0.26 0.22 0.30
Quarterly 0.007 0.39 0.39 0.38 0.35 0.35
Yearly 0.059 0.83 0.76 1.04 0.23 0.84
Panel D: E/P as predictor for the period of 1950-2011 with jumps removed
S&P500 Monthly 0.001 0.16 0.01 0.14 0.18 0.11
Quarterly 0.000 0.02 0.71 0.78 0.61 0.48
Yearly -0.038 -0.38 0.69 -0.05 0.55 0.44
Results for two-sided tests of return predictability for the NYSE/AMEX value-weighted index (CRSP) and the S&P 500 using the Cauchy RT, Cauchy VC, $t_{q}$ ($q=12,16$), and $\tau(\check{\beta})$ tests across different regression frequencies. Panels A–B use the dividend–price ratio (D/P), while Panels C–D use the earnings–price ratio (E/P) as predictors. Statistical significance at the 5% and 1% levels is denoted by ‘‘’’ and ‘‘∗∗’’, respectively.
B Additional Theoretical Analysis
B.1 Wald Test for Joint Predictability Test
Consider a predictive regression with $K$ predictors $x_{t}=[x_{1,t},\ldots,x_{K,t}]^{\prime}$:
$y_{t}$ $=x_{t-1}^{\prime}B+u_{t}$
$=\beta_{1,K}x_{1,t-1}+\cdots+\beta_{K,K}x_{K,t-1}+u_{t},\qquad B=[\beta_{1,K},\ldots,\beta_{K,K}]^{\prime}.$ (S.1)
The objective is to test the joint predictability of the covariates, that is,
$H_{0}:\beta_{1,K}=\cdots=\beta_{K,K}=0.$
Define $z_{t}=[\operatorname{sign}(x_{1,t}),\ldots,\operatorname{sign}(x_{K,t})]^{\prime}$. Clearly, under $H_{0}$,
$\left(\hat{\omega}^{2}\sum_{t=1}^{T}z_{t-1}z_{t-1}^{\prime}\right)^{-1/2}\left(\sum_{t=1}^{T}z_{t-1}y_{t}\right)\to_{d}\mathbb{N}(0,I_{K})$
as long as $\hat{\omega}^{2}\to_{p}\omega^{2}=\int_{0}^{1}\sigma^{2}(r)\,dr$ and the limit of $\sum_{t=1}^{T}z_{t-1}z_{t-1}^{\prime}$ is invertible. Our test statistic is motivated by the convergence, and we propose a Wald-type statistic $W$ defined as
$W=\left(\sum_{t=1}^{T}z_{t-1}y_{t}\right)^{\prime}\left(\hat{\omega}^{2}\sum_{t=1}^{T}z_{t-1}z_{t-1}^{\prime}\right)^{-1}\left(\sum_{t=1}^{T}z_{t-1}y_{t}\right),$
where
$\hat{\omega}^{2}=\frac{1}{T}\sum_{t=1}^{T}\hat{u}_{t}^{2},\qquad\hat{u}_{t}=y_{t}-x_{t-1}^{\prime}\hat{B}$
with the OLS estimator $\hat{B}$. As in the main paper, $\hat{\omega}^{2}$ estimates $\omega^{2}=\int_{0}^{1}\sigma^{2}(r)\,dr$ using OLS residuals.
For our purpose, we assume
Assumption B.1.
  1. (i)
    $\frac{1}{T}\sum_{t=1}^{T}z_{t-1}z_{t-1}^{\prime}\to_{d}M_{z}$, where $M_{z}$ is invertible.
  2. (ii)
    $\left\|\sum_{t=1}^{T}x_{t-1}u_{t}\right\|_{2}=o_{p}\!\left(\sqrt{T\lambda_{T}}\right)$, where $\lambda_{T}$ is the minimum eigenvalue of $\sum_{t=1}^{T}x_{t-1}x_{t-1}^{\prime}$ and the norm $\|\cdot\|_{2}$ is the spectral norm.
  3. (iii)
    let $\varsigma_{T}$ be the minimum singular value of $\sum_{t=1}^{T}z_{t-1}x_{t-1}^{\prime}$ such that $\varsigma_{T}/\sqrt{T}\to_{p}\infty$.
Assumption B.1 is essentially a multivariate extension of the conditions used for Proposition 3.4 in the main paper as we will discuss subsequently. Firtsly, Assumption B.1(i) generalizes the univariate case, in which we always have $\sum_{t=1}^{T}z_{t-1}^{2}=T$ for all $T$ with probability one. Secondly, Assumption B.1(ii) extends Assumption 3.1 in the main paper. Lastly, Assumption B.1(iii) is a multivariate version of the test consistency assumption $\sum_{t=1}^{T}|x_{t-1}|/\sqrt{T}\to_{p}\infty$ used in Proposition 3.4(b) in the main paper.
Note that Assumption B.1 is violated if $x_{t}$ is supported only on the positive (or negative) real line. In this case, one may subtract $m_{t}$, where $m_{t}$ is an $\mathcal{F}_{t-1}$-adapted process taking values in the support of $x_{t}$, and apply the test to the transformed regressor $x_{t}-m_{t}$ (see Footnote 9 of the main paper). Moreover, when an intercept is included, the Wald test can be combined with the first-differencing approach described in Section 4.2 of the main paper.
The proposed Wald test is not straightforward to extend to general linear hypotheses, such as $R\beta=r$ for some $q\times K$ restriction matrix $R$ and $q\times 1$ vector $r$. This is because the asymptotic distribution of the restricted Cauchy-type IV estimator is complex and, in general, depends on the degrees of persistence and heavy-tailedness of the covariates $x_{t}$. Accordingly, our Wald statistic is limited to joint predictability tests for the full set of covariates in the predictive regression, whereas the IVX method of Kostakis et al. (2015) can be applied to test general linear restrictions under their maintained assumptions.
Nonetheless, testing joint significance for a subset of parameters may be feasible by combining our Wald test with a partialling-out approach based on the Frisch–Waugh–Lovell theorem, under additional regularity conditions. We leave such generalizations of our Wald-type approach for future research.
Proposition B.1.
Let Assumptions 2.1, 2.2, and 2.3 in the main paper hold.
(a) Let Assumption B.1(i)-(ii) hold. Then, under $H_{0}:B=\mathbf{0}_{K\times 1}$, $W\to_{d}\chi_{K}^{2}$.
(b) Let Assumption B.1(i)-(iii) hold. Then, under $H_{A}:B\neq\mathbf{0}_{K\times 1}$, $W\to_{p}\infty$.
Proof.
Let $\omega_{T}^{2}=T^{-1}\sum_{t=1}^{T}u_{t}^{2}$. Then $\omega_{T}^{2}\to_{p}\omega^{2}$ as in the proof of Proposition 3.4 of the main paper. Furthermore, we have
$\hat{\omega}^{2}=\omega_{T}^{2}-\left(\sum_{t=1}^{T}x_{t-1}u_{t}\right)^{\prime}\left(T\sum_{t=1}^{T}x_{t-1}x_{t-1}^{\prime}\right)^{-1}\left(\sum_{t=1}^{T}x_{t-1}u_{t}\right)=\omega_{T}^{2}+o_{p}(1)$
by Assumption B.1(ii). Part (a) then follows immediately from Assumption B.1(i).
As for Part (b), we note that under $H_{A}$
$W=_{d}\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}z_{t-1}x_{t-1}^{\prime}B\right)^{\prime}\left(\omega M_{z}\right)^{-1}\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}z_{t-1}x_{t-1}^{\prime}B\right)(1+o_{p}(1)).$
Therefore, $W\to_{p}\infty$ as long as
$\frac{1}{\sqrt{T}}\sum_{t=1}^{T}z_{t-1}x_{t-1}^{\prime}B\to_{p}\infty$
for any fixed $B\neq\mathbf{0}_{K\times 1}$ in $\mathbb{R}^{K}$, which is implied by Assumption B.1(iii). This completes the proof. ∎
B.2 Effects of First Differencing on Testing Power
Corollary 4.2(b) of the main paper establishes the consistency of the test based on the first-differenced model. However, first differencing generally entails a loss of power relative to the test that imposes a zero intercept. In this subsection, we provide a theoretical comparison of the asymptotic power of the two tests.
Specifically, consider the linear predictive regression without intercept,
$y_{t}=\beta x_{t-1}+u_{t},$
where $(x_{t})$ follows the local-to-unity process of Phillips and Magdalinos (2007b),
$x_{t}=\left(1+\frac{c}{T^{\delta}}\right)x_{t-1}+\varepsilon_{t}^{x},\quad c<0,\quad\delta\in[0,1],$
and $(\varepsilon_{t}^{x})$ satisfies Assumption LP therein. We assume that Assumptions 2.1–2.3 in the main paper hold.
We compare the rates of divergence of the test statistics $\tau(\check{\beta})$ and $\tau(\check{\beta}_{e})$, defined as
$\tau(\check{\beta})=\frac{\sum_{t=1}^{T}\text{sign}(x_{t-1})y_{t}}{\hat{\omega}\sqrt{T}}\quad\text{and}\quad\tau(\check{\beta}_{e})=\frac{\sum_{t=2}^{T/2}\text{sign}(x_{2t-2})(y_{2t}-y_{2t-1})}{\hat{\omega}_{d}\sqrt{T/2}},$
where $\hat{\omega}^{2}=T^{-1}\sum_{t=1}^{T}\hat{u}_{t}^{2}$ and $\hat{\omega}_{d}^{2}=T^{-1}\sum_{t=1}^{T}(\hat{u}_{t}^{(d)})^{2}$, with $(\hat{u}_{t})$ denoting the OLS residuals from the original model and $(\hat{u}_{t}^{(d)})$ those from the demeaned model.
It follows from Proposition 3.4(b) and Corollary 4.2(b) of the main paper that, under $\beta\neq 0$,
$\tau(\check{\beta})=\beta\frac{\sum_{t=1}^{T}|x_{t-1}|}{\omega\sqrt{T}}(1+o_{p}(1))+O_{p}(1),$
and
$\tau(\check{\beta}_{e})=\beta\frac{\sum_{t=1}^{T/2}\text{sign}(x_{2t-2})(x_{2t-1}-x_{2t-2})}{\omega\sqrt{T/2}}(1+o_{p}(1))+O_{p}(1),$
provided that Assumption 3.1 holds for both $x_{t-1}$ and $x_{t-1}-\bar{x}_{T}$.
The rate of divergence for $\tau(\check{\beta})$ follows directly from Lemma 3.2 of Phillips and Magdalinos (2007b):
$\frac{1}{T^{1/2+\delta/2}}\tau(\check{\beta})=\frac{\beta}{\omega}M_{1}(1+o_{p}(1)),$
where $M_{k}=E|X_{k}(1)|$ for $k=1,2$, and
$X_{k}(t)=\int_{0}^{t}\exp\bigl(kc(t-r)\bigr)dB_{k}(r),$
with $(B_{k})$ denoting Brownian motions defined in Assumption LP of Phillips and Magdalinos (2007b).
For $\tau(\check{\beta}_{e})$, we decompose
$\tau(\check{\beta}_{e})=\frac{\beta}{\omega/\sqrt{2}}\bigl[\tau_{M}(\check{\beta}_{e})+\tau_{R}(\check{\beta}_{e})\bigr](1+o_{p}(1))+O_{p}(1),$
where
$\tau_{M}(\check{\beta}_{e})=\frac{c}{T^{1/2+\delta}}\sum_{t=1}^{T/2}|x_{2t-2}|,\quad\tau_{R}(\check{\beta}_{e})=\frac{1}{\sqrt{T}}\sum_{t=1}^{T/2}\text{sign}(x_{2t-2})\varepsilon_{2t-1}^{x},$
using the decomposition
$\sum_{t=1}^{T/2}\text{sign}(x_{2t-2})(x_{2t-1}-x_{2t-2})$ $=\frac{c}{T^{\delta}}\sum_{t=1}^{T/2}|x_{2t-2}|+\sum_{t=1}^{T/2}\text{sign}(x_{2t-2})\varepsilon_{2t-1}^{x}.$
Using the second-order representation
$x_{t}=\left(1+\frac{2c}{T^{\delta}}+\frac{c^{2}}{T^{2\delta}}\right)x_{t-2}+\varepsilon_{t}^{x}+\left(1+\frac{c}{T^{\delta}}\right)\varepsilon_{t-1}^{x},$
Lemma 3.2 of Phillips and Magdalinos (2007b) yields
$\frac{1}{T^{1/2-\delta/2}}\tau_{M}(\check{\beta}_{e})=M_{2}(1+o_{p}(1)),$
so that $\tau_{M}(\check{\beta}_{e})$ diverges at rate $T^{1/2-\delta/2}$ whenever $\delta<1$. For the remainder term, $\tau_{R}(\check{\beta}_{e})$ diverges at rate $\sqrt{T}$ if
$E[\text{sign}(x_{2t-2})\varepsilon_{2t-1}^{x}]\neq 0,$
and otherwise $\tau_{R}(\check{\beta}_{e})=O_{p}(1)$.
We summarize the results as follows:
  • $\tau(\check{\beta})$ is always consistent and diverges at rate $T^{1/2+\delta/2}$.
  • $\tau(\check{\beta}_{e})$ is consistent if either $\delta<1$ or $E[\text{sign}(x_{2t-2})\varepsilon_{2t-1}^{x}]\neq 0$. Its divergence rate is $T^{1/2-\delta/2}$ when the expectation is zero, and $\sqrt{T}$ otherwise.
  • Consequently, the divergence rate of $\tau(\check{\beta})$ is uniformly faster than that of $\tau(\check{\beta}_{e})$, implying a loss of power induced by first differencing.
An important observation is that the test based on $\tau(\check{\beta}_{e})$ is inconsistent when $\delta=1$ and $E[\text{sign}(x_{2t-2})\varepsilon_{2t-1}^{x}]=0$. A canonical example is the pure random walk $x_{t}$, i.e., $x_{t}=x_{t-1}+\varepsilon_{t}^{x}$, with $\varepsilon_{t}^{x}\sim\text{i.i.d. }(0,\sigma^{2})$.
C Additional Simulation Results
C.1 Discrete-Time Experiments: One-Sided Tests
Following Section 5.2 of the main paper, we examine the finite-sample performance of the proposed tests in a discrete-time setting with an intercept, comparing them to the IVX test of Kostakis et al. (2015). The data-generating process (DGP) is specified as
$y_{t}$ $=\beta x_{t-1}+\sigma_{t}\varepsilon_{t},$ (S.2)
$x_{t}$ $=\left(1-\frac{\bar{\kappa}}{T}\right)x_{t-1}+\sigma_{t}\eta_{t},$
for $t=2,\ldots,12T$, where $T\in\{20,50,100\}$ corresponds to 20, 50, and 100 years of monthly data (i.e., 240, 600, and 1200 observations, respectively). The constant term in the predictive regression is set to zero. We set $\beta\in\{0.5k:k=0,1,\ldots,5\}$ and $\bar{\kappa}\in\{0,50,100\}$, and consider a one-sided test of $H_{0}\!:\beta=0$ against $H_{A}\!:\beta>0$. The number of replications is 10,000.
The innovation process $\eta_{t}$ follows an MA($q$) process:
$\eta_{t}=\sum_{j=0}^{q}C_{j}v_{t-j},$ (S.3)
where $(\varepsilon_{t},v_{t})$ are jointly normal with correlation 0.98-0.98. Specifically, we consider an MA(3) process with $C_{j}=1/2$ for $j=0,1,2,3$ (see Section 5.2 of the main paper for simulations with MA(1) innovations). The volatility process $\sigma_{t}$ follows the same specifications as in Section 5.2 of the main paper. Specifically, we consider
  • CNST (Constant volatility): $\sigma_{t}^{2}=\sigma_{0}^{2}$, with $\sigma_{0}=1$.
  • SB (Structural break): $\sigma_{t}=\sigma_{0}+(\sigma_{1}-\sigma_{0})1\{t/T\geq 4/5\}$, with $\sigma_{0}=1$ and $\sigma_{1}=4$.
  • RS (Regime switching): $\sigma_{t}=\sigma_{0}(1-s_{t})+\sigma_{1}s_{t}$, where $s_{t}$ is a two-state Markov chain independent of $(y_{t},x_{t})$, with transition matrix
    $P_{t}=\begin{pmatrix}0.8&0.2\\ 0.8&0.2\end{pmatrix}+\begin{pmatrix}0.2&-0.2\\ -0.8&0.8\end{pmatrix}\exp\!\left(-\frac{\bar{\lambda}}{T}t\right),$
    initialized at its invariant distribution, where $\bar{\lambda}=60$, $\sigma_{0}=1$, and $\sigma_{1}=4$.
We implement the hybrid tests based on even and odd observations, denoted by $\tau(\check{\beta}_{e})$ and $\tau(\check{\beta}_{o})$ (see Section 4.2 of the main paper). For comparison, we also include the IVX test of Kostakis et al. (2015), in which the intercept is included in the estimation.
The results, summarized in Tables S.1S.3, indicate that the proposed tests exhibit accurate size control under the null hypothesis across all DGPs, whereas the IVX test is substantially oversized, particularly when volatility is nonstationary or exhibits structural breaks. These findings are consistent with Demetrescu et al. (2023), who show that the IVX method exhibits severe size distortions in one-sided tests when regressors are highly persistent and endogenous. By contrast, our tests exhibit strong size control with reasonably good power in the simulation settings considered here.
Table S.1: Size and Power for Discrete-Time Models with Constant Volatility and MA(3) Innovations (One-Sided Tests)
$\bar{\kappa}=0$ $\bar{\kappa}=50$ $\bar{\kappa}=100$
T 20 50 100 20 50 100 20 50 100
$\beta=0$ $\tau(\check{\beta}_{o})$ 5.0 4.9 4.8 5.2 5.3 4.8 4.5 4.7 5.2
$\tau(\check{\beta}_{e})$ 4.8 4.7 5.2 5.0 4.8 4.8 5.1 5.0 4.9
IVX 14.2 13.5 12.3 10.2 10.5 10.3 9.5 10.3 10.5
$\beta=0.5$ $\tau(\check{\beta}_{o})$ 7.1 6.7 7.0 12.3 12.5 12.2 16.5 16.6 17.2
$\tau(\check{\beta}_{e})$ 7.0 7.0 7.2 12.4 11.5 12.3 16.8 17.2 17.5
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1$ $\tau(\check{\beta}_{o})$ 10.1 9.8 10.3 25.4 25.8 24.6 39.4 39.4 40.5
$\tau(\check{\beta}_{e})$ 10.3 10.0 10.4 25.4 24.2 25.4 40.3 40.3 40.8
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1.5$ $\tau(\check{\beta}_{o})$ 14.2 14.2 14.7 42.4 44.1 43.1 64.4 65.1 66.4
$\tau(\check{\beta}_{e})$ 14.4 14.4 14.4 42.8 42.5 42.8 65.9 66.9 66.5
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2$ $\tau(\check{\beta}_{o})$ 19.0 18.9 19.6 59.7 61.4 61.6 82.9 84.3 84.6
$\tau(\check{\beta}_{e})$ 19.4 19.1 19.1 59.5 60.2 60.6 83.4 84.7 84.8
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2.5$ $\tau(\check{\beta}_{o})$ 24.1 24.1 24.9 73.1 75.7 75.7 92.5 93.8 94.4
$\tau(\check{\beta}_{e})$ 24.5 24.0 24.6 73.7 74.3 75.5 92.8 94.1 94.1
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
This table reports the rejection rates of $\tau(\check{\beta}_{o})$ and $\tau(\check{\beta}_{e})$, as well as the IVX tests, based on simulated data under $\beta\in\{0,0.5,1,1.5,2,2.5\}$ (one-sided tests). The data are generated from the discrete-time model (S.2)–(S.3) with persistence parameter $\bar{\kappa}\in\{0,50,100\}$ and MA(3) innovations. The model is sampled at a monthly frequency, so that a sample of length $T$ years contains $12T$ observations, with $T\in\{20,50,100\}$. The volatility is constant.
Table S.2: Size and Power for Discrete-Time Models with Structural Breaks in Volatility and MA(3) Innovations (One-Sided Tests)
$\bar{\kappa}=0$ $\bar{\kappa}=50$ $\bar{\kappa}=100$
T 20 50 100 20 50 100 20 50 100
$\beta=0$ $\tau(\check{\beta}_{o})$ 4.6 4.9 5.0 4.5 5.0 5.0 4.3 5.0 5.0
$\tau(\check{\beta}_{e})$ 4.8 4.7 4.8 4.6 4.6 4.8 4.9 5.1 5.4
IVX 32.7 32.5 31.2 32.1 33.5 34.3 32.4 34.1 35.5
$\beta=0.5$ $\tau(\check{\beta}_{o})$ 9.8 10.1 11.0 7.9 8.9 9.0 9.0 11.4 12.0
$\tau(\check{\beta}_{e})$ 10.2 10.2 10.4 8.0 8.3 8.8 10.4 11.8 12.4
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1$ $\tau(\check{\beta}_{o})$ 18.9 19.2 20.8 13.9 16.0 15.7 18.1 23.3 25.3
$\tau(\check{\beta}_{e})$ 19.3 20.0 20.0 14.0 14.7 15.8 20.0 24.4 26.2
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1.5$ $\tau(\check{\beta}_{o})$ 29.3 29.7 30.9 22.0 25.4 25.5 31.0 39.5 43.1
$\tau(\check{\beta}_{e})$ 29.4 29.7 30.0 21.9 24.4 25.8 33.9 40.5 43.3
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2$ $\tau(\check{\beta}_{o})$ 38.4 38.1 39.8 31.5 36.6 36.8 45.5 56.6 60.8
$\tau(\check{\beta}_{e})$ 38.5 38.7 38.8 31.8 35.8 37.5 49.0 58.0 60.4
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2.5$ $\tau(\check{\beta}_{o})$ 45.9 45.2 46.9 41.8 47.2 48.6 59.1 70.3 74.9
$\tau(\check{\beta}_{e})$ 45.5 45.7 46.3 41.9 47.2 48.4 62.4 71.8 74.2
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
This table reports the rejection rates of $\tau(\check{\beta}_{o})$ and $\tau(\check{\beta}_{e})$, as well as the IVX tests, based on simulated data under $\beta\in\{0,0.5,1,1.5,2,2.5\}$ (one-sided tests). The data are generated from the discrete-time model (S.2)–(S.3) with persistence parameter $\bar{\kappa}\in\{0,50,100\}$ and MA(3) innovations. The model is sampled at a monthly frequency, so that a sample of length $T$ years contains $12T$ observations, with $T\in\{20,50,100\}$. The volatility exhibits structural breaks.
Table S.3: Size and Power for Discrete-Time Models with Regime Switching in Volatility with MA(3) Innovations (One-Sided Tests)
$\bar{\kappa}=0$ $\bar{\kappa}=50$ $\bar{\kappa}=100$
T 20 50 100 20 50 100 20 50 100
$\beta=0$ $\tau(\check{\beta}_{o})$ 3.8 3.9 4.3 4.3 5.1 5.3 4.6 4.8 4.9
$\tau(\check{\beta}_{e})$ 4.0 4.2 4.3 4.8 4.9 4.7 4.9 4.8 4.5
IVX 15.6 13.6 13.4 9.8 10.5 11.1 9.6 11.2 12.4
$\beta=0.5$ $\tau(\check{\beta}_{o})$ 7.8 7.9 8.8 10.2 11.6 11.5 12.4 14.5 15.8
$\tau(\check{\beta}_{e})$ 8.5 8.7 8.7 10.5 11.4 11.9 12.7 14.4 14.9
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1$ $\tau(\check{\beta}_{o})$ 15.3 15.4 17.0 22.7 24.9 24.1 27.2 34.7 36.4
$\tau(\check{\beta}_{e})$ 16.2 16.8 16.9 22.9 24.8 24.0 28.5 34.4 35.7
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1.5$ $\tau(\check{\beta}_{o})$ 24.1 24.2 26.8 39.3 42.3 41.9 47.5 58.9 61.5
$\tau(\check{\beta}_{e})$ 25.4 25.9 26.3 38.3 41.5 41.5 47.6 59.1 60.6
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2$ $\tau(\check{\beta}_{o})$ 32.9 33.4 35.4 55.1 59.0 59.1 64.5 77.4 80.3
$\tau(\check{\beta}_{e})$ 34.3 35.1 35.0 53.1 58.4 58.5 64.7 77.3 79.3
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2.5$ $\tau(\check{\beta}_{o})$ 40.9 41.0 42.4 67.3 71.8 72.4 76.2 87.7 91.3
$\tau(\check{\beta}_{e})$ 42.3 42.4 42.3 64.5 71.5 71.8 76.0 87.6 89.9
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
This table reports the rejection rates of $\tau(\check{\beta}_{o})$ and $\tau(\check{\beta}_{e})$, as well as the IVX tests, based on simulated data under $\beta\in\{0,0.5,1,1.5,2,2.5\}$ (one-sided tests). The data are generated from the discrete-time model (S.2)–(S.3) with persistence parameter $\bar{\kappa}\in\{0,50,100\}$ and MA(3) innovations. The model is sampled at a monthly frequency, so that a sample of length $T$ years contains $12T$ observations, with $T\in\{20,50,100\}$. The volatility exhibits regime switching.
C.2 Discrete-Time Experiments: Two-Sided Tests
We now consider two-sided tests. The DGP and simulation settings are the same as those for the one-sided tests in Section B.1, but we additionally consider an MA(1) process with $C_{0}=C_{1}=1/\sqrt{2}$. For the volatility specifications, we consider only constant volatility and structural breaks to conserve space, as regime-switching volatility yields qualitatively similar results to the structural break case.
The results, summarized in Tables S.4S.7, indicate that the proposed tests exhibit accurate size control under the null hypothesis across all DGPs, whereas the IVX test is substantially oversized when volatility exhibits structural breaks. Our simulation results further show that the IVX approach exhibits severe size distortions not only in one-sided tests but also in two-sided tests when volatility is nonstationary, whereas our methods exhibit accurate size control with reasonably good power in finite samples.
However, we do not argue that our methods should be preferred to the IVX method in the discrete-time predictive regressions. Under constant volatility—the setting for which the IVX method was originally designed—the IVX method exhibits substantially higher power than our methods. In this sense, as emphasized in the main paper, our methods and the IVX method of Kostakis et al. (2015) (as well as the Cauchy RT test of Choi et al. (2016) and the Cauchy VC test of Ibragimov et al. (2024)) should be viewed as complementary rather than competing approaches.
Table S.4: Size and Power for Discrete-Time Models with Constant Volatility and MA(1) Innovations (Two-Sided Tests)
$\bar{\kappa}=0$ $\bar{\kappa}=50$ $\bar{\kappa}=100$
T 20 50 100 20 50 100 20 50 100
$\beta=0$ $\tau(\check{\beta}_{o})$ 4.9 5.1 5.1 5.2 5.2 4.8 4.7 4.8 5.4
$\tau(\check{\beta}_{e})$ 5.0 5.1 5.0 5.1 5.2 5.0 5.0 4.5 5.0
IVX 7.1 6.6 5.9 5.2 5.5 5.2 4.6 5.1 5.4
$\beta=0.5$ $\tau(\check{\beta}_{o})$ 6.1 6.0 6.3 10.6 10.9 10.6 16.5 16.5 17.2
$\tau(\check{\beta}_{e})$ 6.0 5.9 6.1 10.7 10.2 10.6 16.6 16.9 17.1
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1$ $\tau(\check{\beta}_{o})$ 9.5 9.0 9.7 29.2 29.8 28.4 49.6 49.8 50.4
$\tau(\check{\beta}_{e})$ 9.5 9.2 9.5 29.1 27.8 28.6 50.5 50.7 51.4
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1.5$ $\tau(\check{\beta}_{o})$ 14.9 14.7 15.3 52.9 54.7 53.9 79.6 80.5 81.0
$\tau(\check{\beta}_{e})$ 14.8 14.5 14.6 53.4 53.1 53.8 80.0 81.4 81.4
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2$ $\tau(\check{\beta}_{o})$ 21.2 21.0 21.8 73.0 75.3 75.1 93.0 94.3 95.1
$\tau(\check{\beta}_{e})$ 21.1 20.9 21.2 73.0 74.0 74.8 93.5 94.8 95.0
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2.5$ $\tau(\check{\beta}_{o})$ 27.5 27.1 28.4 85.3 87.3 87.9 97.7 98.4 98.8
$\tau(\check{\beta}_{e})$ 27.8 27.5 28.1 85.0 86.7 87.7 98.0 98.5 98.8
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
This table reports the rejection rates of $\tau(\check{\beta}_{o})$ and $\tau(\check{\beta}_{e})$, as well as the IVX tests, based on simulated data under $\beta\in\{0,0.5,1,1.5,2,2.5\}$ (two-sided tests). The data are generated from the discrete-time model (S.2)–(S.3) with persistence parameter $\bar{\kappa}\in\{0,50,100\}$ and MA(1) innovations. The model is sampled at a monthly frequency, so that a sample of length $T$ years contains $12T$ observations, with $T\in\{20,50,100\}$. The volatility is constant.
Table S.5: Size and Power for Discrete-Time Models with Structural Breaks in Volatility and MA(1) Innovations (Two-Sided Tests)
$\bar{\kappa}=0$ $\bar{\kappa}=50$ $\bar{\kappa}=100$
T 20 50 100 20 50 100 20 50 100
$\beta=0$ $\tau(\check{\beta}_{o})$ 4.9 5.1 5.3 5.0 4.9 5.0 5.0 4.8 5.0
$\tau(\check{\beta}_{e})$ 4.7 5.0 5.3 5.0 5.0 5.1 4.6 4.7 5.1
IVX 20.7 20.9 19.9 24.6 25.8 25.6 25.3 25.5 26.9
$\beta=0.5$ $\tau(\check{\beta}_{o})$ 5.7 6.1 6.2 8.3 8.8 8.2 12.2 11.7 12.3
$\tau(\check{\beta}_{e})$ 6.1 6.1 6.1 8.2 8.1 7.9 12.1 12.3 12.6
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1$ $\tau(\check{\beta}_{o})$ 8.9 9.2 10.1 20.6 19.9 18.9 33.4 33.8 34.3
$\tau(\check{\beta}_{e})$ 9.4 9.6 9.8 19.7 19.2 18.3 34.5 35.3 34.8
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1.5$ $\tau(\check{\beta}_{o})$ 14.5 14.5 15.5 36.8 37.4 36.3 58.8 60.7 62.0
$\tau(\check{\beta}_{e})$ 15.0 15.0 15.3 36.0 35.5 35.5 61.6 61.9 62.3
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2$ $\tau(\check{\beta}_{o})$ 21.3 21.1 21.9 53.0 55.0 54.5 78.2 80.8 81.9
$\tau(\check{\beta}_{e})$ 21.7 21.3 21.6 53.2 53.2 54.5 79.3 81.3 81.8
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2.5$ $\tau(\check{\beta}_{o})$ 27.5 27.2 28.8 66.4 69.5 69.0 88.3 91.4 92.4
$\tau(\check{\beta}_{e})$ 28.2 27.7 27.8 66.8 67.6 69.1 89.5 91.4 92.1
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
This table reports the rejection rates of $\tau(\check{\beta}_{o})$ and $\tau(\check{\beta}_{e})$, as well as the IVX tests, based on simulated data under $\beta\in\{0,0.5,1,1.5,2,2.5\}$ (two-sided tests). The data are generated from the discrete-time model (S.2)–(S.3) with persistence parameter $\bar{\kappa}\in\{0,50,100\}$ and MA(1) innovations. The model is sampled at a monthly frequency, so that a sample of length $T$ years contains $12T$ observations, with $T\in\{20,50,100\}$. The volatility exhibits structural breaks.
Table S.6: Size and Power for Discrete-Time Models with Constant Volatility and MA(3) Innovations (Two-Sided Tests)
$\bar{\kappa}=0$ $\bar{\kappa}=50$ $\bar{\kappa}=100$
T 20 50 100 20 50 100 20 50 100
$\beta=0$ $\tau(\check{\beta}_{o})$ 5.0 4.8 5.1 4.8 4.9 5.0 4.8 4.8 5.1
$\tau(\check{\beta}_{e})$ 4.8 5.1 5.0 5.0 4.9 4.9 4.6 4.9 5.3
IVX 8.0 7.0 6.3 4.9 5.3 5.3 4.8 4.8 5.5
$\beta=0.5$ $\tau(\check{\beta}_{o})$ 7.5 7.2 7.4 6.6 7.4 6.9 6.8 9.0 9.7
$\tau(\check{\beta}_{e})$ 7.2 7.2 7.4 6.8 6.6 7.3 7.6 9.4 10.0
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1$ $\tau(\check{\beta}_{o})$ 13.9 14.0 14.6 13.1 14.7 14.6 17.6 23.1 25.4
$\tau(\check{\beta}_{e})$ 14.0 14.1 14.4 12.8 13.2 15.3 17.8 23.4 25.5
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1.5$ $\tau(\check{\beta}_{o})$ 22.7 22.8 24.1 23.8 27.5 27.8 35.6 45.8 49.6
$\tau(\check{\beta}_{e})$ 22.8 23.4 23.8 23.5 25.6 28.2 35.9 46.3 48.7
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2$ $\tau(\check{\beta}_{o})$ 31.3 31.4 32.8 38.2 43.6 44.7 55.8 68.0 71.9
$\tau(\check{\beta}_{e})$ 31.6 31.9 33.0 37.8 41.1 44.3 56.0 67.8 71.0
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2.5$ $\tau(\check{\beta}_{o})$ 38.5 38.8 40.1 52.3 59.1 60.7 72.8 83.5 86.3
$\tau(\check{\beta}_{e})$ 38.7 39.3 40.4 52.0 56.9 60.5 72.4 83.1 86.0
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
This table reports the rejection rates of $\tau(\check{\beta}_{o})$ and $\tau(\check{\beta}_{e})$, as well as the IVX tests, based on simulated data under $\beta\in\{0,0.5,1,1.5,2,2.5\}$ (two-sided tests). The data are generated from the discrete-time model (S.2)–(S.3) with persistence parameter $\bar{\kappa}\in\{0,50,100\}$ and MA(3) innovations. The model is sampled at a monthly frequency, so that a sample of length $T$ years contains $12T$ observations, with $T\in\{20,50,100\}$. The volatility is constant.
Table S.7: Size and Power for Discrete-Time Models with Structural Breaks in Volatility and MA(3) Innovations (Two-Sided Tests)
$\bar{\kappa}=0$ $\bar{\kappa}=50$ $\bar{\kappa}=100$
T 20 50 100 20 50 100 20 50 100
$\beta=0$ $\tau(\check{\beta}_{o})$ 4.8 5.1 5.2 5.0 5.0 5.0 4.5 4.7 5.1
$\tau(\check{\beta}_{e})$ 4.9 5.0 5.2 5.2 5.1 4.7 5.0 4.8 5.4
IVX 22.0 22.1 20.5 24.2 24.9 25.8 24.4 25.5 27.0
$\beta=0.5$ $\tau(\check{\beta}_{o})$ 6.8 7.3 7.4 5.5 5.9 6.2 5.6 7.3 7.5
$\tau(\check{\beta}_{e})$ 7.1 7.0 7.4 5.7 5.4 5.8 6.3 7.2 8.0
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1$ $\tau(\check{\beta}_{o})$ 13.4 13.7 15.0 8.7 9.9 9.8 11.0 15.0 16.6
$\tau(\check{\beta}_{e})$ 14.0 14.1 14.3 8.7 9.1 9.9 12.5 15.6 17.1
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=1.5$ $\tau(\check{\beta}_{o})$ 22.4 22.9 24.6 14.9 17.0 17.0 21.7 29.2 31.9
$\tau(\check{\beta}_{e})$ 23.2 23.6 23.6 15.1 16.0 17.3 23.4 29.7 32.6
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2$ $\tau(\check{\beta}_{o})$ 31.3 31.6 33.2 22.9 26.9 27.0 34.8 44.9 49.5
$\tau(\check{\beta}_{e})$ 31.6 32.0 32.4 23.0 25.6 27.3 37.6 46.3 49.6
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\beta=2.5$ $\tau(\check{\beta}_{o})$ 38.5 38.7 40.3 32.0 37.3 37.7 48.6 60.5 65.6
$\tau(\check{\beta}_{e})$ 39.1 39.5 39.9 32.3 36.2 38.5 51.6 62.0 65.2
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
This table reports the rejection rates of $\tau(\check{\beta}_{o})$ and $\tau(\check{\beta}_{e})$, as well as the IVX tests, based on simulated data under $\beta\in\{0,0.5,1,1.5,2,2.5\}$ (two-sided tests). The data are generated from the discrete-time model (S.2)–(S.3) with persistence parameter $\bar{\kappa}\in\{0,50,100\}$ and MA(3) innovations. The model is sampled at a monthly frequency, so that a sample of length $T$ years contains $12T$ observations, with $T\in\{20,50,100\}$. The volatility exhibits structural breaks.
C.3 Simulation Results for Joint Predictability Tests
In this subsection, we examine the finite-sample performance of the proposed Wald test in Section A.1 and compare it with the IVX test. Specifically, we consider a joint predictability testing problem with two predictors. The DGP is specified as
$y_{t}$ $=\beta_{1}x_{1,t-1}+\beta_{2}x_{2,t-1}+\sigma_{t}\varepsilon_{t},$ (S.4)
$x_{it}$ $=\left(1-\frac{\bar{\kappa}}{T}\right)x_{i,t-1}+\sigma_{t}\eta_{it},\quad i=1,2,$
for $t=2,\ldots,12T$, where $T\in\{20,50,100\}$ corresponds to 20, 50, and 100 years of monthly data (i.e., 240, 600, and 1200 observations, respectively). The constant term in the predictive regression is set to zero. We set $\beta_{1}=\beta_{2}$ with $\beta_{1}\in\{0.05k:k=0,1,\ldots,5\}$ and $\bar{\kappa}\in\{0,50,100\}$, and consider the joint test of $H_{0}\!:\beta_{1}=\beta_{2}=0$ against $H_{A}\!:\beta_{1}\neq 0\text{ or }\beta_{2}\neq 0$.
The innovation process $\eta_{it}$ follows an MA($q$) process:
$\eta_{it}=\sum_{j=0}^{q}C_{j}v_{i,t-j},$ (S.5)
where $(\varepsilon_{t},v_{1t},v_{2t})$ are jointly normal with the correlation matrix
$\begin{pmatrix}1.0000&-0.7596&-0.2787\\ -0.7596&1.0000&0.1246\\ -0.2787&0.1246&1.0000\end{pmatrix}$.
We consider two specifications for the innovation process: MA(1) with $C_{0}=C_{1}=1/\sqrt{2}$ and MA(3) with $C_{0}=C_{1}=C_{2}=C_{3}=1/2$. The number of replications is 1,000.
The volatility process $\sigma_{t}$ follows the same specifications as in Section B.1. The intercept is set to zero. The simulation design focuses on the case without an intercept in order to isolate the performance of the joint testing procedure from issues arising in predictive regressions with a nonzero intercept. While it is, in principle, possible to combine the Wald-type test with the first-differencing approach to accommodate a nonzero intercept, such an extension introduces an additional robustness–efficiency trade-off, as first differencing generally reduces power, as discussed in Section A.2.
Tables S.8 and S.9 report the results for the Wald test in Section A.1 and the IVX-Wald test of Kostakis et al. (2015) under MA(1) and MA(3) innovations, respectively. As in the tests for individual parameters in Sections B.1 and B.2, our method exhibits accurate size control with good power properties, whereas the IVX method suffers from severe size distortions when volatility exhibits structural breaks or regime switching.
Table S.8: Size and Power for Discrete-Time Models with MA(1) Innovations (Joint Tests)
$\bar{\kappa}=0$ $\bar{\kappa}=50$ $\bar{\kappa}=100$
T 20 50 100 20 50 100 20 50 100
Constant Volatility
$\beta=0$ Wald 6.6 8.3 5.5 6.4 4.6 5.4 5.6 5.4 5.6
IVX 7.6 7.0 6.2 4.8 6.2 5.6 6.0 4.8 5.4
$\beta=0.05$ Wald 98.3 99.6 99.8 44.7 100.0 100.0 25.8 91.9 100.0
IVX 100.0 100.0 100.0 64.5 100.0 100.0 38.4 99.2 100.0
$\beta=0.1$ Wald 99.4 99.7 100.0 97.4 100.0 100.0 76.4 100.0 100.0
IVX 100.0 100.0 100.0 99.8 100.0 100.0 93.2 100.0 100.0
$\beta=0.15$ Wald 99.6 99.7 100.0 100.0 100.0 100.0 98.6 100.0 100.0
IVX 100.0 100.0 100.0 100.0 100.0 100.0 99.9 100.0 100.0
$\beta=0.2$ Wald 99.6 99.8 100.0 100.0 100.0 100.0 100.0 100.0 100.0
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
Structural Breaks in Volatility
$\beta=0$ Wald 7.5 8.6 9.0 6.8 5.5 5.8 6.2 5.5 4.7
IVX 26.9 31.8 28.2 38.6 38.9 41.7 40.0 36.7 39.3
$\beta=0.05$ Wald 97.5 99.7 100.0 30.5 97.4 100.0 17.2 74.5 100.0
IVX 99.7 100.0 100.0 69.3 99.7 100.0 56.4 93.9 100.0
$\beta=0.1$ Wald 99.2 99.8 100.0 84.4 100.0 100.0 56.0 100.0 100.0
IVX 100.0 100.0 100.0 96.1 100.0 100.0 84.2 100.0 100.0
$\beta=0.15$ Wald 99.4 99.8 100.0 98.9 100.0 100.0 87.2 100.0 100.0
IVX 100.0 100.0 100.0 100.0 100.0 100.0 97.3 100.0 100.0
$\beta=0.2$ Wald 99.7 99.9 100.0 100.0 100.0 100.0 99.4 100.0 100.0
IVX 100.0 100.0 100.0 100.0 100.0 100.0 99.8 100.0 100.0
Regime Swithing in Volatility
$\beta=0$ Wald 9.4 9.8 7.8 6.2 5.3 5.2 5.6 4.7 5.5
IVX 12.3 12.2 10.8 15.6 16.7 15.4 17.1 16.8 19.0
$\beta=0.05$ Wald 98.0 99.7 100.0 39.9 99.1 100.0 24.8 85.7 100.0
IVX 99.9 100.0 100.0 68.5 99.9 100.0 44.7 97.8 100.0
$\beta=0.1$ Wald 99.4 99.9 100.0 93.4 100.0 100.0 67.9 100.0 100.0
IVX 100.0 100.0 100.0 99.1 100.0 100.0 89.3 100.0 100.0
$\beta=0.15$ Wald 99.7 99.9 100.0 100.0 100.0 100.0 96.5 100.0 100.0
IVX 100.0 100.0 100.0 100.0 100.0 100.0 99.1 100.0 100.0
$\beta=0.2$ Wald 99.7 99.9 100.0 100.0 100.0 100.0 99.9 100.0 100.0
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
This table reports the rejection rates of the Wald test in Section A.1 and the IVX-Wald test of Kostakis et al. (2015), based on simulated data under $\beta\in\{0,0.05,0.1,0.15,0.2\}$ (joint tests). The data are generated from the discrete-time model (S.4)–(S.5) with persistence parameter $\bar{\kappa}\in\{0,50,100\}$ and MA(1) innovations. The model is sampled at a monthly frequency, so that a sample of length $T$ years contains $12T$ observations, with $T\in\{20,50,100\}$.
Table S.9: Size and Power for Discrete-Time Models with MA(3) Innovations (Joint Tests)
$\bar{\kappa}=0$ $\bar{\kappa}=50$ $\bar{\kappa}=100$
T 20 50 100 20 50 100 20 50 100
Constant Volatility
$\beta=0$ Wald 6.8 7.9 6.4 5.9 3.9 4.7 5.7 6 6.2
IVX 8 7.2 6.6 4.9 5.6 5.7 6.1 5.9 5.4
$\beta=0.05$ Wald 98.9 99.7 100 67.3 100 100 34.7 99.7 100
IVX 100 100 100 88.1 100 100 49 100 100
$\beta=0.1$ Wald 99.3 99.8 100 99.8 100 100 91.4 100 100
IVX 100 100 100 100 100 100 98.3 100 100
$\beta=0.15$ Wald 99.4 99.8 100 100 100 100 99.8 100 100
IVX 100 100 100 100 100 100 100 100 100
$\beta=0.2$ Wald 99.5 99.8 100 100 100 100 100 100 100
IVX 100 100 100 100 100 100 100 100 100
Structural Breaks in Volatility
$\beta=0$ Wald 8.2 9.0 8.6 5.8 5.7 5.7 6.6 6.6 5.6
IVX 28.5 32.2 27.9 34.7 38.6 39.9 41.6 37.8 39.0
$\beta=0.05$ Wald 98.9 99.6 100.0 44.7 99.9 100.0 23.2 93.7 100.0
IVX 100.0 100.0 100.0 79.3 100.0 100.0 60.3 99.3 100.0
$\beta=0.1$ Wald 99.6 99.7 100.0 96.1 100.0 100.0 72.5 100.0 100.0
IVX 100.0 100.0 100.0 99.6 100.0 100.0 91.9 100.0 100.0
$\beta=0.15$ Wald 99.6 99.8 100.0 100.0 100.0 100.0 97.0 100.0 100.0
IVX 100.0 100.0 100.0 100.0 100.0 100.0 99.7 100.0 100.0
$\beta=0.2$ Wald 99.7 99.8 100.0 100.0 100.0 100.0 99.8 100.0 100.0
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
Regime Swithing in Volatility
$\beta=0$ Wald 10.2 9.3 8.7 5.8 4.4 4.8 5.7 5.8 5.5
IVX 12.8 12.5 11.5 13.2 16.1 16.1 17.1 16.5 19.7
$\beta=0.05$ Wald 98.6 99.7 100.0 60.7 100.0 100.0 30.1 98.7 100.0
IVX 100.0 100.0 100.0 86.1 100.0 100.0 51.7 99.7 100.0
$\beta=0.1$ Wald 99.5 99.7 100.0 99.6 100.0 100.0 84.2 100.0 100.0
IVX 100.0 100.0 100.0 99.9 100.0 100.0 96.0 100.0 100.0
$\beta=0.15$ Wald 99.6 99.8 100.0 100.0 100.0 100.0 99.2 100.0 100.0
IVX 100.0 100.0 100.0 100.0 100.0 100.0 99.9 100.0 100.0
$\beta=0.2$ Wald 99.6 99.8 100.0 100.0 100.0 100.0 99.9 100.0 100.0
IVX 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
This table reports the rejection rates of the Wald test in Section A.1 and the IVX-Wald test of Kostakis et al. (2015), based on simulated data under $\beta\in\{0,0.05,0.1,0.15,0.2\}$ (joint tests). The data are generated from the discrete-time model (S.4)–(S.5) with persistence parameter $\bar{\kappa}\in\{0,50,100\}$ and MA(3) innovations. The model is sampled at a monthly frequency, so that a sample of length $T$ years contains $12T$ observations, with $T\in\{20,50,100\}$.
D Additional Empirical Results
D.1 Two-Sided Tests with Intercept
In this subsection, we report empirical results for tests that account for the intercept. We use the same dataset as in Section 6 of the main paper. Table S.10 presents the results. In contrast to Table 9 in the main text, accounting for the intercept may affect the conclusions. In particular, the methods yield different outcomes: (i) the hybrid method with first differencing tends to support no predictability (2 and 1 rejections out of 18 tests for $\tau(\check{\beta}_{o})$ and $\tau(\check{\beta}_{e})$, respectively), and (ii) the IVX method rejects the null hypothesis more frequently than the hybrid method (7 rejections out of 18 tests). These mixed results may reflect the lower power of the hybrid method with first differencing or size distortions of the IVX method. As shown in Section B.2, the IVX method exhibits size distortions under structural breaks in volatility, even for two-sided tests, whereas the hybrid method exhibits good power properties, albeit with reduced power, especially when $\beta$ is close to zero.
Table S.10: Empirical Results on Stock Return Predictability (Two-Sided Tests)
Series Frequency $\tau(\check{\beta}_{o})$ $\tau(\check{\beta}_{e})$ IVX
Panel A: D/P as predictor for the period of 1927-2011
CRSP Monthly 0.51 0.51 1.95
Quarterly 0.71 0.67 2.52
Yearly 0.91 0.73 2.55
S&P500 Monthly 0.68 0.68 2.01
Quarterly 0.77 0.74 2.73∗∗
Yearly 0.68 0.86 1.93
Panel B: D/P as predictor for the period of 1927-2011 with jumps removed
CRSP Monthly -0.16 0.15 2.11
Quarterly 0.22 -0.27 2.79∗∗
Yearly 0.57 0.45 1.42
S&P500 Monthly 1.04 1.05 2.10
Quarterly 2.17 2.21 2.47
Yearly -0.22 -0.01 1.77
Panel C: E/P as predictor for the period of 1950-2011
S&P500 Monthly 0.35 -0.36 0.98
Quarterly 1.65 1.64 1.10
Yearly 1.14 1.21 1.17
Panel D: E/P as predictor for the period of 1950-2011 with jumps removed
S&P500 Monthly 0.88 0.88 0.76
Quarterly -0.14 0.08 0.93
Yearly 0.72 0.50 1.08
Results for two-sided tests of return predictability for the NYSE/AMEX value-weighted index (CRSP) and the S&P 500 using three intercept-robust tests, $\tau(\check{\beta}_{o})$, $\tau(\check{\beta}_{e})$, and the IVX method, across different regression frequencies. Panels A–B use the dividend–price ratio (D/P), while Panels C–D use the earnings–price ratio (E/P) as predictors. Statistical significance at the 5% and 1% levels is denoted by ‘‘’’ and ‘‘∗∗’’, respectively.
D.2 Joint Predictability: Dividend–Price and Earnings–Price Ratios
We now consider the joint predictability of the dividend–price ratio (D/P) and the earnings–price ratio (E/P) for stock returns. We use S&P 500 returns, as both D/P and E/P are readily available for this index. We implement two tests: the Wald-type test in Section A.1 and the IVX-Wald test of Kostakis et al. (2015), and test the null hypothesis of no predictability jointly. Table S.11 presents the results for monthly, quarterly, and yearly S&P 500 series using both the earnings–price and dividend–price ratios over the period 1950–2011. Recursive demeaning is employed as in Choi et al. (2016). In all cases, comparing with the corresponding $\chi^{2}$ critical values at the 1% significance level, we fail to reject the null hypothesis of no predictability.
Table S.11: Empirical Results on Stock Return Predictability (Joint Tests)
Wald IVX Wald IVX
With jumps With jumps removed
Monthly 1.47 1.43 0.72 1.13
Quarterly 0.54 2.47 1.40 0.88
Yearly 2.02 2.35 1.95 0.42
Results for joint tests of return predictability for the S&P 500 using the Wald test in Section A.1, and the IVX-Wald test of Kostakis et al. (2015), across different regression frequencies. The two predictors are the dividend–price ratio and the earnings–price ratio.